RTEMS POSIX API User’s Guide

Edition 4.10.0, for RTEMS 4.10.0

8 February 2011

On-Line Applications Research Corporation

On-Line Applications Research Corporation TEXinfo 2009-08-14.15

COPYRIGHT \begin{matrix}c \\ ◯ \end{matrix} 1988 - 2008. On-Line Applications Research Corporation (OAR).

The authors have used their best efforts in preparing this material. These efforts include the development, research, and testing of the theories and programs to determine their effectiveness. No warranty of any kind, expressed or implied, with regard to the software or the material contained in this document is provided. No liability arising out of the application or use of any product described in this document is assumed. The authors reserve the right to revise this material and to make changes from time to time in the content hereof without obligation to notify anyone of such revision or changes.

The RTEMS Project is hosted at http://www.rtems.com. Any inquiries concerning RTEMS, its related support components, its documentation, or any custom services for RTEMS should be directed to the contacts listed on that site. A current list of RTEMS

Support Providers is at http://www.rtems.com/support.html.

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Process Creation and Execution Manager …. 3

1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.4.1 fork - Create a Process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4.2 execl - Execute a File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.4.3 execv - Execute a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4.4 execle - Execute a File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.4.5 execve - Execute a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.4.6 execlp - Execute a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.4.7 execvp - Execute a File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.4.8 pthread atfork - Register Fork Handlers . . . . . . . . . . . . . . . . . . 11 1.4.9 wait - Wait for Process Termination . . . . . . . . . . . . . . . . . . . . . . 12 1.4.10 waitpid - Wait for Process Termination . . . . . . . . . . . . . . . . . . 13 1.4.11 exit - Terminate a Process. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2 Signal Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.2.1 Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2.2 Signal Delivery . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3.1 Signal Set Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3.2 Blocking Until Signal Generation . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3.3 Sending a Signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.4.1 sigaddset - Add a Signal to a Signal Set . . . . . . . . . . . . . . . . . . 17 2.4.2 sigdelset - Delete a Signal from a Signal Set. . . . . . . . . . . . . . . 18 2.4.3 sigfillset - Fill a Signal Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.4.4 sigismember - Is Signal a Member of a Signal Set . . . . . . . . . 20 2.4.5 sigemptyset - Empty a Signal Set . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.4.6 sigaction - Examine and Change Signal Action. . . . . . . . . . . . 22 2.4.7 pthread kill - Send a Signal to a Thread . . . . . . . . . . . . . . . . . . 23 2.4.8 sigprocmask - Examine and Change Process Blocked Signals

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.4.9 pthread sigmask - Examine and Change Thread Blocked

Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.4.10 kill - Send a Signal to a Process . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.4.11 sigpending - Examine Pending Signals . . . . . . . . . . . . . . . . . . . 27 2.4.12 sigsuspend - Wait for a Signal . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.4.13 pause - Suspend Process Execution . . . . . . . . . . . . . . . . . . . . . . 29 2.4.14 sigwait - Synchronously Accept a Signal . . . . . . . . . . . . . . . . . 30 2.4.15 sigwaitinfo - Synchronously Accept a Signal . . . . . . . . . . . . . 31 2.4.16 sigtimedwait - Synchronously Accept a Signal with Timeout

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.4.17 sigqueue - Queue a Signal to a Process . . . . . . . . . . . . . . . . . . 33 2.4.18 alarm - Schedule Alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.4.19 ualarm - Schedule Alarm in Microseconds . . . . . . . . . . . . . . . 35

3 Process Environment Manager . . . . . . . . . . . . . . . 37

3.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.2.1 Users and Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.2.2 User and Group Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.2.3 Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

3.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.3.1 Accessing User and Group Ids . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.3.2 Accessing Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . 38

3.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.4.1 getpid - Get Process ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.4.2 getppid - Get Parent Process ID . . . . . . . . . . . . . . . . . . . . . . . . . . 40 3.4.3 getuid - Get User ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 3.4.4 geteuid - Get Effective User ID . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.4.5 getgid - Get Real Group ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.4.6 getegid - Get Effective Group ID . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.4.7 setuid - Set User ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.4.8 setgid - Set Group ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.4.9 getgroups - Get Supplementary Group IDs. . . . . . . . . . . . . . . . 47 3.4.10 getlogin - Get User Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 3.4.11 getlogin r - Reentrant Get User Name . . . . . . . . . . . . . . . . . . . 49 3.4.12 getpgrp - Get Process Group ID . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.4.13 setsid - Create Session and Set Process Group ID . . . . . . . . 51 3.4.14 setpgid - Set Process Group ID for Job Control . . . . . . . . . . 52 3.4.15 uname - Get System Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 3.4.16 times - Get process times. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.4.17 getenv - Get Environment Variables . . . . . . . . . . . . . . . . . . . . . 55 3.4.18 setenv - Set Environment Variables . . . . . . . . . . . . . . . . . . . . . . 56 3.4.19 ctermid - Generate Terminal Pathname. . . . . . . . . . . . . . . . . . 57 3.4.20 ttyname - Determine Terminal Device Name. . . . . . . . . . . . . 58 3.4.21 ttyname r - Reentrant Determine Terminal Device Name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.4.22 isatty - Determine if File Descriptor is Terminal . . . . . . . . . 60 3.4.23 sysconf - Get Configurable System Variables . . . . . . . . . . . . . 61

4 Files and Directories Manager . . . . . . . . . . . . . . . 63

4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 4.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

4.2.1 Path Name Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 4.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 4.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

4.4.1 opendir - Open a Directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 4.4.2 readdir - Reads a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.4.3 rewinddir - Resets the readdir() pointer. . . . . . . . . . . . . . . . . . . 67 4.4.4 scandir - Scan a directory for matching entries . . . . . . . . . . . . 68 4.4.5 telldir - Return current location in directory stream. . . . . . . 69 4.4.6 closedir - Ends directory read operation . . . . . . . . . . . . . . . . . . 70 4.4.7 chdir - Changes the current working directory . . . . . . . . . . . . 71 4.4.8 fchdir - Changes the current working directory . . . . . . . . . . . . 72 4.4.9 getcwd - Gets current working directory . . . . . . . . . . . . . . . . . . 73 4.4.10 open - Opens a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.4.11 creat - Create a new file or rewrite an existing one . . . . . . . 76 4.4.12 umask - Sets a file creation mask. . . . . . . . . . . . . . . . . . . . . . . . 77 4.4.13 link - Creates a link to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.4.14 symlink - Creates a symbolic link to a file . . . . . . . . . . . . . . . 79 4.4.15 readlink - Obtain the name of a symbolic link destination

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.4.16 mkdir - Makes a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.4.17 mkfifo - Makes a FIFO special file . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.18 unlink - Removes a directory entry . . . . . . . . . . . . . . . . . . . . . . 83 4.4.19 rmdir - Delete a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.4.20 rename - Renames a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.4.21 stat - Gets information about a file . . . . . . . . . . . . . . . . . . . . . . 86 4.4.22 fstat - Gets file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.4.23 lstat - Gets file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.4.24 access - Check permissions for a file . . . . . . . . . . . . . . . . . . . . . 89 4.4.25 chmod - Changes file mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.4.26 fchmod - Changes permissions of a file . . . . . . . . . . . . . . . . . . . 91 4.4.27 getdents - Get directory entries . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.4.28 chown - Changes the owner and/or group of a file. . . . . . . . 93 4.4.29 utime - Change access and/or modification times of an inode

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.4.30 ftruncate - truncate a file to a specified length . . . . . . . . . . . 95 4.4.31 truncate - truncate a file to a specified length . . . . . . . . . . . . 96 4.4.32 pathconf - Gets configuration values for files . . . . . . . . . . . . . 97 4.4.33 fpathconf - Gets configuration values for files . . . . . . . . . . . . 99 4.4.34 mknod - create a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

5 Input and Output Primitives Manager . . . . . 103

5.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 5.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 5.3 Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 5.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

5.4.1 pipe - Create an Inter-Process Channel . . . . . . . . . . . . . . . . . . 104 5.4.2 dup - Duplicates an open file descriptor. . . . . . . . . . . . . . . . . . 105 5.4.3 dup2 - Duplicates an open file descriptor . . . . . . . . . . . . . . . . 106 5.4.4 close - Closes a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 5.4.5 read - Reads from a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 5.4.6 write - Writes to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 5.4.7 fcntl - Manipulates an open file descriptor . . . . . . . . . . . . . . . 111 5.4.8 lseek - Reposition read/write file offset . . . . . . . . . . . . . . . . . . 113 5.4.9 fsync - Synchronize file complete in-core state with that on

disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 5.4.10 fdatasync - Synchronize file in-core data with that on disk

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 5.4.11 sync - Schedule file system updates . . . . . . . . . . . . . . . . . . . . . 116 5.4.12 mount - Mount a file system . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 5.4.13 unmount - Unmount file systems . . . . . . . . . . . . . . . . . . . . . . . 118 5.4.14 readv - Vectored read from a file . . . . . . . . . . . . . . . . . . . . . . . 119 5.4.15 writev - Vectored write to a file . . . . . . . . . . . . . . . . . . . . . . . . 120 5.4.16 aio read - Asynchronous Read. . . . . . . . . . . . . . . . . . . . . . . . . . 121 5.4.17 aio write - Asynchronous Write . . . . . . . . . . . . . . . . . . . . . . . . 122 5.4.18 lio listio - List Directed I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 5.4.19 aio error - Retrieve Error Status of Asynchronous I/O

Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 5.4.20 aio return - Retrieve Return Status Asynchronous I/O

Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 5.4.21 aio cancel - Cancel Asynchronous I/O Request . . . . . . . . . 126 5.4.22 aio suspend - Wait for Asynchronous I/O Request . . . . . . 127 5.4.23 aio fsync - Asynchronous File Synchronization. . . . . . . . . . 128

6 Device- and Class- Specific Functions Manager

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

6.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.3 Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 6.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

6.4.1 cfgetispeed - Reads terminal input baud rate . . . . . . . . . . . . 130 6.4.2 cfgetospeed - Reads terminal output baud rate. . . . . . . . . . . 131 6.4.3 cfsetispeed - Sets terminal input baud rate. . . . . . . . . . . . . . . 132 6.4.4 cfsetospeed - Sets terminal output baud rate. . . . . . . . . . . . . 133 6.4.5 tcgetattr - Gets terminal attributes . . . . . . . . . . . . . . . . . . . . . . 134 6.4.6 tcsetattr - Set terminal attributes . . . . . . . . . . . . . . . . . . . . . . . 135 6.4.7 tcsendbreak - Sends a break to a terminal . . . . . . . . . . . . . . . 136 6.4.8 tcdrain - Waits for all output to be transmitted to the

terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 6.4.9 tcflush - Discards terminal data . . . . . . . . . . . . . . . . . . . . . . . . . 138 6.4.10 tcflow - Suspends/restarts terminal output. . . . . . . . . . . . . . 139 6.4.11 tcgetpgrp - Gets foreground process group ID . . . . . . . . . . 140 6.4.12 tcsetpgrp - Sets foreground process group ID . . . . . . . . . . . 141

7 Language-Specific Services for the C

Programming Language Manager . . . . . . . . . 143

7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.3 Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

7.4.1 setlocale - Set the Current Locale . . . . . . . . . . . . . . . . . . . . . . . 144 7.4.2 fileno - Obtain File Descriptor Number for this File . . . . . . 145 7.4.3 fdopen - Associate Stream with File Descriptor . . . . . . . . . . 146 7.4.4 flockfile - Acquire Ownership of File Stream . . . . . . . . . . . . . 147 7.4.5 ftrylockfile - Poll to Acquire Ownership of File Stream … 148 7.4.6 funlockfile - Release Ownership of File Stream . . . . . . . . . . . 149 7.4.7 getc unlocked - Get Character without Locking . . . . . . . . . . 150 7.4.8 getchar unlocked - Get Character from stdin without Locking

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.4.9 putc unlocked - Put Character without Locking . . . . . . . . . 152 7.4.10 putchar unlocked - Put Character to stdin without Locking

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.4.11 setjmp - Save Context for Non-Local Goto. . . . . . . . . . . . . . 154 7.4.12 longjmp - Non-Local Jump to a Saved Context . . . . . . . . . 155 7.4.13 sigsetjmp - Save Context with Signal Status for Non-Local

Goto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.4.14 siglongjmp - Non-Local Jump with Signal Status to a Saved

Context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 7.4.15 tzset - Initialize Time Conversion Information . . . . . . . . . . 158 7.4.16 strtok r - Reentrant Extract Token from String . . . . . . . . . 159 7.4.17 asctime r - Reentrant struct tm to ASCII Time Conversion

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 7.4.18 ctime r - Reentrant time t to ASCII Time Conversion . . 161 7.4.19 gmtime r - Reentrant UTC Time Conversion . . . . . . . . . . . 162 7.4.20 localtime r - Reentrant Local Time Conversion . . . . . . . . . 163 7.4.21 rand r - Reentrant Random Number Generation. . . . . . . . 164

8 System Databases Manager . . . . . . . . . . . . . . . . . 165

8.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 8.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 8.3 Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 8.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

8.4.1 getgrgid - Get Group File Entry for ID . . . . . . . . . . . . . . . . . . 166 8.4.2 getgrgid r - Reentrant Get Group File Entry . . . . . . . . . . . . 167 8.4.3 getgrnam - Get Group File Entry for Name. . . . . . . . . . . . . . 168 8.4.4 getgrnam r - Reentrant Get Group File Entry for Name . . 169 8.4.5 getpwuid - Get Password File Entry for UID. . . . . . . . . . . . . 170 8.4.6 getpwuid r - Reentrant Get Password File Entry for UID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.4.7 getpwnam - Password File Entry for Name . . . . . . . . . . . . . . 172 8.4.8 getpwnam r - Reentrant Get Password File Entry for Name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

9 Semaphore Manager . . . . . . . . . . . . . . . . . . . . . . . . . 175

9.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 9.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

9.2.1 Theory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 9.2.2 "sem t" Structure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 9.2.3 Building a Semaphore Attribute Set . . . . . . . . . . . . . . . . . . . . . 175

9.3 Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

9.3.1 Using as a Binary Semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 9.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

9.4.1 sem init - Initialize an unnamed semaphore. . . . . . . . . . . . . . 177 9.4.2 sem destroy - Destroy an unnamed semaphore . . . . . . . . . . . 178 9.4.3 sem open - Open a named semaphore . . . . . . . . . . . . . . . . . . . 179 9.4.4 sem close - Close a named semaphore . . . . . . . . . . . . . . . . . . . 180 9.4.5 sem unlink - Unlink a semaphore . . . . . . . . . . . . . . . . . . . . . . . . 181 9.4.6 sem wait - Wait on a Semaphore . . . . . . . . . . . . . . . . . . . . . . . . 182 9.4.7 sem trywait - Non-blocking Wait on a Semaphore . . . . . . . 183 9.4.8 sem timedwait - Wait on a Semaphore for a Specified Time

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 9.4.9 sem post - Unlock a Semaphore . . . . . . . . . . . . . . . . . . . . . . . . . 185 9.4.10 sem getvalue - Get the value of a semaphore . . . . . . . . . . . 186

10 Mutex Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

10.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 10.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

10.2.1 Mutex Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 10.2.2 PTHREAD MUTEX INITIALIZER . . . . . . . . . . . . . . . . . . . 187

10.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 10.4 Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

10.4.1 pthread mutexattr init - Initialize a Mutex Attribute Set

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 10.4.2 pthread mutexattr destroy - Destroy a Mutex Attribute Set

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 10.4.3 pthread mutexattr setprotocol - Set the Blocking Protocol

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 10.4.4 pthread mutexattr getprotocol - Get the Blocking Protocol

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 10.4.5 pthread mutexattr setprioceiling - Set the Priority Ceiling

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 10.4.6 pthread mutexattr getprioceiling - Get the Priority Ceiling

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 10.4.7 pthread mutexattr setpshared - Set the Visibility. . . . . . . 195 10.4.8 pthread mutexattr getpshared - Get the Visibility . . . . . . 196 10.4.9 pthread mutex init - Initialize a Mutex. . . . . . . . . . . . . . . . . 197 10.4.10 pthread mutex destroy - Destroy a Mutex. . . . . . . . . . . . . 198 10.4.11 pthread mutex lock - Lock a Mutex. . . . . . . . . . . . . . . . . . . 199 10.4.12 pthread mutex trylock - Poll to Lock a Mutex . . . . . . . . 200 10.4.13 pthread mutex timedlock - Lock a Mutex with Timeout

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 10.4.14 pthread mutex unlock - Unlock a Mutex . . . . . . . . . . . . . . 202 10.4.15 pthread mutex setprioceiling - Dynamically Set the

Priority Ceiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 10.4.16 pthread mutex getprioceiling - Get the Current Priority

Ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

11 Condition Variable Manager . . . . . . . . . . . . . . . 205

11.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 11.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 11.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 11.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

11.4.1 pthread condattr init - Initialize a Condition Variable

Attribute Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 11.4.2 pthread condattr destroy - Destroy a Condition Variable

Attribute Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 11.4.3 pthread condattr setpshared - Set Process Shared Attribute

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 11.4.4 pthread condattr getpshared - Get Process Shared Attribute

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 11.4.5 pthread cond init - Initialize a Condition Variable . . . . . . 210 11.4.6 pthread cond destroy - Destroy a Condition Variable … 211 11.4.7 pthread cond signal - Signal a Condition Variable . . . . . . 212 11.4.8 pthread cond broadcast - Broadcast a Condition Variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 11.4.9 pthread cond wait - Wait on a Condition Variable. . . . . . 214 11.4.10 pthread cond timedwait - Wait with Timeout a Condition

Variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

12 Memory Management Manager . . . . . . . . . . . 217

12.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 12.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 12.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 12.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

12.4.1 mlockall - Lock the Address Space of a Process . . . . . . . . . 218 12.4.2 munlockall - Unlock the Address Space of a Process …. 219 12.4.3 mlock - Lock a Range of the Process Address Space …. 220 12.4.4 munlock - Unlock a Range of the Process Address Space

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 12.4.5 mmap - Map Process Addresses to a Memory Object … 222 12.4.6 munmap - Unmap Previously Mapped Addresses . . . . . . . 223 12.4.7 mprotect - Change Memory Protection . . . . . . . . . . . . . . . . . 224 12.4.8 msync - Memory Object Synchronization . . . . . . . . . . . . . . . 225 12.4.9 shm open - Open a Shared Memory Object. . . . . . . . . . . . . 226 12.4.10 shm unlink - Remove a Shared Memory Object . . . . . . . 227

13 Scheduler Manager . . . . . . . . . . . . . . . . . . . . . . . . . 229

13.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 13.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

13.2.1 Priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 13.2.2 Scheduling Policies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

13.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 13.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230

13.4.1 sched get priority min - Get Minimum Priority Value . . 231 13.4.2 sched get priority max - Get Maximum Priority Value . . 232 13.4.3 sched rr get interval - Get Timeslicing Quantum . . . . . . . 233 13.4.4 sched yield - Yield the Processor . . . . . . . . . . . . . . . . . . . . . . . 234

14 Clock Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

14.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 14.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 14.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 14.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

14.4.1 clock gettime - Obtain Time of Day . . . . . . . . . . . . . . . . . . . . 235 14.4.2 clock settime - Set Time of Day . . . . . . . . . . . . . . . . . . . . . . . . 237 14.4.3 clock getres - Get Clock Resolution . . . . . . . . . . . . . . . . . . . . 238 14.4.4 sleep - Delay Process Execution . . . . . . . . . . . . . . . . . . . . . . . . 239 14.4.5 usleep - Delay Process Execution in Microseconds . . . . . . 240 14.4.6 nanosleep - Delay with High Resolution . . . . . . . . . . . . . . . . 241 14.4.7 gettimeofday - Get the Time of Day . . . . . . . . . . . . . . . . . . . . 242 14.4.8 time - Get time in seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243

15 Timer Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

15.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 15.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 15.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 15.4 System Calls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

15.4.1 timer create - Create a Per-Process Timer . . . . . . . . . . . . . . 246 15.4.2 timer delete - Delete a Per-Process Timer . . . . . . . . . . . . . . 247 15.4.3 timer settime - Set Next Timer Expiration . . . . . . . . . . . . . 248 15.4.4 timer gettime - Get Time Remaining on Timer . . . . . . . . . 249 15.4.5 timer getoverrun - Get Timer Overrun Count . . . . . . . . . . 250

16 Message Passing Manager. . . . . . . . . . . . . . . . . . 251

16.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 16.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251

16.2.1 Theory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 16.2.2 Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 16.2.3 Message Queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 16.2.4 Building a Message Queue Attribute Set. . . . . . . . . . . . . . . . 252 16.2.5 Notification of a Message on the Queue. . . . . . . . . . . . . . . . . 252 16.2.6 POSIX Interpretation Issues . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

16.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 16.3.1 Opening or Creating a Message Queue . . . . . . . . . . . . . . . . . 253 16.3.2 Closing a Message Queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 16.3.3 Removing a Message Queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 16.3.4 Sending a Message to a Message Queue. . . . . . . . . . . . . . . . . 253 16.3.5 Receiving a Message from a Message Queue . . . . . . . . . . . . 254 16.3.6 Notification of Receipt of a Message on an Empty Queue

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 16.3.7 Setting the Attributes of a Message Queue . . . . . . . . . . . . . 254 16.3.8 Getting the Attributes of a Message Queue . . . . . . . . . . . . . 255

16.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 16.4.1 mq open - Open a Message Queue . . . . . . . . . . . . . . . . . . . . . 256 16.4.2 mq close - Close a Message Queue. . . . . . . . . . . . . . . . . . . . . . 258 16.4.3 mq unlink - Remove a Message Queue. . . . . . . . . . . . . . . . . . 259 16.4.4 mq send - Send a Message to a Message Queue . . . . . . . . . 260 16.4.5 mq receive - Receive a Message from a Message Queue . . 261 16.4.6 mq notify - Notify Process that a Message is Available . . 262 16.4.7 mq setattr - Set Message Queue Attributes . . . . . . . . . . . . . 263 16.4.8 mq getattr - Get Message Queue Attributes . . . . . . . . . . . . 264

17 Thread Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265

17.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 17.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265

17.2.1 Thread Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 17.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 17.4 Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266

17.4.1 pthread attr init - Initialize a Thread Attribute Set …. 267 17.4.2 pthread attr destroy - Destroy a Thread Attribute Set . . 268 17.4.3 pthread attr setdetachstate - Set Detach State . . . . . . . . . 269 17.4.4 pthread attr getdetachstate - Get Detach State . . . . . . . . 270 17.4.5 pthread attr setstacksize - Set Thread Stack Size . . . . . . . 271 17.4.6 pthread attr getstacksize - Get Thread Stack Size . . . . . . 272 17.4.7 pthread attr setstackaddr - Set Thread Stack Address . . 273 17.4.8 pthread attr getstackaddr - Get Thread Stack Address

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 17.4.9 pthread attr setscope - Set Thread Scheduling Scope … 275 17.4.10 pthread attr getscope - Get Thread Scheduling Scope

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 17.4.11 pthread attr setinheritsched - Set Inherit Scheduler Flag

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 17.4.12 pthread attr getinheritsched - Get Inherit Scheduler Flag

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 17.4.13 pthread attr setschedpolicy - Set Scheduling Policy. … 279 17.4.14 pthread attr getschedpolicy - Get Scheduling Policy … 280 17.4.15 pthread attr setschedparam - Set Scheduling Parameters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 17.4.16 pthread attr getschedparam - Get Scheduling Parameters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 17.4.17 pthread create - Create a Thread . . . . . . . . . . . . . . . . . . . . . 283 17.4.18 pthread exit - Terminate the Current Thread. . . . . . . . . . 285 17.4.19 pthread detach - Detach a Thread . . . . . . . . . . . . . . . . . . . . 286 17.4.20 pthread join - Wait for Thread Termination . . . . . . . . . . . 287 17.4.21 pthread self - Get Thread ID . . . . . . . . . . . . . . . . . . . . . . . . . 288 17.4.22 pthread equal - Compare Thread IDs . . . . . . . . . . . . . . . . . 289 17.4.23 pthread once - Dynamic Package Initialization . . . . . . . . 290 17.4.24 pthread setschedparam - Set Thread Scheduling

Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 17.4.25 pthread getschedparam - Get Thread Scheduling

Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292

18 Key Manager. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

18.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 18.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 18.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 18.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

18.4.1 pthread key create - Create Thread Specific Data Key . . 294 18.4.2 pthread key delete - Delete Thread Specific Data Key . . 295 18.4.3 pthread setspecific - Set Thread Specific Key Value. . . . . 296 18.4.4 pthread getspecific - Get Thread Specific Key Value …. 297

19 Thread Cancellation Manager . . . . . . . . . . . . . 299

19.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 19.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 19.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 19.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299

19.4.1 pthread cancel - Cancel Execution of a Thread . . . . . . . . . 300 19.4.2 pthread setcancelstate - Set Cancelability State . . . . . . . . 301 19.4.3 pthread setcanceltype - Set Cancelability Type . . . . . . . . . 302 19.4.4 pthread testcancel - Create Cancellation Point . . . . . . . . . 303 19.4.5 pthread cleanup push - Establish Cancellation Handler

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 19.4.6 pthread cleanup pop - Remove Cancellation Handler … 305

20 Services Provided by C Library (libc) …. 307

20.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 20.2 Standard Utility Functions (stdlib.h) . . . . . . . . . . . . . . . . . . . . . . . . 307 20.3 Character Type Macros and Functions (ctype.h). . . . . . . . . . . . . 308 20.4 Input and Output (stdio.h) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 20.5 Strings and Memory (string.h) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 20.6 Signal Handling (signal.h). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 20.7 Time Functions (time.h) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 20.8 Locale (locale.h) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 20.9 Reentrant Versions of Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 20.10 Miscellaneous Macros and Functions . . . . . . . . . . . . . . . . . . . . . . . 312 20.11 Variable Argument Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 20.12 Reentrant System Calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313

21 Services Provided by the Math Library (libm)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

21.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 21.2 Standard Math Functions (math.h) . . . . . . . . . . . . . . . . . . . . . . . . . 315

22 Status of Implementation . . . . . . . . . . . . . . . . . . 317

Command and Variable Index . . . . . . . . . . . . . . . . . . 319

Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323

Preface

This is the User’s Guide for the POSIX API support provided in RTEMS.

The functionality described in this document is based on the following standards:

• POSIX 1003.1b-1993.

• POSIX 1003.1h/D3.

• Open Group Single UNIX Specification.

Much of the POSIX API standard is actually implemented in the Cygnus Newlib ANSI C Library. Please refer to documentation on Newlib for more information on the functionality it supplies.

This manual is still under construction and improvements are welcomed from users.

Acknowledgements

The Institute of Electrical and Electronics Engineers, Inc and The Open Group, have given us permission to reprint portions of their documentation.

Portions of this text are reprinted and reproduced in electronic form from IEEE Std 1003.1, 2004 Edition, Standard for Information Technology Operating System Interface (POSIX), The Open Group Base Specifications Issue 6, Copyright 2001-2004 by the Institute of Electrical and Electronics Engineers, Inc and The Open Group. In the event of any discrepancy between this version and the original IEEE and The Open Group Standard, the original IEEE and The Open Group Standard is the referee document. The original Standard can be obtained online at http://www.opengroup.org/unix/online.html

(http://www.opengroup.org/unix/online.html).

This notice shall appear on any product containing this material.

1 Process Creation and Execution Manager

1.1 Introduction

The process creation and execution manager provides the functionality associated with the creation and termination of processes.

The directives provided by the process creation and execution manager are:

• fork - Create a Process

• execl - Execute a File

• execv - Execute a File

• execle - Execute a File

• execve - Execute a File

• execlp - Execute a File

• execvp - Execute a File

• pthread_atfork - Register Fork Handlers • wait - Wait for Process Termination

• waitpid - Wait for Process Termination • _exit - Terminate a Process

1.2 Background

POSIX process functionality can not be completely supported by RTEMS. This is because RTEMS provides no memory protection and implements a single process, multi-threaded execution model. In this light, RTEMS provides none of the routines that are associated with the creation of new processes. However, since the entire RTEMS application (e.g. executable) is logically a single POSIX process, RTEMS is able to provide implementations of many operations on processes. The rule of thumb is that those routines provide a meaningful result. For example, getpid() returns the node number.

1.3 Operations

The only functionality method defined by this manager which is supported by RTEMS is the _exit service. The implementation of _exit shuts the application down and is equivalent

to invoking either exit or rtems_shutdown_executive.

1.4 Directives

This section details the process creation and execution manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

1.4.1 fork - Create a Process CALLING SEQUENCE:

     #include <sys/types.h>
     int fork( void );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

NOTES:

NONE

1.4.2 execl - Execute a File

CALLING SEQUENCE:

     int execl(
       const char *path,
       const char *arg,
       ...
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.3 execv - Execute a File

CALLING SEQUENCE:

     int execv(
       const char *path,
       char const *argv[],
       ...
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.4 execle - Execute a File

CALLING SEQUENCE:

     int execle(
       const char *path,
       const char *arg,
       ...
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.5 execve - Execute a File

CALLING SEQUENCE:

     int execve(
       const char *path,
       char *const argv[],
       char *const envp[]
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.6 execlp - Execute a File

CALLING SEQUENCE:

     int execlp(
       const char *file,
       const char *arg,
       ...
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.7 execvp - Execute a File

CALLING SEQUENCE:

     int execvp(
       const char *file,
       char *const argv[]
       ...
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.8 pthread atfork - Register Fork Handlers

CALLING SEQUENCE:

     #include <sys/types.h>
     int pthread_atfork(
       void (*prepare)(void),
       void (*parent)(void),
       void (*child)(void)
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.9 wait - Wait for Process Termination CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/wait.h>
     int wait(
       int *stat_loc
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

NOTES:

NONE

1.4.10 waitpid - Wait for Process Termination

CALLING SEQUENCE:

     int wait(
       pid_t  pid,
       int   *stat_loc,
       int    options
     );

STATUS CODES:

ENOSYS This routine is not supported by RTEMS.

DESCRIPTION:

This routine is not supported by RTEMS.

1.4.11 exit - Terminate a Process CALLING SEQUENCE:

     void _exit(
       int status
     );

STATUS CODES:

NONE

DESCRIPTION:

The _exit() function terminates the calling process.

NOTES:

In RTEMS, a process is equivalent to the entire application on a single processor. Invoking this service terminates the application.

2 Signal Manager

2.1 Introduction

The signal manager provides the functionality associated with the generation, delivery, and management of process-oriented signals. The directives provided by the signal manager are:

• sigaddset - Add a Signal to a Signal Set

• sigdelset - Delete a Signal from a Signal Set

• sigfillset - Fill a Signal Set

• sigismember - Is Signal a Member of a Signal Set

• sigemptyset - Empty a Signal Set

• sigaction - Examine and Change Signal Action

• pthread_kill - Send a Signal to a Thread

• sigprocmask - Examine and Change Process Blocked Signals • pthread_sigmask - Examine and Change Thread Blocked Signals • kill - Send a Signal to a Process

• sigpending - Examine Pending Signals

• sigsuspend - Wait for a Signal

• pause - Suspend Process Execution

• sigwait - Synchronously Accept a Signal

• sigwaitinfo - Synchronously Accept a Signal

• sigtimedwait - Synchronously Accept a Signal with Timeout • sigqueue - Queue a Signal to a Process

• alarm - Schedule Alarm

• ualarm - Schedule Alarm in Microseconds

2.2 Background

2.2.1 Signals

POSIX signals are an asynchronous event mechanism. Each process and thread has a set of signals associated with it. Individual signals may be enabled (e.g. unmasked) or blocked (e.g. ignored) on both a per-thread and process level. Signals which are enabled have a signal handler associated with them. When the signal is generated and conditions are met, then the signal handler is invoked in the proper process or thread context asynchronous relative to the logical thread of execution. If a signal has been blocked when it is generated, then it is queued and kept pending until the thread or process unblocks the signal or explicitly checks for it. Traditional, non-real-time POSIX signals do not queue. Thus if a process or thread has blocked a particular signal, then multiple occurrences of that signal are recorded as a single occurrence of that signal.

One can check for the set of outstanding signals that have been blocked. Services are provided to check for outstanding process or thread directed signals.

2.2.2 Signal Delivery

Signals which are directed at a thread are delivered to the specified thread.

Signals which are directed at a process are delivered to a thread which is selected based on the following algorithm:

1. If the action for this signal is currently SIG_IGN, then the signal is simply ignored.

2. If the currently executing thread has the signal unblocked, then the signal is delivered to it.

3. If any threads are currently blocked waiting for this signal (sigwait()), then the signal is delivered to the highest priority thread waiting for this signal.

4. If any other threads are willing to accept delivery of the signal, then the signal is delivered to the highest priority thread of this set. In the event, multiple threads of the same priority are willing to accept this signal, then priority is given first to ready threads, then to threads blocked on calls which may be interrupted, and finally to threads blocked on non-interruptible calls.

5. In the event the signal still can not be delivered, then it is left pending. The first

      thread to unblock the signal (sigprocmask() or pthread_sigprocmask()) or to

wait for this signal (sigwait()) will be the recipient of the signal.

2.3 Operations

2.3.1 Signal Set Management

Each process and each thread within that process has a set of individual signals and handlers associated with it. Services are provided to construct signal sets for the purposes of building signal sets – type sigset_t – that are used to provide arguments to the services that mask, unmask, and check on pending signals.

2.3.2 Blocking Until Signal Generation

A thread may block until receipt of a signal. The "sigwait" and "pause" families of services block until the requested signal is received or if using sigtimedwait() until the specified timeout period has elapsed.

2.3.3 Sending a Signal

This is accomplished via one of a number of services that sends a signal to either a process or thread. Signals may be directed at a process by the service kill() or at a thread by the

service pthread_kill()

2.4 Directives

This section details the signal manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

2.4.1 sigaddset - Add a Signal to a Signal Set CALLING SEQUENCE:

     #include <signal.h>
     int sigaddset(
       sigset_t *set,
       int       signo
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function adds the signo to the specified signal set.

NOTES:

NONE

2.4.2 sigdelset - Delete a Signal from a Signal Set CALLING SEQUENCE:

     #include <signal.h>
     int sigdelset(
       sigset_t *set,
       int       signo
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function deletes the signo to the specified signal set.

NOTES:

NONE

2.4.3 sigfillset - Fill a Signal Set CALLING SEQUENCE:

     #include <signal.h>
     int sigfillset(
       sigset_t *set
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function fills the specified signal set such that all signals are set.

NOTES:

NONE

2.4.4 sigismember - Is Signal a Member of a Signal Set CALLING SEQUENCE:

     #include <signal.h>
     int sigismember(
       const sigset_t *set,
       int             signo
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function returns returns 1 if signo is a member of set and 0 otherwise.

NOTES:

NONE

2.4.5 sigemptyset - Empty a Signal Set CALLING SEQUENCE:

     #include <signal.h>
     int sigemptyset(
       sigset_t *set
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function fills the specified signal set such that all signals are cleared.

NOTES:

NONE

2.4.6 sigaction - Examine and Change Signal Action

CALLING SEQUENCE:

     #include <signal.h>
     int sigaction(
       int                     sig,
       const struct sigaction *act,
       struct sigaction       *oact
     );

STATUS CODES:

EINVAL Invalid argument passed.

ENOTSUP Realtime Signals Extension option not supported.

DESCRIPTION:

This function is used to change the action taken by a process on receipt of the specfic signal sig. The new action is specified by act and the previous action is returned via oact.

NOTES:

The signal number cannot be SIGKILL.

2.4.7 pthread kill - Send a Signal to a Thread CALLING SEQUENCE:

     #include <signal.h>
     int pthread_kill(
       pthread_t thread,
       int       sig
     );

STATUS CODES:

ESRCH The thread indicated by the parameter thread is invalid. EINVAL Invalid argument passed.

DESCRIPTION:

This functions sends the specified signal sig to thread.

NOTES:

NONE

2.4.8 sigprocmask - Examine and Change Process Blocked Signals

CALLING SEQUENCE:

     #include <signal.h>
     int sigprocmask(
       int             how,
       const sigset_t *set,
       sigset_t       *oset
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function is used to alter the set of currently blocked signals on a process wide basis. A blocked signal will not be received by the process. The behavior of this function is dependent on the value of how which may be one of the following:

SIG_BLOCK The set of blocked signals is set to the union of set and those signals currently blocked.

SIG_UNBLOCK The signals specific in set are removed from the currently blocked set.

SIG_SETMASK The set of currently blocked signals is set to set.

If oset is not NULL, then the set of blocked signals prior to this call is returned in oset.

NOTES:

It is not an error to unblock a signal which is not blocked.

2.4.9 pthread sigmask - Examine and Change Thread Blocked Signals

CALLING SEQUENCE:

     #include <signal.h>
     int pthread_sigmask(
       int             how,
       const sigset_t *set,
       sigset_t       *oset
     );

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

This function is used to alter the set of currently blocked signals for the calling thread. A blocked signal will not be received by the process. The behavior of this function is dependent on the value of how which may be one of the following:

SIG_BLOCK The set of blocked signals is set to the union of set and those signals currently blocked.

SIG_UNBLOCK The signals specific in set are removed from the currently blocked set.

SIG_SETMASK The set of currently blocked signals is set to set.

If oset is not NULL, then the set of blocked signals prior to this call is returned in oset.

NOTES:

It is not an error to unblock a signal which is not blocked.

2.4.10 kill - Send a Signal to a Process

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <signal.h>
     int kill(
       pid_t pid,
       int   sig
     );

STATUS CODES:

EINVAL Invalid argument passed.

EPERM Process does not have permission to send the signal to any receiving process.

ESRCH The process indicated by the parameter pid is invalid.

DESCRIPTION:

This function sends the signal sig to the process pid.

2.4.11 sigpending - Examine Pending Signals CALLING SEQUENCE:

     #include <signal.h>
     int sigpending(
       const sigset_t *set
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EFAULT Invalid address for set.

DESCRIPTION:

This function allows the caller to examine the set of currently pending signals. A pending signal is one which has been raised but is currently blocked. The set of pending signals is returned in set.

NOTES:

NONE

2.4.12 sigsuspend - Wait for a Signal CALLING SEQUENCE:

     #include <signal.h>
     int sigsuspend(
       const sigset_t *sigmask
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINTR Signal interrupted this function.

DESCRIPTION:

This function temporarily replaces the signal mask for the process with that specified by sigmask and blocks the calling thread until the signal is raised.

NOTES:

NONE

2.4.13 pause - Suspend Process Execution CALLING SEQUENCE:

     #include <signal.h>
     int pause( void );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINTR Signal interrupted this function.

DESCRIPTION:

This function causes the calling thread to be blocked until an unblocked signal is received.

NOTES:

NONE

2.4.14 sigwait - Synchronously Accept a Signal

CALLING SEQUENCE:

     #include <signal.h>
     int sigwait(
       const sigset_t *set,
       int            *sig
     );

STATUS CODES:

EINVAL Invalid argument passed.

EINTR Signal interrupted this function.

DESCRIPTION:

This function selects a pending signal based on the set specified in set, atomically clears it from the set of pending signals, and returns the signal number for that signal in sig.

2.4.15 sigwaitinfo - Synchronously Accept a Signal

CALLING SEQUENCE:

     #include <signal.h>
     int sigwaitinfo(
       const sigset_t *set,
       siginfo_t      *info
     );

STATUS CODES:

EINTR Signal interrupted this function.

DESCRIPTION:

This function selects a pending signal based on the set specified in set, atomically clears it from the set of pending signals, and returns information about that signal in info.

2.4.16 sigtimedwait - Synchronously Accept a Signal with Timeout

CALLING SEQUENCE:

     #include <signal.h>
     int sigtimedwait(
       const sigset_t        *set,
       siginfo_t             *info,
       const struct timespec *timeout
     );

STATUS CODES:

EAGAIN Timed out while waiting for the specified signal set.

EINVAL Nanoseconds field of the timeout argument is invalid.

EINTR Signal interrupted this function.

DESCRIPTION:

This function selects a pending signal based on the set specified in set, atomically clears it from the set of pending signals, and returns information about that signal in info. The calling thread will block up to timeout waiting for the signal to arrive.

NOTES:

If timeout is NULL, then the calling thread will wait forever for the specified signal set.

2.4.17 sigqueue - Queue a Signal to a Process

CALLING SEQUENCE:

     #include <signal.h>
     int sigqueue(
       pid_t              pid,
       int                signo,
       const union sigval value
     );

STATUS CODES:

EAGAIN No resources available to queue the signal. The process has already queued SIGQUEUE MAX signals that are still pending at the receiver or the systemwide resource limit has been exceeded.

EINVAL The value of the signo argument is an invalid or unsupported signal number.

EPERM The process does not have the appropriate privilege to send the signal to the receiving process.

ESRCH The process pid does not exist.

DESCRIPTION:

This function sends the signal specified by signo to the process pid

2.4.18 alarm - Schedule Alarm

CALLING SEQUENCE:

     #include <unistd.h>
     unsigned int alarm(
       unsigned int seconds
     );

STATUS CODES:

This call always succeeds.

If there was a previous alarm() request with time remaining, then this routine returns the number of seconds until that outstanding alarm would have fired. If no previous alarm() request was outstanding, then zero is returned.

DESCRIPTION:

The alarm() service causes the SIGALRM signal to be generated after the number of seconds specified by seconds has elapsed.

NOTES:

Alarm requests do not queue. If alarm is called while a previous request is outstanding, the call will result in rescheduling the time at which the SIGALRM signal will be generated.

If the notification signal, SIGALRM, is not caught or ignored, the calling process is terminated.

2.4.19 ualarm - Schedule Alarm in Microseconds

CALLING SEQUENCE:

     #include <unistd.h>
     useconds_t ualarm(
       useconds_t useconds,
       useconds_t interval
     );

STATUS CODES:

This call always succeeds.

If there was a previous ualarm() request with time remaining, then this routine returns the number of seconds until that outstanding alarm would have fired. If no previous alarm() request was outstanding, then zero is returned.

DESCRIPTION:

The ualarm() service causes the SIGALRM signal to be generated after the number of microseconds specified by useconds has elapsed.

When interval is non-zero, repeated timeout notification occurs with a period in microseconds specified by interval.

NOTES:

Alarm requests do not queue. If alarm is called while a previous request is outstanding, the call will result in rescheduling the time at which the SIGALRM signal will be generated.

If the notification signal, SIGALRM, is not caught or ignored, the calling process is terminated.

3 Process Environment Manager

3.1 Introduction

The process environment manager is responsible for providing the functions related to user and group Id management. The directives provided by the process environment manager are:

• getpid - Get Process ID

• getppid - Get Parent Process ID

• getuid - Get User ID

• geteuid - Get Effective User ID

• getgid - Get Real Group ID

• getegid - Get Effective Group ID

• setuid - Set User ID

• setgid - Set Group ID

• getgroups - Get Supplementary Group IDs

• getlogin - Get User Name

• getlogin_r - Reentrant Get User Name

• getpgrp - Get Process Group ID

• setsid - Create Session and Set Process Group ID

• setpgid - Set Process Group ID for Job Control

• uname - Get System Name

• times - Get Process Times

• getenv - Get Environment Variables

• setenv - Set Environment Variables

• ctermid - Generate Terminal Pathname

• ttyname - Determine Terminal Device Name

• ttyname_r - Reentrant Determine Terminal Device Name • isatty - Determine if File Descriptor is Terminal

• sysconf - Get Configurable System Variables

3.2 Background

3.2.1 Users and Groups

RTEMS provides a single process, multi-threaded execution environment. In this light, the notion of user and group is somewhat without meaning. But RTEMS does provide services to provide a synthetic version of user and group. By default, a single user and group is associated with the application. Thus unless special actions are taken, every thread in the application shares the same user and group Id. The initial rationale for providing user and group Id functionality in RTEMS was for the filesystem infrastructure to implement file permission checks. The effective user/group Id capability has since been used to implement permissions checking by the ftpd server.

In addition to the "real" user and group Ids, a process may have an effective user/group Id. This allows a process to function using a more limited permission set for certain operations.

3.2.2 User and Group Names

POSIX considers user and group Ids to be a unique integer that may be associated with a name. This is usually accomplished via a file named /etc/passwd for user Id mapping and /etc/groups for group Id mapping. Again, although RTEMS is effectively a single process and thus single user system, it provides limited support for user and group names. When configured with an appropriate filesystem, RTEMS will access the appropriate files to map user and group Ids to names.

If these files do not exist, then RTEMS will synthesize a minimal version so this family of services return without error. It is important to remember that a design goal of the RTEMS POSIX services is to provide useable and meaningful results even though a full process model is not available.

3.2.3 Environment Variables

POSIX allows for variables in the run-time environment. These are name/value pairs that make be dynamically set and obtained by programs. In a full POSIX environment with command line shell and multiple processes, environment variables may be set in one process – such as the shell – and inherited by child processes. In RTEMS, there is only one process and thus only one set of environment variables across all processes.

3.3 Operations

3.3.1 Accessing User and Group Ids

The user Id associated with the current thread may be obtain using the getuid() service. Similarly, the group Id may be obtained using the getgid() service.

3.3.2 Accessing Environment Variables

The value associated with an environment variable may be obtained using the getenv() service and set using the putenv() service.

3.4 Directives

This section details the process environment manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

3.4.1 getpid - Get Process ID CALLING SEQUENCE:

     int getpid( void );

STATUS CODES:

The process Id is returned.

DESCRIPTION:

This service returns the process Id.

NOTES:

NONE

3.4.2 getppid - Get Parent Process ID CALLING SEQUENCE:

     int getppid( void );

STATUS CODES:

The parent process Id is returned.

DESCRIPTION:

This service returns the parent process Id.

NOTES:

NONE

3.4.3 getuid - Get User ID CALLING SEQUENCE:

     int getuid( void );

STATUS CODES:

The effective user Id is returned.

DESCRIPTION:

This service returns the effective user Id.

NOTES:

NONE

3.4.4 geteuid - Get Effective User ID CALLING SEQUENCE:

     int geteuid( void );

STATUS CODES:

The effective group Id is returned.

DESCRIPTION:

This service returns the effective group Id.

NOTES:

NONE

3.4.5 getgid - Get Real Group ID CALLING SEQUENCE:

     int getgid( void );

STATUS CODES:

The group Id is returned.

DESCRIPTION:

This service returns the group Id.

NOTES:

NONE

3.4.6 getegid - Get Effective Group ID CALLING SEQUENCE:

     int getegid( void );

STATUS CODES:

The effective group Id is returned.

DESCRIPTION:

This service returns the effective group Id.

NOTES:

NONE

3.4.7 setuid - Set User ID CALLING SEQUENCE:

     int setuid(
       uid_t uid
     );

STATUS CODES:

This service returns 0.

DESCRIPTION:

This service sets the user Id to uid.

NOTES:

NONE

3.4.8 setgid - Set Group ID CALLING SEQUENCE:

     int setgid(
       gid_t  gid
     );

STATUS CODES:

This service returns 0.

DESCRIPTION:

This service sets the group Id to gid.

NOTES:

NONE

3.4.9 getgroups - Get Supplementary Group IDs CALLING SEQUENCE:

     int getgroups(
       int    gidsetsize,
       gid_t  grouplist[]
     );

STATUS CODES:

NA

DESCRIPTION:

This service is not implemented as RTEMS has no notion of supplemental groups.

NOTES:

If supported, this routine would only be allowed for the super-user.

3.4.10 getlogin - Get User Name CALLING SEQUENCE:

     char *getlogin( void );

STATUS CODES:

Returns a pointer to a string containing the name of the current user.

DESCRIPTION:

This routine returns the name of the current user.

NOTES:

This routine is not reentrant and subsequent calls to getlogin() will overwrite the same buffer.

3.4.11 getlogin r - Reentrant Get User Name

CALLING SEQUENCE:

     int getlogin_r(
       char   *name,
       size_t  namesize
     );

STATUS CODES:

EINVAL The arguments were invalid.

DESCRIPTION:

This is a reentrant version of the getlogin() service. The caller specified their own buffer, name, as well as the length of this buffer, namesize.

3.4.12 getpgrp - Get Process Group ID CALLING SEQUENCE:

     pid_t getpgrp( void );

STATUS CODES:

The procress group Id is returned.

DESCRIPTION:

This service returns the current progress group Id.

NOTES:

This routine is implemented in a somewhat meaningful way for RTEMS but is truly not functional.

3.4.13 setsid - Create Session and Set Process Group ID CALLING SEQUENCE:

     pid_t setsid( void );

STATUS CODES:

EPERM The application does not have permission to create a process group.

DESCRIPTION:

This routine always returns EPERM as RTEMS has no way to create new processes and thus no way to create a new process group.

NOTES:

NONE

3.4.14 setpgid - Set Process Group ID for Job Control CALLING SEQUENCE:

     int setpgid(
       pid_t pid,
       pid_t pgid
     );

STATUS CODES:

ENOSYS The routine is not implemented.

DESCRIPTION:

This service is not implemented for RTEMS as process groups are not supported.

NOTES:

NONE

3.4.15 uname - Get System Name

CALLING SEQUENCE:

     int uname(
       struct utsname *name
     );

STATUS CODES:

EPERM The provided structure pointer is invalid.

DESCRIPTION:

This service returns system information to the caller. It does this by filling in the struct utsname format structure for the caller.

NOTES:

The information provided includes the operating system (RTEMS in all configurations), the node number, the release as the RTEMS version, and the CPU family and model. The CPU model name will indicate the multilib executive variant being used.

3.4.16 times - Get process times

CALLING SEQUENCE:

     #include <sys/time.h>
     clock_t times(
       struct tms *ptms
     );

STATUS CODES:

This routine returns the number of clock ticks that have elapsed since the system was initialized (e.g. the application was started).

DESCRIPTION:

times stores the current process times in ptms. The format of struct tms is as defined in <sys/times.h>. RTEMS fills in the field tms_utime with the number of ticks that the calling thread has executed and the field tms_stime with the number of clock ticks since system boot (also returned). All other fields in the ptms are left zero.

NOTES:

RTEMS has no way to distinguish between user and system time so this routine returns the most meaningful information possible.

3.4.17 getenv - Get Environment Variables CALLING SEQUENCE:

     char *getenv(
       const char *name
     );

STATUS CODES:

NULL when no match pointer to value when successful

DESCRIPTION:

This service searches the set of environment variables for a string that matches the specified name. If found, it returns the associated value.

NOTES:

The environment list consists of name value pairs that are of the form name = value.

3.4.18 setenv - Set Environment Variables

CALLING SEQUENCE:

     int setenv(
       const char *name,
       const char *value,
       int overwrite
     );

STATUS CODES:

Returns 0 if successful and -1 otherwise.

DESCRIPTION:

This service adds the variable name to the environment with value. If name is not already exist, then it is created. If name exists and overwrite is zero, then the previous value is not overwritten.

3.4.19 ctermid - Generate Terminal Pathname

CALLING SEQUENCE:

     char *ctermid(
       char *s
     );

STATUS CODES:

Returns a pointer to a string indicating the pathname for the controlling terminal.

DESCRIPTION:

This service returns the name of the terminal device associated with this process. If s is NULL, then a pointer to a static buffer is returned. Otherwise, s is assumed to have a buffer of sufficient size to contain the name of the controlling terminal.

NOTES:

By default on RTEMS systems, the controlling terminal is /dev/console. Again this implementation is of limited meaning, but it provides true and useful results which should be sufficient to ease porting applications from a full POSIX implementation to the reduced profile supported by RTEMS.

3.4.20 ttyname - Determine Terminal Device Name CALLING SEQUENCE:

     char *ttyname(
       int fd
     );

STATUS CODES:

Pointer to a string containing the terminal device name or NULL is returned on any error.

DESCRIPTION:

This service returns a pointer to the pathname of the terminal device that is open on the file descriptor fd. If fd is not a valid descriptor for a terminal device, then NULL is returned.

NOTES:

This routine uses a static buffer.

3.4.21 ttyname r - Reentrant Determine Terminal Device Name CALLING SEQUENCE:

     int ttyname_r(
       int   fd,
       char *name,
       int   namesize
     );

STATUS CODES:

This routine returns -1 and sets errno as follows: EBADF If not a valid descriptor for a terminal device. EINVAL If name is NULL or namesize are insufficient.

DESCRIPTION:

This service the pathname of the terminal device that is open on the file descriptor fd.

NOTES:

NONE

3.4.22 isatty - Determine if File Descriptor is Terminal CALLING SEQUENCE:

     int isatty(
       int fd
     );

STATUS CODES:

Returns 1 if fd is a terminal device and 0 otherwise.

DESCRIPTION:

This service returns 1 if fd is an open file descriptor connected to a terminal and 0 otherwise.

NOTES:

3.4.23 sysconf - Get Configurable System Variables

CALLING SEQUENCE:

     long sysconf(
       int name
     );

STATUS CODES:

The value returned is the actual value of the system resource. If the requested configuration name is a feature flag, then 1 is returned if the available and 0 if it is not. On any other error condition, -1 is returned.

DESCRIPTION:

This service is the mechanism by which an application determines values for system limits or options at runtime.

NOTES:

Much of the information that may be obtained via sysconf has equivalent macros in <unistd.h. However, those macros reflect conservative limits which may have been altered by application configuration.

4 Files and Directories Manager

4.1 Introduction

The files and directories manager is ... The directives provided by the files and directories manager are:

• opendir - Open a Directory

• readdir - Reads a directory

• rewinddir - Resets the readdir() pointer

• scandir - Scan a directory for matching entries

• telldir - Return current location in directory stream

• closedir - Ends directory read operation

• getdents - Get directory entries

• chdir - Changes the current working directory

• fchdir - Changes the current working directory

• getcwd - Gets current working directory

• open - Opens a file

• creat - Create a new file or rewrite an existing one

• umask - Sets a file creation mask

• link - Creates a link to a file

• symlink - Creates a symbolic link to a file

• readlink - Obtain the name of the link destination

• mkdir - Makes a directory

• mkfifo - Makes a FIFO special file

• unlink - Removes a directory entry

• rmdir - Delete a directory

• rename - Renames a file

• stat - Gets information about a file.

• fstat - Gets file status

• lstat - Gets file status

• access - Check permissions for a file.

• chmod - Changes file mode

• fchmod - Changes permissions of a file

• chown - Changes the owner and/ or group of a file

• utime - Change access and/or modification times of an inode • ftruncate - Truncate a file to a specified length

• truncate - Truncate a file to a specified length

• pathconf - Gets configuration values for files

• fpathconf - Get configuration values for files

• mknod - Create a directory

4.2 Background

4.2.1 Path Name Evaluation

A pathname is a string that consists of no more than PATH_MAX bytes, including the terminating null character. A pathname has an optional beginning slash, followed by zero or more filenames separated by slashes. If the pathname refers to a directory, it may also have one or more trailing slashes. Multiple successive slahes are considered to be the same as one slash.

POSIX allows a pathname that begins with precisely two successive slashes to be interpreted in an implementation-defined manner. RTEMS does not currently recognize this as a special condition. Any number of successive slashes is treated the same as a single slash. POSIX requires that an implementation treat more than two leading slashes as a single slash.

4.3 Operations

There is currently no text in this section.

4.4 Directives

This section details the files and directories manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

4.4.1 opendir - Open a Directory

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <dirent.h>
     int opendir(
       const char *dirname
     );

STATUS CODES:

EACCES Search permission was denied on a component of the path prefix of dirname, or read permission is denied

EMFILE Too many file descriptors in use by process

ENFILE Too many files are currently open in the system.

ENOENT Directory does not exist, or name is an empty string.

ENOMEM Insufficient memory to complete the operation.

ENOTDIR name is not a directory.

DESCRIPTION:

This routine opens a directory stream corresponding to the directory specified by the dirname argument. The directory stream is positioned at the first entry.

NOTES:

The routine is implemented in Cygnus newlib.

4.4.2 readdir - Reads a directory

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <dirent.h>
     int readdir(
       DIR *dirp
     );

STATUS CODES:

EBADF Invalid file descriptor

DESCRIPTION:

The readdir() function returns a pointer to a structure dirent representing the next directory entry from the directory stream pointed to by dirp. On end-of-file, NULL is returned.

The readdir() function may (or may not) return entries for . or .. Your program should tolerate reading dot and dot-dot but not require them.

The data pointed to be readdir() may be overwritten by another call to readdir() for the same directory stream. It will not be overwritten by a call for another directory.

NOTES:

If ptr is not a pointer returned by malloc(), calloc(), or realloc() or has been deallocated with free() or realloc(), the results are not portable and are probably disastrous.

The routine is implemented in Cygnus newlib.

4.4.3 rewinddir - Resets the readdir() pointer CALLING SEQUENCE:

     #include <sys/types.h>
     #include <dirent.h>
     void rewinddir(
       DIR *dirp
     );

STATUS CODES:

No value is returned.

DESCRIPTION:

The rewinddir() function resets the position associated with the directory stream pointed to by dirp. It also causes the directory stream to refer to the current state of the directory.

NOTES:

NONE If dirp is not a pointer by opendir(), the results are undefined. The routine is implemented in Cygnus newlib.

4.4.4 scandir - Scan a directory for matching entries

CALLING SEQUENCE:

     #include <dirent.h>
     int scandir(
       const char       *dir,
       struct dirent ***namelist,
       int  (*select)(const struct dirent *),
       int  (*compar)(const struct dirent **, const struct dirent **)
     );

STATUS CODES:

ENOMEM Insufficient memory to complete the operation.

DESCRIPTION:

The scandir() function scans the directory dir, calling select() on each directory entry. Entries for which select() returns non-zero are stored in strings allocated via malloc(), sorted using qsort() with the comparison function compar(), and collected in array namelist which is allocated via malloc(). If select is NULL, all entries are selected.

NOTES:

The routine is implemented in Cygnus newlib.

4.4.5 telldir - Return current location in directory stream CALLING SEQUENCE:

     #include <dirent.h>
     off_t telldir(
       DIR *dir
     );

STATUS CODES:

EBADF Invalid directory stream descriptor dir.

DESCRIPTION:

The telldir() function returns the current location associated with the directory stream

dir.

NOTES:

The routine is implemented in Cygnus newlib.

4.4.6 closedir - Ends directory read operation

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <dirent.h>
     int closedir(
       DIR *dirp
     );

STATUS CODES:

EBADF Invalid file descriptor

DESCRIPTION:

The directory stream associated with dirp is closed. The value in dirp may not be usable after a call to closedir().

NOTES:

NONE

The argument to closedir() must be a pointer returned by opendir(). If it is not, the results are not portable and most likely unpleasant.

The routine is implemented in Cygnus newlib.

4.4.7 chdir - Changes the current working directory

CALLING SEQUENCE:

     #include <unistd.h>
     int chdir(
       const char *path
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following:

EACCES Search permission is denied for a directory in a file’s path prefix.

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified pathname was not a directory when directory was expected.

DESCRIPTION:

The chdir() function causes the directory named by path to become the current working directory; that is, the starting point for searches of pathnames not beginning with a slash.

If chdir() detects an error, the current working directory is not changed.

4.4.8 fchdir - Changes the current working directory

CALLING SEQUENCE:

     #include <unistd.h>
     int fchdir(
       int fd
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following:

EACCES Search permission is denied for a directory in a file’s path prefix.

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified pathname was not a directory when directory was expected.

DESCRIPTION:

The fchdir() function causes the directory named by fd to become the current working directory; that is, the starting point for searches of pathnames not beginning with a slash.

If fchdir() detects an error, the current working directory is not changed.

4.4.9 getcwd - Gets current working directory CALLING SEQUENCE:

     #include <unistd.h>
     int getcwd( void );

STATUS CODES:

EINVAL Invalid argument ERANGE Result is too large EACCES Search permission is denied for a directory in a file’s path prefix.

DESCRIPTION:

The getcwd() function copies the absolute pathname of the current working directory to the character array pointed to by buf. The size argument is the number of bytes available

in buf

NOTES:

There is no way to determine the maximum string length that fetcwd() may need to return. Applications should tolerate getting ERANGE and allocate a larger buffer. It is possible for getcwd() to return EACCES if, say, login puts the process into a directory without read access. The 1988 standard uses int instead of size_t for the second parameter.

4.4.10 open - Opens a file

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     #include <fcntl.h>
     int open(
       const char *path,
       int         oflag,
       mode_t      mode
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix.

EEXIST The named file already exists.

EINTR Function was interrupted by a signal.

EISDIR Attempt to open a directory for writing or to rename a file to be a directory.

EMFILE Too many file descriptors are in use by this process.

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENFILE Too many files are currently open in the system.

ENOENT A file or directory does not exist.

ENOSPC No space left on disk.

ENOTDIR A component of the specified pathname was not a directory when a directory was expected.

ENXIO No such device. This error may also occur when a device is not ready, for example, a tape drive is off-line.

EROFS Read-only file system.

DESCRIPTION:

The open function establishes a connection between a file and a file descriptor. The file descriptor is a small integer that is used by I/O functions to reference the file. The path argument points to the pathname for the file.

The oflag argument is the bitwise inclusive OR of the values of symbolic constants. The programmer must specify exactly one of the following three symbols:

O RDONLY Open for reading only.

O WRONLY Open for writing only.

O RDWR Open for reading and writing.

Any combination of the following symbols may also be used.

O APPEND Set the file offset to the end-of-file prior to each write.

O CREAT If the file does not exist, allow it to be created. This flag indicates that the mode argument is present in the call to open.

O EXCL This flag may be used only if O CREAT is also set. It causes the call to open to fail if the file already exists.

O NOCTTY If path identifies a terminal, this flag prevents that teminal from becoming the controlling terminal for thi9s process. See Chapter 8 for a description of terminal I/O.

O NONBLOCK Do no wait for the device or file to be ready or available. After the file is open, the read and write calls return immediately. If the process would be delayed in the read or write opermation, -1 is returned and errno is set to EAGAIN instead of blocking the caller.

O TRUNC This flag should be used only on ordinary files opened for writing. It causes the file to be tuncated to zero length..

Upon successful completion, open returns a non-negative file descriptor.

4.4.11 creat - Create a new file or rewrite an existing one CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     #include <fcntl.h>
     int creat(
       const char *path,
       mode_t      mode
     );

STATUS CODES:

EEXIST path already exists and O CREAT and O EXCL were used. EISDIR path refers to a directory and the access requested involved writing ETXTBSY path refers to an executable image which is currently being executed

and write access was requested

EFAULT path points outside your accessible address space EACCES The requested access to the file is not allowed, or one of the directories

in path did not allow search (execute) permission.

ENAMETOOLONG path was too long. ENOENT A directory component in path does not exist or is a dangling sym-

bolic link.

ENOTDIR A component used as a directory in path is not, in fact, a directory. EMFILE The process alreadyh has the maximum number of files open. ENFILE The limit on the total number of files open on the system has been

reached.

ENOMEM Insufficient kernel memory was available. EROFS path refers to a file on a read-only filesystem and write access was

requested

DESCRIPTION:

creat attempts to create a file and return a file descriptor for use in read, write, etc.

NOTES:

NONE The routine is implemented in Cygnus newlib.

4.4.12 umask - Sets a file creation mask.

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     mode_t umask(
       mode_t cmask
     );

STATUS CODES:

DESCRIPTION:

The umask() function sets the process file creation mask to cmask. The file creation mask is used during open(), creat(), mkdir(), mkfifo() calls to turn off permission bits in the mode argument. Bit positions that are set in cmask are cleared in the mode of the created file.

NOTES:

NONE

The cmask argument should have only permission bits set. All other bits should be zero.

In a system which supports multiple processes, the file creation mask is inherited across fork() and exec() calls. This makes it possible to alter the default permission bits of created files. RTEMS does not support multiple processes so this behavior is not possible.

4.4.13 link - Creates a link to a file CALLING SEQUENCE:

     #include <unistd.h>
     int link(
       const char *existing,
       const char *new
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix EEXIST The named file already exists. EMLINK The number of links would exceed LINK_MAX. ENAMETOOLONG Length of a filename string exceeds PATH MAX and

POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist. ENOSPC No space left on disk. ENOTDIR A component of the specified pathname was not a directory when a

directory was expected.

EPERM Operation is not permitted. Process does not have the appropriate

priviledges or permissions to perform the requested operations.

EROFS Read-only file system. EXDEV Attempt to link a file to another file system.

DESCRIPTION:

The link() function atomically creates a new link for an existing file and increments the link count for the file. If the link() function fails, no directories are modified. The existing argument should not be a directory. The caller may (or may not) need permission to access the existing file.

NOTES:

NONE

4.4.14 symlink - Creates a symbolic link to a file CALLING SEQUENCE:

     #include <unistd.h>
     int symlink(
       const char *topath,
       const char *frompath
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix EEXIST The named file already exists. ENAMETOOLONG Length of a filename string exceeds PATH MAX and

POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist. ENOSPC No space left on disk. ENOTDIR A component of the specified pathname was not a directory when a

directory was expected.

EPERM Operation is not permitted. Process does not have the appropriate

priviledges or permissions to perform the requested operations.

EROFS Read-only file system.

DESCRIPTION:

The symlink() function creates a symbolic link from the frombath to the topath. The symbolic link will be interpreted at run-time. If the symlink() function fails, no directories are modified. The caller may (or may not) need permission to access the existing file.

NOTES:

NONE

4.4.15 readlink - Obtain the name of a symbolic link destination

CALLING SEQUENCE:

     #include <unistd.h>
     int readlink(
       const char *path,
       char       *buf,
       size_t      bufsize
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the prefix pathname was not a directory when a directory was expected.

ELOOP Too many symbolic links were encountered in the pathname.

EINVAL The pathname does not refer to a symbolic link

EFAULT An invalid pointer was passed into the readlink() routine.

DESCRIPTION:

The readlink() function places the symbolic link destination into buf argument and returns the number of characters copied.

If the symbolic link destination is longer than bufsize characters the name will be truncated.

4.4.16 mkdir - Makes a directory

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int mkdir(
       const char *path,
       mode_t      mode
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

EEXIST The name file already exist.

EMLINK The number of links would exceed LINK MAX

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOSPC No space left on disk.

ENOTDIR A component of the specified pathname was not a directory when a directory was expected.

EROFS Read-only file system.

DESCRIPTION:

The mkdir() function creates a new diectory named path. The permission bits (modified by the file creation mask) are set from mode. The owner and group IDs for the directory are set from the effective user ID and group ID.

The new directory may (or may not) contain entries for.. and .. but is otherwise empty.

4.4.17 mkfifo - Makes a FIFO special file

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>

     int mkfifo(
       const char *path,
       mode_t      mode
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

EEXIST The named file already exists.

ENOENT A file or directory does not exist.

ENOSPC No space left on disk.

ENOTDIR A component of the specified path was not a directory when a directory was expected.

EROFS Read-only file system.

DESCRIPTION:

The mkfifo() function creates a new FIFO special file named path. The permission bits (modified by the file creation mask) are set from mode. The owner and group IDs for the FIFO are set from the efective user ID and group ID.

4.4.18 unlink - Removes a directory entry

CALLING SEQUENCE:

     #include <unistd.h>
     int unlink(
       const char path
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

EBUSY The directory is in use.

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified path was not a directory when a directory was expected.

EPERM Operation is not permitted. Process does not have the appropriate priviledges or permissions to perform the requested operations.

EROFS Read-only file system.

DESCRIPTION:

The unlink function removes the link named by path and decrements the link count of the file referenced by the link. When the link count goes to zero and no process has the file open, the space occupied by the file is freed and the file is no longer accessible.

4.4.19 rmdir - Delete a directory CALLING SEQUENCE:

     #include <unistd.h>
     int rmdir(
       const char *pathname
     );

STATUS CODES:

EPERM The filesystem containing pathname does not support the removal of

directories.

EFAULT pathname points ouside your accessible address space. EACCES Write access to the directory containing pathname was not allowed

for the process’s effective uid, or one of the directories in pathname did not allow search (execute) permission.

EPERM The directory containing pathname has the stickybit (S ISVTX) set and the process’s effective uid is neither the uid of the file to be delected nor that of the director containing it.

ENAMETOOLONG pathname was too long. ENOENT A dirctory component in pathname does not exist or is a dangling

symbolic link.

ENOTDIR pathname, or a component used as a directory in pathname, is not,

in fact, a directory.

ENOTEMPTY pathname contains entries other than . and .. . EBUSY pathname is the current working directory or root directory of some

process

EBUSY pathname is the current directory or root directory of some process. ENOMEM Insufficient kernel memory was available EROGS pathname refers to a file on a read-only filesystem. ELOOP pathname contains a reference to a circular symbolic link

DESCRIPTION:

rmdir deletes a directory, which must be empty

NOTES:

NONE

4.4.20 rename - Renames a file CALLING SEQUENCE:

     #include <unistd.h>
     int rename(
       const char *old,
       const char *new
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix. EBUSY The directory is in use. EEXIST The named file already exists. EINVAL Invalid argument. EISDIR Attempt to open a directory for writing or to rename a file to be a

directory.

EMLINK The number of links would exceed LINK MAX. ENAMETOOLONG Length of a filename string exceeds PATH MAX and

POSIX NO TRUNC is in effect.

ENOENT A file or directory does no exist. ENOSPC No space left on disk. ENOTDIR A component of the specified pathname was not a directory when a

directory was expected.

ENOTEMPTY Attempt to delete or rename a non-empty directory. EROFS Read-only file system EXDEV Attempt to link a file to another file system.

DESCRIPTION:

The rename() function causes the file known bo old to now be known as new. Ordinary files may be renamed to ordinary files, and directories may be renamed to directories; however, files cannot be converted using rename(). The new pathname may not contain a path prefix of old.

NOTES:

If a file already exists by the name new, it is removed. The rename() function is atomic. If the rename() detects an error, no files are removed. This guarantees that the rename("x", "x") does not remove x. You may not rename dot or dot-dot. The routine is implemented in Cygnus newlib using link() and unlink().

4.4.21 stat - Gets information about a file

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int stat(
       const char  *path,
       struct stat *buf
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix.

EBADF Invalid file descriptor.

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified pathname was not a directory when a directory was expected.

DESCRIPTION:

The path argument points to a pathname for a file. Read, write, or execute permission for the file is not required, but all directories listed in path must be searchable. The stat() function obtains information about the named file and writes it to the area pointed to by

buf.

4.4.22 fstat - Gets file status

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int fstat(
       int          fildes,
       struct stat *buf
     );

STATUS CODES:

EBADF Invalid file descriptor

DESCRIPTION:

The fstat() function obtains information about the file associated with fildes and writes it to the area pointed to by the buf argument.

NOTES:

If the filesystem object referred to by fildes is a link, then the information returned in buf refers to the destination of that link. This is in contrast to lstat() which does not follow the link.

4.4.23 lstat - Gets file status

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int lstat(
       int          fildes,
       struct stat *buf
     );

STATUS CODES:

EBADF Invalid file descriptor

DESCRIPTION:

The lstat() function obtains information about the file associated with fildes and writes it to the area pointed to by the buf argument.

NOTES:

If the filesystem object referred to by fildes is a link, then the information returned in buf refers to the link itself. This is in contrast to fstat() which follows the link.

The lstat() routine is defined by BSD 4.3 and SVR4 and not included in POSIX 1003.1b-

4.4.24 access - Check permissions for a file

CALLING SEQUENCE:

     #include <unistd.h>
     int access(
       const char *pathname,
       int         mode
     );

STATUS CODES:

EACCES The requested access would be denied, either to the file itself or one of the directories in pathname.

EFAULT pathname points outside your accessible address space.

EINVAL Mode was incorrectly specified.

ENAMETOOLONG pathname is too long.

ENOENT A directory component in pathname would have been accessible but does not exist or was a dangling symbolic link.

ENOTDIR A component used as a directory in pathname is not, in fact, a directory.

ENOMEM Insufficient kernel memory was available.

DESCRIPTION:

Access checks whether the process would be allowed to read, write or test for existence of the file (or other file system object) whose name is pathname. If pathname is a symbolic link permissions of the file referred by this symbolic link are tested.

Mode is a mask consisting of one or more of R OK, W OK, X OK and F OK.

4.4.25 chmod - Changes file mode.

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int chmod(
       const char *path,
       mode_t      mode
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified pathname was not a directory when a directory was expected.

EPERM Operation is not permitted. Process does not have the appropriate priviledges or permissions to perform the requested operations.

EROFS Read-only file system.

DESCRIPTION:

Set the file permission bits, the set user ID bit, and the set group ID bit for the file named by path to mode. If the effective user ID does not match the owner of the file and the calling process does not have the appropriate privileges, chmod() returns -1 and sets errno

to EPERM.

4.4.26 fchmod - Changes permissions of a file CALLING SEQUENCE:

     #include <sys/types.h>
     #include <sys/stat.h>
     int fchmod(
       int    fildes,
       mode_t mode
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix. EBADF The descriptor is not valid. EFAULT path points outside your accessible address space. EIO A low-level I/o error occurred while modifying the inode. ELOOP path contains a circular reference ENAMETOOLONG Length of a filename string exceeds PATH MAX and

POSIX NO TRUNC is in effect.

ENOENT A file or directory does no exist. ENOMEM Insufficient kernel memory was avaliable. ENOTDIR A component of the specified pathname was not a directory when a

directory was expected.

EPERM The effective UID does not match the owner of the file, and is not

zero

EROFS Read-only file system

DESCRIPTION:

The mode of the file given by path or referenced by filedes is changed.

NOTES:

NONE

4.4.27 getdents - Get directory entries

CALLING SEQUENCE:

     #include <unistd.h>
     #include <linux/dirent.h>
     #include <linux/unistd.h>
     long getdents(
       int   dd_fd,
       char *dd_buf,
       int   dd_len
     );

STATUS CODES:

A successful call to getdents returns th the number of bytes read. On end of directory, 0 is returned. When an error occurs, -1 is returned, and errno is set appropriately.

EBADF Invalid file descriptor fd.

EFAULT Argument points outside the calling process’s address space.

EINVAL Result buffer is too small.

ENOENT No such directory.

ENOTDIR File descriptor does not refer to a directory.

DESCRIPTION:

getdents reads several dirent structures from the directory pointed by fd into the memory area pointed to by dirp. The parameter count is the size of the memory area.

4.4.28 chown - Changes the owner and/or group of a file.

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <unistd.h>
     int chown(
       const char *path,
       uid_t       owner,
       gid_t       group
     );

STATUS CODES:

EACCES Search permission is denied for a directory in a file’s path prefix

EINVAL Invalid argument

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist.

ENOTDIR A component of the specified pathname was not a directory when a directory was expected.

EPERM Operation is not permitted. Process does not have the appropriate priviledges or permissions to perform the requested operations.

EROFS Read-only file system.

DESCRIPTION:

The user ID and group ID of the file named by path are set to owner and path, respectively.

For regular files, the set group ID (S ISGID) and set user ID (S ISUID) bits are cleared.

Some systems consider it a security violation to allow the owner of a file to be changed, If users are billed for disk space usage, loaning a file to another user could result in incorrect billing. The chown() function may be restricted to privileged users for some or all files. The group ID can still be changed to one of the supplementary group IDs.

NOTES:

This function may be restricted for some file. The pathconf function can be used to test

the _PC_CHOWN_RESTRICTED flag.

4.4.29 utime - Change access and/or modification times of an inode

CALLING SEQUENCE:

     #include <sys/types.h>
     int utime(
       const char     *filename,
       struct utimbuf *buf
     );

STATUS CODES:

EACCES Permission to write the file is denied

ENOENT Filename does not exist

DESCRIPTION:

Utime changes the access and modification times of the inode specified by filename to the actime and modtime fields of buf respectively. If buf is NULL, then the access and modification times of the file are set to the current time.

4.4.30 ftruncate - truncate a file to a specified length CALLING SEQUENCE:

     #include <unistd.h>
     int ftrunctate(
       int    fd,
       size_t length
     );

STATUS CODES:

ENOTDIR A component of the path prefix is not a directory. EINVAL The pathname contains a character with the high-order bit set. ENAMETOOLONG A component of a pathname exceeded 255 characters, or an entire

path name exceeded 1023 characters.

ENOENT The named file does not exist. EACCES The named file is not writable by the user. EACCES Search permission is denied for a component of the path prefix. ELOOP Too many symbolic links were encountered in translating the path-

name

EISDIR The named file is a directory. EROFS The named file resides on a read-only file system ETXTBSY The file is a pure procedure (shared text) file that is being executed EIO An I/O error occurred updating the inode. EFAULT Path points outside the process’s allocated address space. EBADF The fd is not a valid descriptor.

DESCRIPTION:

truncate() causes the file named by path or referenced by fd to be truncated to at most length bytes in size. If the file previously was larger than this size, the extra data is lost. With ftruncate(), the file must be open for writing.

NOTES:

NONE

4.4.31 truncate - truncate a file to a specified length CALLING SEQUENCE:

     #include <unistd.h>
     int trunctate(
       const char *path,
       size_t      length
     );

STATUS CODES:

ENOTDIR A component of the path prefix is not a directory. EINVAL The pathname contains a character with the high-order bit set. ENAMETOOLONG A component of a pathname exceeded 255 characters, or an entire

path name exceeded 1023 characters.

ENOENT The named file does not exist. EACCES The named file is not writable by the user. EACCES Search permission is denied for a component of the path prefix. ELOOP Too many symbolic links were encountered in translating the path-

name

EISDIR The named file is a directory. EROFS The named file resides on a read-only file system ETXTBSY The file is a pure procedure (shared text) file that is being executed EIO An I/O error occurred updating the inode. EFAULT Path points outside the process’s allocated address space. EBADF The fd is not a valid descriptor.

DESCRIPTION:

truncate() causes the file named by path or referenced by fd to be truncated to at most length bytes in size. If the file previously was larger than this size, the extra data is lost. With ftruncate(), the file must be open for writing.

NOTES:

NONE

4.4.32 pathconf - Gets configuration values for files

CALLING SEQUENCE:

#include <unistd.h>
int pathconf(
const char *path,
int         name
);

STATUS CODES:

EINVAL Invalid argument

EACCES Permission to write the file is denied

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist

ENOTDIR A component of the specified path was not a directory whan a directory was expected.

DESCRIPTION:

pathconf() gets a value for the configuration option name for the open file descriptor

filedes.

The possible values for name are:

PC LINK MAX returns the maximum number of links to the file. If filedes or path refer to a directory, then the value applies to the whole directory. The corresponding macro is _POSIX_LINK_MAX.

PC MAX CANON returns the maximum length of a formatted input line, where filedes or path must refer to a terminal. The corresponding macro

is _POSIX_MAX_CANON.

PC MAX INPUT returns the maximum length of an input line, where filedes or path must refer to a terminal. The corresponding macro is _POSIX_MAX_

INPUT.

PC NAME MAX returns the maximum length of a filename in the directory path or filedes. The process is allowed to create. The corresponding macro

is _POSIX_NAME_MAX.

PC PATH MAX returns the maximum length of a relative pathname when path or filedes is the current working directory. The corresponding macro

is _POSIX_PATH_MAX.

PC PIPE BUF returns the size of the pipe buffer, where filedes must refer to a pipe or FIFO and path must refer to a FIFO. The corresponding

macro is _POSIX_PIPE_BUF.

PC CHOWN RESTRICTED returns nonzero if the chown(2) call may not be used on this file. If filedes or path refer to a directory, then this applies to all files in that directory. The corresponding macro is _POSIX_CHOWN_

RESTRICTED.

NOTES:

Files with name lengths longer than the value returned for name equal _PC_NAME_MAX may exist in the given directory.

4.4.33 fpathconf - Gets configuration values for files

CALLING SEQUENCE:

#include <unistd.h>
int fpathconf(
int filedes,
int name
);

STATUS CODES:

EINVAL Invalid argument

EACCES Permission to write the file is denied

ENAMETOOLONG Length of a filename string exceeds PATH MAX and POSIX NO TRUNC is in effect.

ENOENT A file or directory does not exist

ENOTDIR A component of the specified path was not a directory whan a directory was expected.

DESCRIPTION:

pathconf() gets a value for the configuration option name for the open file descriptor

filedes.

The possible values for name are:

PC LINK MAX returns the maximum number of links to the file. If filedes or path refer to a directory, then the value applies to the whole directory. The corresponding macro is POSIX LINK MAX.

PC MAX CANON returns the maximum length of a formatted input line, where filedes or path must refer to a terminal. The corresponding macro

is _POSIX_MAX_CANON.

PC MAX INPUT returns the maximum length of an input line, where filedes or path must refer to a terminal. The corresponding macro is _POSIX_MAX_

INPUT.

PC NAME MAX returns the maximum length of a filename in the directory path or filedes. The process is allowed to create. The corresponding macro

is _POSIX_NAME_MAX.

PC PATH MAX returns the maximum length of a relative pathname when path or filedes is the current working directory. The corresponding macro

is _POSIX_PATH_MAX.

PC PIPE BUF returns the size of the pipe buffer, where filedes must refer to a pipe or FIFO and path must refer to a FIFO. The corresponding

macro is _POSIX_PIPE_BUF.

PC CHOWN RESTRICTED returns nonzero if the chown() call may not be used on this file. If filedes or path refer to a directory, then this applies to all files in that directory. The corresponding macro is _POSIX_CHOWN_

RESTRICTED.

4.4.34 mknod - create a directory

CALLING SEQUENCE:

     #include <unistd.h>
     #include <fcntl.h>
     #include <sys/types.h>
     #include <sys/stat.h>
     long mknod(
       const char *pathname,
       mode_t      mode,
       dev_t       dev
     );

STATUS CODES:

mknod returns zero on success, or -1 if an error occurred (in which case, errno is set appropriately).

ENAMETOOLONG pathname was too long.

ENOENT A directory component in pathname does not exist or is a dangling symbolic link.

ENOTDIR A component used in the directory pathname is not, in fact, a directory.

ENOMEM Insufficient kernel memory was available

EROFS pathname refers to a file on a read-only filesystem.

ELOOP pathname contains a reference to a circular symbolic link, ie a symbolic link whose expansion contains a reference to itself.

ENOSPC The device containing pathname has no room for the new node.

DESCRIPTION:

mknod attempts to create a filesystem node (file, device special file or named pipe) named

pathname, specified by mode and dev.

mode specifies both the permissions to use and the type of node to be created.

It should be a combination (using bitwise OR) of one of the file types listed below and the permissions for the new node.

The permissions are modified by the process’s umask in the usual way: the permissions of the created node are (mode & ~umask).

The file type should be one of S_IFREG, S_IFCHR, S_IFBLK and S_IFIFO to specify a normal file (which will be created empty), character special file, block special file or FIFO (named pipe), respectively, or zero, which will create a normal file.

If the file type is S_IFCHR or S_IFBLK then dev specifies the major and minor numbers of the newly created device special file; otherwise it is ignored.

The newly created node will be owned by the effective uid of the process. If the directory containing the node has the set group id bit set, or if the filesystem is mounted with BSD group semantics, the new node will inherit the group ownership from its parent directory; otherwise it will be owned by the effective gid of the process.

5 Input and Output Primitives Manager

5.1 Introduction

The input and output primitives manager is ... The directives provided by the input and output primitives manager are:

• pipe - Create an Inter-Process Channel

• dup - Duplicates an open file descriptor

• dup2 - Duplicates an open file descriptor

• close - Closes a file

• read - Reads from a file

• write - Writes to a file

• fcntl - Manipulates an open file descriptor

• lseek - Reposition read/write file offset

• fsync - Synchronize file complete in-core state with that on disk • fdatasync - Synchronize file in-core data with that on disk

• sync - Schedule file system updates

• mount - Mount a file system

• unmount - Unmount file systems

• readv - Vectored read from a file

• writev - Vectored write to a file

• aio_read - Asynchronous Read

• aio_write - Asynchronous Write

• lio_listio - List Directed I/O

• aio_error - Retrieve Error Status of Asynchronous I/O Operation • aio_return - Retrieve Return Status Asynchronous I/O Operation • aio_cancel - Cancel Asynchronous I/O Request

• aio_suspend - Wait for Asynchronous I/O Request

• aio_fsync - Asynchronous File Synchronization

5.2 Background

There is currently no text in this section.

5.3 Operations

There is currently no text in this section.

5.4 Directives

This section details the input and output primitives manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

5.4.1 pipe - Create an Inter-Process Channel CALLING SEQUENCE:

     int pipe(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.2 dup - Duplicates an open file descriptor

CALLING SEQUENCE:

     #include <unistd.h>
     int dup(
       int fildes
     );

STATUS CODES:

EBADF Invalid file descriptor.

EINTR Function was interrupted by a signal.

EMFILE The process already has the maximum number of file descriptors open and tried to open a new one.

DESCRIPTION:

The dup function returns the lowest numbered available file descriptor. This new desciptor refers to the same open file as the original descriptor and shares any locks.

5.4.3 dup2 - Duplicates an open file descriptor

CALLING SEQUENCE:

     #include <unistd.h>
     int dup2(
       int fildes,
       int fildes2
     );

STATUS CODES:

EBADF Invalid file descriptor.

EINTR Function was interrupted by a signal.

EMFILE The process already has the maximum number of file descriptors open and tried to open a new one.

DESCRIPTION:

dup2 creates a copy of the file descriptor oldfd.

The old and new descriptors may be used interchangeably. They share locks, file position pointers and flags; for example, if the file position is modified by using lseek on one of the descriptors, the position is also changed for the other.

5.4.4 close - Closes a file

CALLING SEQUENCE:

     #include <unistd.h>
     int close(
       int fildes
     );

STATUS CODES:

EBADF Invalid file descriptor

EINTR Function was interrupted by a signal.

DESCRIPTION:

The close() function deallocates the file descriptor named by fildes and makes it available for reuse. All outstanding record locks owned by this process for the file are unlocked.

NOTES:

A signal can interrupt the close() function. In that case, close() returns -1 with errno set to EINTR. The file may or may not be closed.

5.4.5 read - Reads from a file

CALLING SEQUENCE:

     #include <unistd.h>
     int read(
       int           fildes,
       void         *buf,
       unsigned int  nbyte
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following:

EAGAIN The O NONBLOCK flag is set for a file descriptor and the process would be delayed in the I/O operation.

EBADF Invalid file descriptor

EINTR Function was interrupted by a signal.

EIO Input or output error

EINVAL Bad buffer pointer

DESCRIPTION:

The read() function reads nbyte bytes from the file associated with fildes into the buffer pointed to by buf.

The read() function returns the number of bytes actually read and placed in the buffer. This will be less than nbyte if:

• The number of bytes left in the file is less than nbyte.

• The read() request was interrupted by a signal.

• The file is a pipe or FIFO or special file with less than nbytes immediately available for reading.

When attempting to read from any empty pipe or FIFO:

• If no process has the pipe open for writing, zero is returned to indicate end-of-file.

• If some process has the pipe open for writing and O NONBLOCK is set, -1 is returned and errno is set to EAGAIN.

• If some process has the pipe open for writing and O NONBLOCK is clear, read() waits for some data to be written or the pipe to be closed.

When attempting to read from a file other than a pipe or FIFO and no data is available.

• If O NONBLOCK is set, -1 is returned and errno is set to EAGAIN.

• If O NONBLOCK is clear, read() waits for some data to become available.

• The O NONBLOCK flag is ignored if data is available.

Chapter 5: Input and Output Primitives Manager 109

NOTES:^{NONE} 5.4.6 write - Writes to a file

CALLING SEQUENCE:

     #include <unistd.h>
     int write(
       int           fildes,
       const void   *buf,
       unsigned int  nbytes
     );

STATUS CODES:

EAGAIN The O NONBLOCK flag is set for a file descriptor and the process would be delayed in the I/O operation.

EBADF Invalid file descriptor

EFBIG An attempt was made to write to a file that exceeds the maximum file size

EINTR The function was interrupted by a signal.

EIO Input or output error.

ENOSPC No space left on disk.

EPIPE Attempt to write to a pope or FIFO with no reader.

EINVAL Bad buffer pointer

DESCRIPTION:

The write() function writes nbyte from the array pointed to by buf into the file associated

with fildes.

If nybte is zero and the file is a regular file, the write() function returns zero and has no other effect. If nbyte is zero and the file is a special file, te results are not portable.

The write() function returns the number of bytes written. This number will be less than nbytes if there is an error. It will never be greater than nbytes.

5.4.7 fcntl - Manipulates an open file descriptor

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <fcntl.h>
     #include <unistd.h>
     int fcntl(
       int fildes,
       int cmd
     );

STATUS CODES:

EACCESS Search permission is denied for a direcotry in a file’s path prefix.

EAGAIN The O NONBLOCK flag is set for a file descriptor and the process would be delayed in the I/O operation.

EBADF Invalid file descriptor

EDEADLK An fcntl with function F SETLKW would cause a deadlock.

EINTR The functioin was interrupted by a signal.

EINVAL Invalid argument

EMFILE Too many file descriptor or in use by the process.

ENOLCK No locks available

DESCRIPTION:

fcntl() performs one of various miscellaneous operations on fd. The operation in question is determined by cmd:

F DUPFD Makes arg be a copy of fd, closing fd first if necessary. The same functionality can be more easily achieved by using dup2(). The old and new descriptors may be used interchangeably. They share locks, file position pointers and flags; for example, if the file position is modified by using lseek() on one of the descriptors, the position is also changed for the other. The two descriptors do not share the close-on-exec flag, however. The close-on-exec flag of the copy is off, meaning that it will be closed on exec. On success, the new descriptor is returned.

F GETFD Read the close-on-exec flag. If the low-order bit is 0, the file will remain open across exec, otherwise it will be closed.

F SETFD Set the close-on-exec flag to the value specified by arg (only the least significant bit is used).

F GETFL Read the descriptor’s flags (all flags (as set by open()) are returned).

F SETFL Set the descriptor’s flags to the value specified by arg. Only O_

APPEND and O_NONBLOCK may be set.

The flags are shared between copies (made with dup() etc.) of the same file descriptor. The flags and their semantics are described in open().

F GETLK, F SETLK and F SETLKW Manage discretionary file locks. The third argument arg is a pointer to a struct flock (that may be overwritten by this call).

F GETLK Return the flock structure that prevents us from obtaining the lock, or set the l_type field of the lock to F_UNLCK if there is no obstruction.

F SETLK The lock is set (when l_type is F_RDLCK or F_WRLCK) or cleared (when it is F_UNLCK. If lock is held by someone else, this call returns -1 and sets errno to EACCES or EAGAIN.

F SETLKW Like F_SETLK, but instead of returning an error we wait for the lock to be released.

F GETOWN Get the process ID (or process group) of the owner of a socket. Process groups are returned as negative values.

F SETOWN Set the process or process group that owns a socket.

For these commands, ownership means receiving SIGIO or SIGURG signals. Process groups are specified using negative values.

NOTES:

The errors returned by dup2 are different from those returned by F_DUPFD.

5.4.8 lseek - Reposition read/write file offset

CALLING SEQUENCE:

     #include <sys/types.h>
     #include <unistd.h>
     int lseek(
       int    fildes,
       off_t  offset,
       int    whence
     );

STATUS CODES:

EBADF fildes is not an open file descriptor.

ESPIPE fildes is associated with a pipe, socket or FIFO.

EINVAL whence is not a proper value.

DESCRIPTION:

The lseek function repositions the offset of the file descriptor fildes to the argument offset according to the directive whence. The argument fildes must be an open file descriptor. Lseek repositions the file pointer fildes as follows:

• If whence is SEEK SET, the offset is set to offset bytes.

• If whence is SEEK CUR, the offset is set to its current location plus offset bytes.

• If whence is SEEK END, the offset is set to the size of the file plus offset bytes.

The lseek function allows the file offset to be set beyond the end of the existing end-of-file of the file. If data is later written at this point, subsequent reads of the data in the gap return bytes of zeros (until data is actually written into the gap).

Some devices are incapable of seeking. The value of the pointer associated with such a device is undefined.

5.4.9 fsync - Synchronize file complete in-core state with that on

disk CALLING SEQUENCE:

     int fsync(
       int fd
     );

STATUS CODES:

On success, zero is returned. On error, -1 is returned, and errno is set appropriately. EBADF fd is not a valid descriptor open for writing EINVAL fd is bound to a special file which does not support support synchro-

nization

EROFS fd is bound to a special file which does not support support synchro-

nization

EIO An error occurred during synchronization

DESCRIPTION:

fsync copies all in-core parts of a file to disk.

NOTES:

NONE

5.4.10 fdatasync - Synchronize file in-core data with that on disk

CALLING SEQUENCE:

     int fdatasync(
       int fd
     );

STATUS CODES:

On success, zero is returned. On error, -1 is returned, and errno is set appropriately.

EBADF fd is not a valid file descriptor open for writing.

EINVAL fd is bound to a special file which does not support synchronization.

EIO An error occurred during synchronization.

EROFS fd is bound to a special file which dows not support synchronization.

DESCRIPTION:

fdatasync flushes all data buffers of a file to disk (before the system call returns). It resembles fsync but is not required to update the metadata such as access time.

Applications that access databases or log files often write a tiny data fragment (e.g., one line in a log file) and then call fsync immediately in order to ensure that the written data is physically stored on the harddisk. Unfortunately, fsync will always initiate two write operations: one for the newly written data and another one in order to update the modification time stored in the inode. If the modification time is not a part of the transaction concept fdatasync can be used to avoid unnecessary inode disk write operations.

5.4.11 sync - Schedule file system updates CALLING SEQUENCE:

     void sync(void);

STATUS CODES:

NONE

DESCRIPTION:

The sync service causes all information in memory that updates file systems to be scheduled for writing out to all file systems.

NOTES:

The writing of data to the file systems is only guaranteed to be scheduled upon return. It is not necessarily complete upon return from sync.

5.4.12 mount - Mount a file system

CALLING SEQUENCE:

     #include <libio.h>
     int mount(
       rtems_filesystem_mount_table_entry_t **mt_entry,
       rtems_filesystem_operations_table    *fs_ops,
       rtems_filesystem_options_t            fsoptions,
       char                                 *device,
       char                                 *mount_point
     );

STATUS CODES:

EXXX

DESCRIPTION:

The mount routines mounts the filesystem class which uses the filesystem operations specified by fs_ops and fsoptions. The filesystem is mounted at the directory mount_point and the mode of the mounted filesystem is specified by fsoptions. If this filesystem class requires a device, then the name of the device must be specified by device.

If this operation succeeds, the mount table entry for the mounted filesystem is returned in

mt_entry.

5.4.13 unmount - Unmount file systems CALLING SEQUENCE:

     #include <libio.h>
     int unmount(
       const char *mount_path
     );

STATUS CODES:

EXXX

DESCRIPTION:

The unmount routine removes the attachment of the filesystem specified by mount_path.

NOTES:

NONE

5.4.14 readv - Vectored read from a file

CALLING SEQUENCE:

     #include <sys/uio.h>
     ssize_t readv(
       int                 fd,
       const struct iovec *iov,
       int                 iovcnt
     );

STATUS CODES:

In addition to the errors detected by Input and Output Primitives Manager read - Reads

from a file, read(), this routine may return -1 and sets errno based upon the following errors:

EINVAL The sum of the iov_len values in the iov array overflowed an ssize_

                   t.

EINVAL The iovcnt argument was less than or equal to 0, or greater than

                   IOV_MAX.

DESCRIPTION:

The readv() function is equivalent to read() except as described here. The readv() function shall place the input data into the iovcnt buffers specified by the members of the

iov array: iov[0], iov[1], ..., iov[iovcnt-1].

Each iovec entry specifies the base address and length of an area in memory where data should be placed. The readv() function always fills an area completely before proceeding to the next.

5.4.15 writev - Vectored write to a file

CALLING SEQUENCE:

     #include <sys/uio.h>
     ssize_t writev(
       int                 fd,
       const struct iovec *iov,
       int                 iovcnt
     );

STATUS CODES:

In addition to the errors detected by Input and Output Primitives Manager write -

Write to a file, write(), this routine may return -1 and sets errno based upon the following errors:

EINVAL The sum of the iov_len values in the iov array overflowed an ssize_

                   t.

EINVAL The iovcnt argument was less than or equal to 0, or greater than

                   IOV_MAX.

DESCRIPTION:

The writev() function is equivalent to write(), except as noted here. The writev() function gathers output data from the iovcnt buffers specified by the members of the iov

array: iov[0], iov[1], ..., iov[iovcnt-1]. The iovcnt argument is valid if greater

than 0 and less than or equal to IOV_MAX.

Each iovec entry specifies the base address and length of an area in memory from which data should be written. The writev() function always writes a complete area before proceeding to the next.

If fd refers to a regular file and all of the iov_len members in the array pointed to by iov are 0, writev() returns 0 and has no other effect. For other file types, the behavior is unspecified by POSIX.

5.4.16 aio read - Asynchronous Read CALLING SEQUENCE:

     int aio_read(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.17 aio write - Asynchronous Write CALLING SEQUENCE:

     int aio_write(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.18 lio listio - List Directed I/O CALLING SEQUENCE:

     int lio_listio(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.19 aio error - Retrieve Error Status of Asynchronous I/O

Operation CALLING SEQUENCE:

     int aio_error(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.20 aio return - Retrieve Return Status Asynchronous I/O

Operation CALLING SEQUENCE:

     int aio_return(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.21 aio cancel - Cancel Asynchronous I/O Request CALLING SEQUENCE:

     int aio_cancel(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.22 aio suspend - Wait for Asynchronous I/O Request CALLING SEQUENCE:

     int aio_suspend(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

5.4.23 aio fsync - Asynchronous File Synchronization CALLING SEQUENCE:

     int aio_fsync(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

6 Device- and Class- Specific Functions Manager

6.1 Introduction

The device- and class- specific functions manager is ... The directives provided by the device- and class- specific functions manager are:

• cfgetispeed - Reads terminal input baud rate

• cfgetospeed - Reads terminal output baud rate

• cfsetispeed - Sets terminal input baud rate

• cfsetospeed - Set terminal output baud rate

• tcgetattr - Gets terminal attributes

• tcsetattr - Set terminal attributes

• tcsendbreak - Sends a break to a terminal

• tcdrain - Waits for all output to be transmitted to the terminal • tcflush - Discards terminal data

• tcflow - Suspends/restarts terminal output

• tcgetpgrp - Gets foreground process group ID

• tcsetpgrp - Sets foreground process group ID

6.2 Background

There is currently no text in this section.

6.3 Operations

There is currently no text in this section.

6.4 Directives

This section details the device- and class- specific functions manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

6.4.1 cfgetispeed - Reads terminal input baud rate

CALLING SEQUENCE:

     #include <termios.h>
     int cfgetispeed(
       const struct termios *p
     );

STATUS CODES:

The cfgetispeed() function returns a code for baud rate.

DESCRIPTION:

The cfsetispeed() function stores a code for the terminal speed stored in a struct termios. The codes are defined in <termios.h> by the macros BO, B50, B75, B110, B134, B150, B200, B300, B600, B1200, B1800, B2400, B4800, B9600, B19200, and B38400.

The cfsetispeed() function does not do anything to the hardware. It merely stores a value for use by tcsetattr().

NOTES:

Baud rates are defined by symbols, such as B110, B1200, B2400. The actual number returned for any given speed may change from system to system.

6.4.2 cfgetospeed - Reads terminal output baud rate

CALLING SEQUENCE:

     #include <termios.h>
     int cfgetospeed(
       const struct termios *p
     );

STATUS CODES:

The cfgetospeed() function returns the termios code for the baud rate.

DESCRIPTION:

The cfgetospeed() function returns a code for the terminal speed stored in a struct termios. The codes are defined in <termios.h> by the macros BO, B50, B75, B110, B134, B150, B200, B300, B600, B1200, B1800, B2400, B4800, B9600, B19200, and B38400.

The cfgetospeed() function does not do anything to the hardware. It merely returns the value stored by a previous call to tcgetattr().

NOTES:

Baud rates are defined by symbols, such as B110, B1200, B2400. The actual number returned for any given speed may change from system to system.

6.4.3 cfsetispeed - Sets terminal input baud rate

CALLING SEQUENCE:

     #include <termios.h>
     int cfsetispeed(
       struct termios *p,
       speed_t         speed
     );

STATUS CODES:

The cfsetispeed() function returns a zero when successful and returns -1 when an error occurs.

DESCRIPTION:

The cfsetispeed() function stores a code for the terminal speed stored in a struct termios. The codes are defined in <termios.h> by the macros B0, B50, B75, B110, B134, B150, B200, B300, B600, B1200, B1800, B2400, B4800, B9600, B19200, and B38400.

NOTES:

This function merely stores a value in the termios structure. It does not change the terminal speed until a tcsetattr() is done. It does not detect impossible terminal speeds.

6.4.4 cfsetospeed - Sets terminal output baud rate

CALLING SEQUENCE:

     #include <termios.h>
     int cfsetospeed(
       struct termios *p,
       speed_t         speed
     );

STATUS CODES:

The cfsetospeed() function returns a zero when successful and returns -1 when an error occurs.

DESCRIPTION:

The cfsetospeed() function stores a code for the terminal speed stored in a struct termios. The codes are defiined in <termios.h> by the macros B0, B50, B75, B110, B134, B150, B200, B300, B600, B1200, B1800, B2400, B4800, B9600, B19200, and B38400.

The cfsetospeed() function does not do anything to the hardware. It merely stores a value for use by tcsetattr().

NOTES:

This function merely stores a value in the termios structure. It does not change the terminal speed until a tcsetattr() is done. It does not detect impossible terminal speeds.

6.4.5 tcgetattr - Gets terminal attributes

CALLING SEQUENCE:

     #include <termios.h>
     #include <unistd.h>
     int tcgetattr(
       int             fildes,
       struct termios *p
     );

STATUS CODES:

EBADF Invalid file descriptor

ENOOTY Terminal control function attempted for a file that is not a terminal.

DESCRIPTION:

The tcgetattr() gets the parameters associated with the terminal referred to by fildes and stores them into the termios() structure pointed to by termios_p.

6.4.6 tcsetattr - Set terminal attributes

CALLING SEQUENCE:

#include <termios.h>
#include <unistd.h>
int tcsetattr(
  int                   fildes,
  int                   options,
  const struct termios *tp
);

STATUS CODES:

E The

DESCRIPTION:

6.4.7 tcsendbreak - Sends a break to a terminal CALLING SEQUENCE:

     int tcsendbreak(
       int fd
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

6.4.8 tcdrain - Waits for all output to be transmitted to the

terminal. CALLING SEQUENCE:

     #include <termios.h>
     #include <unistd.h>
     int tcdrain(
       int fildes
     );

STATUS CODES:

EBADF Invalid file descriptor EINTR Function was interrupted by a signal ENOTTY Terminal control function attempted for a file that is not a terminal.

DESCRIPTION:

The tcdrain() function waits until all output written to fildes has been transmitted.

NOTES:

NONE

6.4.9 tcflush - Discards terminal data CALLING SEQUENCE:

     int tcflush(
       int fd
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

6.4.10 tcflow - Suspends/restarts terminal output. CALLING SEQUENCE:

     int tcflow(
       int fd
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

6.4.11 tcgetpgrp - Gets foreground process group ID CALLING SEQUENCE:

     int tcgetpgrp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

6.4.12 tcsetpgrp - Sets foreground process group ID CALLING SEQUENCE:

     int tcsetpgrp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

This routine is not currently supported by RTEMS but could be in a future version.

7 Language-Specific Services for the C Programming Language Manager

7.1 Introduction

The language-specific services for the C programming language manager is ... The directives provided by the language-specific services for the C programming language manager are:

• setlocale - Set the Current Locale

• fileno - Obtain File Descriptor Number for this File

• fdopen - Associate Stream with File Descriptor

• flockfile - Acquire Ownership of File Stream

• ftrylockfile - Poll to Acquire Ownership of File Stream

• funlockfile - Release Ownership of File Stream

• getc_unlocked - Get Character without Locking

• getchar_unlocked - Get Character from stdin without Locking

• putc_unlocked - Put Character without Locking

• putchar_unlocked - Put Character to stdin without Locking

• setjmp - Save Context for Non-Local Goto

• longjmp - Non-Local Jump to a Saved Context

• sigsetjmp - Save Context with Signal Status for Non-Local Goto • siglongjmp - Non-Local Jump with Signal Status to a Saved Context • tzset - Initialize Time Conversion Information

• strtok_r - Reentrant Extract Token from String

• asctime_r - Reentrant struct tm to ASCII Time Conversion

• ctime_r - Reentrant time t to ASCII Time Conversion

• gmtime_r - Reentrant UTC Time Conversion

• localtime_r - Reentrant Local Time Conversion

• rand_r - Reentrant Random Number Generation

7.2 Background

There is currently no text in this section.

7.3 Operations

There is currently no text in this section.

7.4 Directives

This section details the language-specific services for the C programming language manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

7.4.1 setlocale - Set the Current Locale CALLING SEQUENCE:

     int setlocale(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.2 fileno - Obtain File Descriptor Number for this File CALLING SEQUENCE:

     int fileno(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.3 fdopen - Associate Stream with File Descriptor CALLING SEQUENCE:

     int fdopen(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.4 flockfile - Acquire Ownership of File Stream CALLING SEQUENCE:

     int flockfile(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.5 ftrylockfile - Poll to Acquire Ownership of File Stream CALLING SEQUENCE:

     int ftrylockfile(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.6 funlockfile - Release Ownership of File Stream CALLING SEQUENCE:

     int funlockfile(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.7 getc unlocked - Get Character without Locking CALLING SEQUENCE:

     int getc_unlocked(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.8 getchar unlocked - Get Character from stdin without

Locking CALLING SEQUENCE:

     int getchar_unlocked(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.9 putc unlocked - Put Character without Locking CALLING SEQUENCE:

     int putc_unlocked(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.10 putchar unlocked - Put Character to stdin without Locking CALLING SEQUENCE:

     int putchar_unlocked(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.11 setjmp - Save Context for Non-Local Goto CALLING SEQUENCE:

     int setjmp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.12 longjmp - Non-Local Jump to a Saved Context CALLING SEQUENCE:

     int longjmp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.13 sigsetjmp - Save Context with Signal Status for Non-Local

Goto CALLING SEQUENCE:

     int sigsetjmp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.14 siglongjmp - Non-Local Jump with Signal Status to a

Saved Context CALLING SEQUENCE:

     int siglongjmp(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.15 tzset - Initialize Time Conversion Information CALLING SEQUENCE:

     int tzset(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.16 strtok r - Reentrant Extract Token from String CALLING SEQUENCE:

     int strtok_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.17 asctime r - Reentrant struct tm to ASCII Time Conversion CALLING SEQUENCE:

     int asctime_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.18 ctime r - Reentrant time t to ASCII Time Conversion CALLING SEQUENCE:

     int ctime_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.19 gmtime r - Reentrant UTC Time Conversion CALLING SEQUENCE:

     int gmtime_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.20 localtime r - Reentrant Local Time Conversion CALLING SEQUENCE:

     int localtime_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

7.4.21 rand r - Reentrant Random Number Generation CALLING SEQUENCE:

     int rand_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8 System Databases Manager

8.1 Introduction

The system databases manager is ... The directives provided by the system databases manager are:

• getgrgid - Get Group File Entry for ID

• getgrgid_r - Reentrant Get Group File Entry

• getgrnam - Get Group File Entry for Name

• getgrnam_r - Reentrant Get Group File Entry for Name • getpwuid - Get Password File Entry for UID

• getpwuid_r - Reentrant Get Password File Entry for UID • getpwnam - Get Password File Entry for Name

• getpwnam_r - Reentrant Get Password File Entry for Name

8.2 Background

There is currently no text in this section.

8.3 Operations

There is currently no text in this section.

8.4 Directives

This section details the system databases manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

8.4.1 getgrgid - Get Group File Entry for ID CALLING SEQUENCE:

     int getgrgid(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.2 getgrgid r - Reentrant Get Group File Entry CALLING SEQUENCE:

     int getgrgid_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.3 getgrnam - Get Group File Entry for Name CALLING SEQUENCE:

     int getgrnam(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.4 getgrnam r - Reentrant Get Group File Entry for Name CALLING SEQUENCE:

     int getgrnam_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.5 getpwuid - Get Password File Entry for UID CALLING SEQUENCE:

     int getpwuid(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.6 getpwuid r - Reentrant Get Password File Entry for UID CALLING SEQUENCE:

     int getpwuid_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.7 getpwnam - Password File Entry for Name CALLING SEQUENCE:

     int getpwnam(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

8.4.8 getpwnam r - Reentrant Get Password File Entry for Name CALLING SEQUENCE:

     int getpwnam_r(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

9 Semaphore Manager

9.1 Introduction

The semaphore manager provides functions to allocate, delete, and control semaphores. This manager is based on the POSIX 1003.1 standard. The directives provided by the semaphore manager are:

• sem_init - Initialize an unnamed semaphore

• sem_destroy - Destroy an unnamed semaphore

• sem_open - Open a named semaphore

• sem_close - Close a named semaphore

• sem_unlink - Remove a named semaphore

• sem_wait - Lock a semaphore

• sem_trywait - Lock a semaphore

• sem_timedwait - Wait on a Semaphore for a Specified Time • sem_post - Unlock a semaphore

• sem_getvalue - Get the value of a semeaphore

9.2 Background

9.2.1 Theory

Semaphores are used for synchronization and mutual exclusion by indicating the availability and number of resources. The task (the task which is returning resources) notifying other tasks of an event increases the number of resources held by the semaphore by one. The task (the task which will obtain resources) waiting for the event decreases the number of resources held by the semaphore by one. If the number of resources held by a semaphore is insufficient (namely 0), the task requiring resources will wait until the next time resources are returned to the semaphore. If there is more than one task waiting for a semaphore, the tasks will be placed in the queue.

9.2.2 "sem t" Structure

The sem_t structure is used to represent semaphores. It is passed as an argument to the semaphore directives and is defined as follows:

     typedef int sem_t;

9.2.3 Building a Semaphore Attribute Set

9.3 Operations

9.3.1 Using as a Binary Semaphore

Although POSIX supports mutexes, they are only visible between threads. To work between processes, a binary semaphore must be used.

Creating a semaphore with a limit on the count of 1 effectively restricts the semaphore to being a binary semaphore. When the binary semaphore is available, the count is 1. When the binary semaphore is unavailable, the count is 0.

Since this does not result in a true binary semaphore, advanced binary features like the Priority Inheritance and Priority Ceiling Protocols are not available.

There is currently no text in this section.

9.4 Directives

This section details the semaphore manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

9.4.1 sem init - Initialize an unnamed semaphore

CALLING SEQUENCE:

     int sem_init(
       sem_t        *sem,
       int           pshared,
       unsigned int  value
     );

STATUS CODES:

EINVAL The value argument exceeds SEM VALUE MAX

ENOSPC A resource required to initialize the semaphore has been exhausted The limit on semaphores (SEM VALUE MAX) has been reached

ENOSYS The function sem init is not supported by this implementation

EPERM The process lacks appropriate privileges to initialize the semaphore

DESCRIPTION:

The sem init function is used to initialize the unnamed semaphore referred to by "sem". The value of the initialized semaphore is the parameter "value". The semaphore remains valid until it is destroyed.

ADD MORE HERE XXX

NOTES:

If the functions completes successfully, it shall return a value of zero. Otherwise, it shall return a value of -1 and set "errno" to specify the error that occurred.

Multiprocessing is currently not supported in this implementation.

9.4.2 sem destroy - Destroy an unnamed semaphore

CALLING SEQUENCE:

     int sem_destroy(
       sem_t *sem
     );

STATUS CODES:

EINVAL The value argument exceeds SEM VALUE MAX

ENOSYS The function sem init is not supported by this implementation

EBUSY There are currently processes blocked on the semaphore

DESCRIPTION:

The sem destroy function is used to destroy an unnamed semaphore refered to by "sem". sem destroy can only be used on a semaphore that was created using sem init.

NOTES:

If the functions completes successfully, it shall return a value of zero. Otherwise, it shall return a value of -1 and set "errno" to specify the error that occurred.

Multiprocessing is currently not supported in this implementation.

9.4.3 sem open - Open a named semaphore

CALLING SEQUENCE:

     int sem_open(
       const char *name,
       int         oflag
     );

ARGUMENTS:

The following flag bit may be set in oflag:

O_CREAT - Creates the semaphore if it does not already exist. If O CREAT is set and the semaphore already exists then O CREAT has no effect. Otherwise, sem open() creates a semaphore. The O CREAT flag requires the third and fourth argument: mode and value of type mode t and unsigned int, respectively.

O_EXCL - If O EXCL and O CREAT are set, all call to sem open() shall fail if the semaphore name exists

STATUS CODES:

EACCES Valid name specified but oflag permissions are denied, or the semaphore name specified does not exist and permission to create the named semaphore is denied.

EEXIST O CREAT and O EXCL are set and the named semaphore already exists.

EINTR The sem open() operation was interrupted by a signal.

EINVAL The sem open() operation is not supported for the given name.

EMFILE Too many semaphore descriptors or file descriptors in use by this process.

ENAMETOOLONG The length of the name exceed PATH MAX or name component is longer than NAME MAX while POSIX NO TRUNC is in effect.

ENOENT O CREAT is not set and the named semaphore does not exist.

ENOSPC There is insufficient space for the creation of a new named semaphore.

ENOSYS The function sem open() is not supported by this implementation.

DESCRIPTION:

The sem open() function establishes a connection between a specified semaphore and a process. After a call to sem open with a specified semaphore name, a process can reference to semaphore by the associated name using the address returned by the call. The oflag arguments listed above control the state of the semaphore by determining if the semaphore is created or accessed by a call to sem open().

9.4.4 sem close - Close a named semaphore

CALLING SEQUENCE:

     int sem_close(
       sem_t *sem_close
     );

STATUS CODES:

EACCES The semaphore argument is not a valid semaphore descriptor.

ENOSYS The function sem close is not supported by this implementation.

DESCRIPTION:

The sem close() function is used to indicate that the calling process is finished using the named semaphore indicated by sem. The function sem close deallocates any system resources that were previously allocated by a sem open system call. If sem close() completes successfully it returns a 1, otherwise a value of -1 is return and errno is set.

9.4.5 sem unlink - Unlink a semaphore

CALLING SEQUENCE:

     int sem_unlink(
       const char *name
     );

STATUS CODES:

EACCESS Permission is denied to unlink a semaphore.

ENAMETOOLONG The length of the strong name exceed NAME MAX while POSIX NO TRUNC is in effect.

ENOENT The name of the semaphore does not exist.

ENOSPC There is insufficient space for the creation of a new named semaphore.

ENOSYS The function sem unlink is not supported by this implementation.

DESCRIPTION:

The sem unlink() function shall remove the semaphore name by the string name. If a process is currently accessing the name semaphore, the sem unlink command has no effect. If one or more processes have the semaphore open when the sem unlink function is called, the destruction of semaphores shall be postponed until all reference to semaphore are destroyed by calls to sem close, exit(), or exec. After all references have been destroyed, it returns immediately.

If the termination is successful, the function shall return 0. Otherwise, a -1 is returned and the errno is set.

9.4.6 sem wait - Wait on a Semaphore

CALLING SEQUENCE:

     int sem_wait(
       sem_t *sem
     );

STATUS CODES:

EINVAL The "sem" argument does not refer to a valid semaphore

DESCRIPTION:

This function attempts to lock a semaphore specified by sem. If the semaphore is available, then the semaphore is locked (i.e., the semaphore value is decremented). If the semaphore is unavailable (i.e., the semaphore value is zero), then the function will block until the semaphore becomes available. It will then successfully lock the semaphore. The semaphore remains locked until released by a sem_post() call.

If the call is unsuccessful, then the function returns -1 and sets errno to the appropriate error code.

9.4.7 sem trywait - Non-blocking Wait on a Semaphore

CALLING SEQUENCE:

     int sem_trywait(
       sem_t *sem
     );

STATUS CODES:

EAGAIN The semaphore is not available (i.e., the semaphore value is zero), so the semaphore could not be locked.

EINVAL The sem argument does not refewr to a valid semaphore

DESCRIPTION:

This function attempts to lock a semaphore specified by sem. If the semaphore is available, then the semaphore is locked (i.e., the semaphore value is decremented) and the function returns a value of 0. The semaphore remains locked until released by a sem_post() call. If the semaphore is unavailable (i.e., the semaphore value is zero), then the function will return a value of -1 immediately and set errno to EAGAIN.

If the call is unsuccessful, then the function returns -1 and sets errno to the appropriate error code.

9.4.8 sem timedwait - Wait on a Semaphore for a Specified Time

CALLING SEQUENCE:

     int sem_timedwait(
       sem_t                 *sem,
       const struct timespec *abstime
     );

STATUS CODES:

EAGAIN The semaphore is not available (i.e., the semaphore value is zero), so the semaphore could not be locked.

EINVAL The sem argument does not refewr to a valid semaphore

DESCRIPTION:

This function attemtps to lock a semaphore specified by sem, and will wait for the semaphore until the absolute time specified by abstime. If the semaphore is available, then the semaphore is locked (i.e., the semaphore value is decremented) and the function returns a value of 0. The semaphore remains locked until released by a sem_post() call. If the semaphore is unavailable, then the function will wait for the semaphore to become available for the amount of time specified by timeout.

If the semaphore does not become available within the interval specified by timeout, then the function returns -1 and sets errno to EAGAIN. If any other error occurs, the function returns -1 and sets errno to the appropriate error code.

9.4.9 sem post - Unlock a Semaphore

CALLING SEQUENCE:

     int sem_post(
       sem_t *sem
     );

STATUS CODES:

EINVAL The sem argument does not refer to a valid semaphore

DESCRIPTION:

This function attempts to release the semaphore specified by sem. If other tasks are waiting on the semaphore, then one of those tasks (which one depends on the scheduler being used) is allowed to lock the semaphore and return from its sem_wait(), sem_trywait(), or sem_ timedwait() call. If there are no other tasks waiting on the semaphore, then the semaphore value is simply incremented. sem_post() returns 0 upon successful completion.

If an error occurs, the function returns -1 and sets errno to the appropriate error code.

9.4.10 sem getvalue - Get the value of a semaphore

CALLING SEQUENCE:

     int sem_getvalue(
       sem_t *sem,
       int   *sval
     );

STATUS CODES:

EINVAL The "sem" argument does not refer to a valid semaphore

ENOSYS The function sem getvalue is not supported by this implementation

DESCRIPTION:

The sem getvalue functions sets the location referenced by the "sval" argument to the value of the semaphore without affecting the state of the semaphore. The updated value represents a semaphore value that occurred at some point during the call, but is not necessarily the actual value of the semaphore when it returns to the calling process.

If "sem" is locked, the value returned by sem getvalue will be zero or a negative number whose absolute value is the number of processes waiting for the semaphore at some point during the call.

NOTES:

If the functions completes successfully, it shall return a value of zero. Otherwise, it shall return a value of -1 and set "errno" to specify the error that occurred.

10 Mutex Manager

10.1 Introduction

The mutex manager implements the functionality required of the mutex manager as defined by POSIX 1003.1b-1996. This standard requires that a compliant operating system provide the facilties to ensure that threads can operate with mutual exclusion from one another and defines the API that must be provided. The services provided by the mutex manager are:

• pthread_mutexattr_init - Initialize a Mutex Attribute Set

• pthread_mutexattr_destroy - Destroy a Mutex Attribute Set

• pthread_mutexattr_setprotocol - Set the Blocking Protocol
• pthread_mutexattr_getprotocol - Get the Blocking Protocol
• pthread_mutexattr_setprioceiling - Set the Priority Ceiling
• pthread_mutexattr_getprioceiling - Get the Priority Ceiling
• pthread_mutexattr_setpshared - Set the Visibility
• pthread_mutexattr_getpshared - Get the Visibility
• pthread_mutex_init - Initialize a Mutex
• pthread_mutex_destroy - Destroy a Mutex
• pthread_mutex_lock - Lock a Mutex
• pthread_mutex_trylock - Poll to Lock a Mutex
• pthread_mutex_timedlock - Lock a Mutex with Timeout
• pthread_mutex_unlock - Unlock a Mutex

• pthread_mutex_setprioceiling - Dynamically Set the Priority Ceiling • pthread_mutex_getprioceiling - Dynamically Get the Priority Ceiling

10.2 Background

10.2.1 Mutex Attributes

Mutex attributes are utilized only at mutex creation time. A mutex attribute structure may be initialized and passed as an argument to the mutex_init routine. Note that the priority ceiling of a mutex may be set at run-time.

blocking protcol is the XXX priority ceiling is the XXX pshared is the XXX

10.2.2 PTHREAD MUTEX INITIALIZER

This is a special value that a variable of type pthread_mutex_t may be statically initialized to as shown below:

     pthread_mutex_t my_mutex = PTHREAD_MUTEX_INITIALIZER;

This indicates that my_mutex will be automatically initialized by an implicit call to pthread_ mutex_init the first time the mutex is used. Note that the mutex will be initialized with default attributes.

10.3 Operations

There is currently no text in this section.

10.4 Services

This section details the mutex manager’s services. A subsection is dedicated to each of this manager’s services and describes the calling sequence, related constants, usage, and status codes.

10.4.1 pthread mutexattr init - Initialize a Mutex Attribute Set CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_init(
       pthread_mutexattr_t *attr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

DESCRIPTION:

The pthread_mutexattr_init routine initializes the mutex attributes object specified by attr with the default value for all of the individual attributes.

NOTES:

XXX insert list of default attributes here.

10.4.2 pthread mutexattr destroy - Destroy a Mutex Attribute

Set CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_destroy(
       pthread_mutexattr_t *attr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid. EINVAL The attribute set is not initialized.

DESCRIPTION:

The pthread_mutex_attr_destroy routine is used to destroy a mutex attributes object. The behavior of using an attributes object after it is destroyed is implementation dependent.

NOTES:

NONE

10.4.3 pthread mutexattr setprotocol - Set the Blocking Protocol

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_setprotocol(
       pthread_mutexattr_t *attr,
       int                  protocol
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The protocol argument is invalid.

DESCRIPTION:

The pthread_mutexattr_setprotocol routine is used to set value of the protocol at-

tribute. This attribute controls the order in which threads waiting on this mutex will receive it.

The protocol can be one of the following:

PTHREAD_PRIO_NONE in which case blocking order is FIFO.

PTHREAD_PRIO_INHERIT

in which case blocking order is priority with the priority inheritance protocol in effect.

PTHREAD_PRIO_PROTECT

in which case blocking order is priority with the priority ceiling protocol in effect.

NOTES:

There is currently no way to get simple priority blocking ordering with POSIX mutexes even though this could easily by supported by RTEMS.

10.4.4 pthread mutexattr getprotocol - Get the Blocking Protocol

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_getprotocol(
       pthread_mutexattr_t *attr,
       int                 *protocol
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The protocol pointer argument is invalid.

DESCRIPTION:

The pthread_mutexattr_getprotocol routine is used to obtain the value of the protocol

attribute. This attribute controls the order in which threads waiting on this mutex will receive it.

10.4.5 pthread mutexattr setprioceiling - Set the Priority Ceiling

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_setprioceiling(
       pthread_mutexattr_t *attr,
       int                  prioceiling
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The prioceiling argument is invalid.

DESCRIPTION:

The pthread_mutexattr_setprioceiling routine is used to set value of the prioceiling

attribute. This attribute specifies the priority that is the ceiling for threads obtaining this mutex. Any task obtaining this mutex may not be of greater priority that the ceiling. If it is of lower priority, then its priority will be elevated to prioceiling.

10.4.6 pthread mutexattr getprioceiling - Get the Priority Ceiling

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_getprioceiling(
       const pthread_mutexattr_t *attr,
       int                       *prioceiling
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The prioceiling pointer argument is invalid.

DESCRIPTION:

The pthread_mutexattr_getprioceiling routine is used to obtain the value of the prioceiling attribute. This attribute specifies the priority ceiling for this mutex.

10.4.7 pthread mutexattr setpshared - Set the Visibility

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_setpshared(
       pthread_mutexattr_t *attr,
       int                  pshared
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The pshared argument is invalid.

DESCRIPTION:

10.4.8 pthread mutexattr getpshared - Get the Visibility

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutexattr_getpshared(
       const pthread_mutexattr_t *attr,
       int                       *pshared
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The pshared pointer argument is invalid.

DESCRIPTION:

10.4.9 pthread mutex init - Initialize a Mutex

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_init(
       pthread_mutex_t           *mutex,
       const pthread_mutexattr_t *attr
     );

STATUS CODES:

EINVAL The attribute set is not initialized.

EINVAL The specified protocol is invalid.

EAGAIN The system lacked the necessary resources to initialize another mutex.

ENOMEM Insufficient memory exists to initialize the mutex.

EBUSY Attempted to reinialize the object reference by mutex, a previously initialized, but not yet destroyed.

DESCRIPTION:

10.4.10 pthread mutex destroy - Destroy a Mutex

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_destroy(
       pthread_mutex_t *mutex
     );

STATUS CODES:

EINVAL The specified mutex is invalid.

EBUSY Attempted to destroy the object reference by mutex, while it is locked or referenced by another thread.

DESCRIPTION:

10.4.11 pthread mutex lock - Lock a Mutex

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_lock(
       pthread_mutex_t *mutex
     );

STATUS CODES:

EINVAL The specified mutex is invalid.

EINVAL  The  mutex  has  the  protocol  attribute  of

PTHREAD PRIO PROTECT and the priority of the calling thread is higher than the current priority ceiling.

EDEADLK The current thread already owns the mutex.

DESCRIPTION:

10.4.12 pthread mutex trylock - Poll to Lock a Mutex

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_trylock(
       pthread_mutex_t *mutex
     );

STATUS CODES:

EINVAL The specified mutex is invalid.

EINVAL  The  mutex  has  the  protocol  attribute  of

PTHREAD PRIO PROTECT and the priority of the calling thread is higher than the current priority ceiling.

EDEADLK The current thread already owns the mutex.

DESCRIPTION:

10.4.13 pthread mutex timedlock - Lock a Mutex with Timeout

CALLING SEQUENCE:

     #include <pthread.h>
     #include <time.h>
     int pthread_mutex_timedlock(
       pthread_mutex_t       *mutex,
       const struct timespec *timeout
     );

STATUS CODES:

EINVAL The specified mutex is invalid.

EINVAL The nanoseconds field of timeout is invalid.

EINVAL  The  mutex  has  the  protocol  attribute  of

PTHREAD PRIO PROTECT and the priority of the calling thread is higher than the current priority ceiling.

EDEADLK The current thread already owns the mutex.

ETIMEDOUT The calling thread was unable to obtain the mutex within the specified timeout period.

DESCRIPTION:

10.4.14 pthread mutex unlock - Unlock a Mutex CALLING SEQUENCE:

#include <pthread.h>
int pthread_mutex_unlock(
  pthread_mutex_t *mutex
);

STATUS CODES:

EINVAL The specified mutex is invalid.

DESCRIPTION: NOTES:

10.4.15 pthread mutex setprioceiling - Dynamically Set the Priority Ceiling

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_setprioceiling(
       pthread_mutex_t *mutex,
       int              prioceiling,
       int             *oldceiling
     );

STATUS CODES:

EINVAL The oldceiling pointer parameter is invalid.

EINVAL The prioceiling parameter is an invalid priority.

EINVAL The specified mutex is invalid.

DESCRIPTION:

10.4.16 pthread mutex getprioceiling - Get the Current Priority Ceiling

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_mutex_getprioceiling(
       pthread_mutex_t *mutex,
       int             *prioceiling
     );

STATUS CODES:

EINVAL The prioceiling pointer parameter is invalid.

EINVAL The specified mutex is invalid.

DESCRIPTION:

11 Condition Variable Manager

11.1 Introduction

The condition variable manager ... The directives provided by the condition variable manager are:

• pthread_condattr_init - Initialize a Condition Variable Attribute Set • pthread_condattr_destroy - Destroy a Condition Variable Attribute Set

• pthread_condattr_setpshared - Set Process Shared Attribute
• pthread_condattr_getpshared - Get Process Shared Attribute

• pthread_cond_init - Initialize a Condition Variable

• pthread_cond_destroy - Destroy a Condition Variable

• pthread_cond_signal - Signal a Condition Variable

• pthread_cond_broadcast - Broadcast a Condition Variable

• pthread_cond_wait - Wait on a Condition Variable

• pthread_cond_timedwait - With with Timeout a Condition Variable

11.2 Background

There is currently no text in this section.

11.3 Operations

There is currently no text in this section.

11.4 Directives

This section details the condition variable manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

11.4.1 pthread condattr init - Initialize a Condition Variable Attribute Set

CALLING SEQUENCE:

#include <pthread.h>
int pthread_condattr_init(
  pthread_condattr_t *attr
);

STATUS CODES:

ENOMEM Insufficient memory is available to initialize the condition variable attributes object.

DESCRIPTION:

11.4.2 pthread condattr destroy - Destroy a Condition Variable

Attribute Set CALLING SEQUENCE:

#include <pthread.h>
int pthread_condattr_destroy(
  pthread_condattr_t *attr
);

STATUS CODES:

EINVAL The attribute object specified is invalid.

DESCRIPTION: NOTES:

11.4.3 pthread condattr setpshared - Set Process Shared Attribute

CALLING SEQUENCE:

#include <pthread.h>
int pthread_condattr_setpshared(
  pthread_condattr_t *attr,
  int                 pshared
);

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

11.4.4 pthread condattr getpshared - Get Process Shared Attribute

CALLING SEQUENCE:

#include <pthread.h>
int pthread_condattr_getpshared(
  const pthread_condattr_t *attr,
  int                      *pshared
);

STATUS CODES:

EINVAL Invalid argument passed.

DESCRIPTION:

11.4.5 pthread cond init - Initialize a Condition Variable

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_cond_init(
       pthread_cond_t           *cond,
       const pthread_condattr_t *attr
     );

STATUS CODES:

EAGAIN The system lacked a resource other than memory necessary to create the initialize the condition variable object.

ENOMEM Insufficient memory is available to initialize the condition variable object.

EBUSY The specified condition variable has already been initialized.

EINVAL The specified attribute value is invalid.

DESCRIPTION:

11.4.6 pthread cond destroy - Destroy a Condition Variable CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_cond_destroy(
       pthread_cond_t *cond
     );

STATUS CODES:

EINVAL The specified condition variable is invalid. EBUSY The specified condition variable is currently in use.

DESCRIPTION: NOTES:

11.4.7 pthread cond signal - Signal a Condition Variable

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_cond_signal(
       pthread_cond_t *cond
     );

STATUS CODES:

EINVAL The specified condition variable is not valid.

DESCRIPTION:

NOTES:

This routine should not be invoked from a handler from an asynchronous signal handler or an interrupt service routine.

11.4.8 pthread cond broadcast - Broadcast a Condition Variable

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_cond_broadcast(
       pthread_cond_t *cond
     );

STATUS CODES:

EINVAL The specified condition variable is not valid.

DESCRIPTION:

NOTES:

This routine should not be invoked from a handler from an asynchronous signal handler or an interrupt service routine.

11.4.9 pthread cond wait - Wait on a Condition Variable

CALLING SEQUENCE:

#include <pthread.h>
int pthread_cond_wait(
pthread_cond_t *cond,
pthread_mutex_t *mutex
);

STATUS CODES:

EINVAL The specified condition variable or mutex is not initialized OR different mutexes were specified for concurrent pthread cond wait() and pthread cond timedwait() operations on the same condition variable OR the mutex was not owned by the current thread at the time of the call.

DESCRIPTION:

11.4.10 pthread cond timedwait - Wait with Timeout a Condition Variable

CALLING SEQUENCE:

#include <pthread.h>
int pthread_cond_timedwait(
pthread_cond_t        *cond,
pthread_mutex_t       *mutex,
const struct timespec *abstime
);

STATUS CODES:

EINVAL The specified condition variable or mutex is not initialized OR different mutexes were specified for concurrent pthread cond wait() and pthread cond timedwait() operations on the same condition variable OR the mutex was not owned by the current thread at the time of the call.

ETIMEDOUT The specified time has elapsed without the condition variable being satisfied.

DESCRIPTION:

12 Memory Management Manager

12.1 Introduction

The memory management manager is ... The directives provided by the memory management manager are:

• mlockall - Lock the Address Space of a Process

• munlockall - Unlock the Address Space of a Process • mlock - Lock a Range of the Process Address Space

• munlock - Unlock a Range of the Process Address Space • mmap - Map Process Addresses to a Memory Object

• munmap - Unmap Previously Mapped Addresses

• mprotect - Change Memory Protection

• msync - Memory Object Synchronization

• shm_open - Open a Shared Memory Object

• shm_unlink - Remove a Shared Memory Object

12.2 Background

There is currently no text in this section.

12.3 Operations

There is currently no text in this section.

12.4 Directives

This section details the memory management manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

12.4.1 mlockall - Lock the Address Space of a Process CALLING SEQUENCE:

     int mlockall(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.2 munlockall - Unlock the Address Space of a Process CALLING SEQUENCE:

     int munlockall(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.3 mlock - Lock a Range of the Process Address Space CALLING SEQUENCE:

     int mlock(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.4 munlock - Unlock a Range of the Process Address Space CALLING SEQUENCE:

     int munlock(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.5 mmap - Map Process Addresses to a Memory Object CALLING SEQUENCE:

     int mmap(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.6 munmap - Unmap Previously Mapped Addresses CALLING SEQUENCE:

     int munmap(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.7 mprotect - Change Memory Protection CALLING SEQUENCE:

     int mprotect(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.8 msync - Memory Object Synchronization CALLING SEQUENCE:

     int msync(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.9 shm open - Open a Shared Memory Object CALLING SEQUENCE:

     int shm_open(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

12.4.10 shm unlink - Remove a Shared Memory Object CALLING SEQUENCE:

     int shm_unlink(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

13 Scheduler Manager

13.1 Introduction

The scheduler manager ...

The directives provided by the scheduler manager are:

• sched_get_priority_min - Get Minimum Priority Value • sched_get_priority_max - Get Maximum Priority Value • sched_rr_get_interval - Get Timeslicing Quantum • sched_yield - Yield the Processor

13.2 Background

13.2.1 Priority

In the RTEMS implementation of the POSIX API, the priorities range from the low prior-

ity of sched_get_priority_min() to the highest priority of sched_get_priority_max().

Numerically higher values represent higher priorities.

13.2.2 Scheduling Policies

The following scheduling policies are available:

SCHED FIFO Priority-based, preemptive scheduling with no timeslicing. This is equivalent to what is called "manual round-robin" scheduling.

SCHED RR Priority-based, preemptive scheduling with timeslicing. Time quantums are maintained on a per-thread basis and are not reset at each context switch. Thus, a thread which is preempted and subsequently resumes execution will attempt to complete the unused portion of its time quantum.

SCHED OTHER Priority-based, preemptive scheduling with timeslicing. Time quantums are maintained on a per-thread basis and are reset at each context switch.

SCHED SPORADIC Priority-based, preemptive scheduling utilizing three additional parameters: budget, replenishment period, and low priority. Under this policy, the thread is allowed to execute for "budget" amount of time before its priority is lowered to "low priority". At the end of each replenishment period, the thread resumes its initial priority and has its budget replenished.

13.3 Operations

There is currently no text in this section.

13.4 Directives

This section details the scheduler manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

13.4.1 sched get priority min - Get Minimum Priority Value CALLING SEQUENCE:

     #include <sched.h>
     int sched_get_priority_min(
       int policy
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINVAL The indicated policy is invalid.

DESCRIPTION:

This routine return the minimum (numerically and logically lowest) priority for the specified

policy.

NOTES:

NONE

13.4.2 sched get priority max - Get Maximum Priority Value CALLING SEQUENCE:

     #include <sched.h>
     int sched_get_priority_max(
       int policy
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINVAL The indicated policy is invalid.

DESCRIPTION:

This routine return the maximum (numerically and logically highest) priority for the spec-

ified policy.

NOTES:

NONE

13.4.3 sched rr get interval - Get Timeslicing Quantum CALLING SEQUENCE:

     #include <sched.h>
     int sched_rr_get_interval(
       pid_t            pid,
       struct timespec *interval
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: ESRCH The indicated process id is invalid. EINVAL The specified interval pointer parameter is invalid.

DESCRIPTION:

This routine returns the length of the timeslice quantum in the interval parameter for the specified pid.

NOTES:

The pid argument should be 0 to indicate the calling process.

13.4.4 sched yield - Yield the Processor CALLING SEQUENCE:

     #include <sched.h>
     int sched_yield( void );

STATUS CODES:

This routine always returns zero to indicate success.

DESCRIPTION:

This call forces the calling thread to yield the processor to another thread. Normally this is used to implement voluntary round-robin task scheduling.

NOTES:

NONE

14 Clock Manager

14.1 Introduction

The clock manager provides services two primary classes of services. The first focuses on obtaining and setting the current date and time. The other category of services focus on allowing a thread to delay for a specific length of time.

The directives provided by the clock manager are:

• clock_gettime - Obtain Time of Day

• clock_settime - Set Time of Day

• clock_getres - Get Clock Resolution

• sleep - Delay Process Execution

• usleep - Delay Process Execution in Microseconds • nanosleep - Delay with High Resolution

• gettimeofday - Get the Time of Day

• time - Get time in seconds

14.2 Background

There is currently no text in this section.

14.3 Operations

There is currently no text in this section.

14.4 Directives

This section details the clock manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

14.4.1 clock gettime - Obtain Time of Day

CALLING SEQUENCE:

     #include <time.h>
     int clock_gettime(
       clockid_t        clock_id,
       struct timespec *tp
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following:

EINVAL The tp pointer parameter is invalid.

EINVAL The clock id specified is invalid.

236 RTEMS POSIX API User’s Guide

DESCRIPTION: NOTES:

NONE

14.4.2 clock settime - Set Time of Day CALLING SEQUENCE:

     #include <time.h>
     int clock_settime(
       clockid_t              clock_id,
       const struct timespec *tp
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINVAL The tp pointer parameter is invalid. EINVAL The clock id specified is invalid. EINVAL The contents of the tp structure are invalid.

DESCRIPTION: NOTES:

NONE

14.4.3 clock getres - Get Clock Resolution CALLING SEQUENCE:

     #include <time.h>
     int clock_getres(
       clockid_t        clock_id,
       struct timespec *res
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following: EINVAL The res pointer parameter is invalid. EINVAL The clock id specified is invalid.

DESCRIPTION: NOTES:

If res is NULL, then the resolution is not returned.

14.4.4 sleep - Delay Process Execution CALLING SEQUENCE:

     #include <unistd.h>
     unsigned int sleep(
       unsigned int seconds
     );

STATUS CODES:

This routine returns the number of unslept seconds.

DESCRIPTION:

The sleep() function delays the calling thread by the specified number of seconds.

NOTES:

This call is interruptible by a signal.

14.4.5 usleep - Delay Process Execution in Microseconds

CALLING SEQUENCE:

     #include <time.h>
     useconds_t usleep(
       useconds_t useconds
     );

STATUS CODES:

This routine returns the number of unslept seconds.

DESCRIPTION:

The sleep() function delays the calling thread by the specified number of seconds.

The usleep() function suspends the calling thread from execution until either the number of microseconds specified by the useconds argument has elapsed or a signal is delivered to the calling thread and its action is to invoke a signal-catching function or to terminate the process.

Because of other activity, or because of the time spent in processing the call, the actual length of time the thread is blocked may be longer than the amount of time specified.

NOTES:

This call is interruptible by a signal.

The Single UNIX Specification allows this service to be implemented using the same timer as that used by the alarm() service. This is NOT the case for RTEMS and this call has no interaction with the SIGALRM signal.

14.4.6 nanosleep - Delay with High Resolution

CALLING SEQUENCE:

     #include <time.h>
     int nanosleep(
       const struct timespec *rqtp,
       struct timespec       *rmtp
     );

STATUS CODES:

On error, this routine returns -1 and sets errno to one of the following:

EINTR The routine was interrupted by a signal.

EAGAIN The requested sleep period specified negative seconds or nanoseconds.

EINVAL The requested sleep period specified an invalid number for the nanoseconds field.

DESCRIPTION:

NOTES:

This call is interruptible by a signal.

14.4.7 gettimeofday - Get the Time of Day CALLING SEQUENCE:

     #include <sys/time.h>
     #include <unistd.h>
     int gettimeofday(
       struct timeval  *tp,
       struct timezone *tzp
     );

STATUS CODES:

On error, this routine returns -1 and sets errno as appropriate. EPERM settimeofdat is called by someone other than the superuser. EINVAL Timezone (or something else) is invalid. EFAULT One of tv or tz pointed outside your accessible address space

DESCRIPTION:

This routine returns the current time of day in the tp structure.

NOTES:

Currently, the timezone information is not supported. The tzp argument is ignored.

14.4.8 time - Get time in seconds CALLING SEQUENCE:

     #include <time.h>
     int time(
       time_t *tloc
     );

STATUS CODES:

This routine returns the number of seconds since the Epoch.

DESCRIPTION:

time returns the time since 00:00:00 GMT, January 1, 1970, measured in seconds If tloc in non null, the return value is also stored in the memory pointed to by t.

NOTES:

NONE

15 Timer Manager

15.1 Introduction

The timer manager is ... The services provided by the timer manager are:

• timer_create - Create a Per-Process Timer • timer_delete - Delete a Per-Process Timer

• timer_settime - Set Next Timer Expiration • timer_gettime - Get Time Remaining on Timer • timer_getoverrun - Get Timer Overrun Count

15.2 Background

15.3 Operations

15.4 System Calls

This section details the timer manager’s services. A subsection is dedicated to each of this manager’s services and describes the calling sequence, related constants, usage, and status codes.

15.4.1 timer create - Create a Per-Process Timer

CALLING SEQUENCE:

#include <time.h>
#include <signal.h>
int timer_create(
  clockid_t        clock_id,
  struct sigevent *evp,
  timer_t         *timerid
);

STATUS CODES:

EXXX -

DESCRIPTION:

15.4.2 timer delete - Delete a Per-Process Timer CALLING SEQUENCE:

     #include <time.h>
     int timer_delete(
       timer_t timerid
     );

STATUS CODES:

EXXX -

DESCRIPTION: NOTES:

15.4.3 timer settime - Set Next Timer Expiration

CALLING SEQUENCE:

#include <time.h>
int timer_settime(
timer_t                  timerid,
int                      flags,
const struct itimerspec *value,
struct itimerspec       *ovalue
);

STATUS CODES:

EXXX -

DESCRIPTION:

15.4.4 timer gettime - Get Time Remaining on Timer

CALLING SEQUENCE:

#include <time.h>
int timer_gettime(
  timer_t            timerid,
  struct itimerspec *value
);

STATUS CODES:

EXXX -

DESCRIPTION:

15.4.5 timer getoverrun - Get Timer Overrun Count CALLING SEQUENCE:

     #include <time.h>
     int timer_getoverrun(
       timer_t   timerid
     );

STATUS CODES:

EXXX -

DESCRIPTION: NOTES:

16 Message Passing Manager

16.1 Introduction

The message passing manager is the means to provide communication and synchronization capabilities using POSIX message queues.

The directives provided by the message passing manager are:

• mq_open - Open a Message Queue

• mq_close - Close a Message Queue

• mq_unlink - Remove a Message Queue

• mq_send - Send a Message to a Message Queue

• mq_receive - Receive a Message from a Message Queue • mq_notify - Notify Process that a Message is Available • mq_setattr - Set Message Queue Attributes

• mq_getattr - Get Message Queue Attributes

16.2 Background

16.2.1 Theory

Message queues are named objects that operate with readers and writers. In addition, a message queue is a priority queue of discrete messages. POSIX message queues offer a certain, basic amount of application access to, and control over, the message queue geometry that can be changed.

16.2.2 Messages

A message is a variable length buffer where information can be stored to support communication. The length of the message and the information stored in that message are user-defined and can be actual data, pointer(s), or empty. There is a maximum acceptable length for a message that is associated with each message queue.

16.2.3 Message Queues

Message queues are named objects similar to the pipes of POSIX. They are a means of communicating data between multiple processes and for passing messages among tasks and ISRs. Message queues can contain a variable number of messages from 0 to an upper limit that is user defined. The maximum length of the message can be set on a per message queue basis. Normally messages are sent and received from the message queue in FIFO order. However, messages can also be prioritized and a priority queue established for the passing of messages. Synchronization is needed when a task waits for a message to arrive at a queue. Also, a task may poll a queue for the arrival of a message.

The message queue descriptor mqd_t represents the message queue. It is passed as an argument to all of the message queue functions.

16.2.4 Building a Message Queue Attribute Set

The mq attr structure is used to define the characteristics of the message queue.

     typedef struct mq_attr{
       long mq_flags;
       long mq_maxmsg;
       long mq_msgsize;
       long mq_curmsgs;
     };

All of these attributes are set when the message queue is created using mq open. The mq flags field is not used in the creation of a message queue, it is only used by mq setattr and mq getattr. The structure mq attr is passed as an argument to mq setattr and mq getattr.

The mq flags contain information affecting the behavior of the message queue. The O NONBLOCK mq flag is the only flag that is defined. In mq setattr, the mq flag can be set to dynamically change the blocking and non-blocking behavior of the message queue. If the non-block flag is set then the message queue is non-blocking, and requests to send and receive messages do not block waiting for resources. For a blocking message queue, a request to send might have to wait for an empty message queue, and a request to receive might have to wait for a message to arrive on the queue. Both mq maxmsg and mq msgsize affect the sizing of the message queue. mq maxmsg specifies how many messages the queue can hold at any one time. mq msgsize specifies the size of any one message on the queue. If either of these limits is exceeded, an error message results.

Upon return from mq getattr, the mq curmsgs is set according to the current state of the message queue. This specifies the number of messages currently on the queue.

16.2.5 Notification of a Message on the Queue

Every message queue has the ability to notify one (and only one) process whenever the queue’s state changes from empty (0 messages) to nonempty. This means that the process does not have to block or constantly poll while it waits for a message. By calling mq notify, you can attach a notification request to a message queue. When a message is received by an empty queue, if there are no processes blocked and waiting for the message, then the queue notifies the requesting process of a message arrival. There is only one signal sent by the message queue, after that the notification request is de-registered and another process can attach its notification request. After receipt of a notification, a process must re-register if it wishes to be notified again.

If there is a process blocked and waiting for the message, that process gets the message, and notification is not sent. It is also possible for another process to receive the message after the notification is sent but before the notified process has sent its receive request.

Only one process can have a notification request attached to a message queue at any one time. If another process attempts to register a notification request, it fails. You can deregister for a message queue by passing a NULL to mq notify, this removes any notification request attached to the queue. Whenever the message queue is closed, all notification attachments are removed.

16.2.6 POSIX Interpretation Issues

There is one significant point of interpretation related to the RTEMS implementation of POSIX message queues:

What happens to threads already blocked on a message queue when the mode of that same message queue is changed from blocking to non-blocking?

The RTEMS POSIX implementation decided to unblock all waiting tasks with an EAGAIN status just as if a non-blocking version of the same operation had returned unsatisfied. This case is not discussed in the POSIX standard and other implementations may have chosen alternative behaviors.

16.3 Operations

16.3.1 Opening or Creating a Message Queue

If the message queue already exists, mq open() opens it, if the message queue does not exist, mq open() creates it. When a message queue is created, the geometry of the message queue is contained in the attribute structure that is passed in as an argument. This includes mq msgsize that dictates the maximum size of a single message, and the mq maxmsg that dictates the maximum number of messages the queue can hold at one time. The blocking or non-blocking behavior of the queue can also specified.

16.3.2 Closing a Message Queue

The mq close() function is used to close the connection made to a message queue that was made during mq open. The message queue itself and the messages on the queue are persistent and remain after the queue is closed.

16.3.3 Removing a Message Queue

The mq unlink() function removes the named message queue. If the message queue is not open when mq unlink is called, then the queue is immediately eliminated. Any messages that were on the queue are lost, and the queue can not be opened again. If processes have the queue open when mq unlink is called, the removal of the queue is delayed until the last process using the queue has finished. However, the name of the message queue is removed so that no other process can open it.

16.3.4 Sending a Message to a Message Queue

The mq send() function adds the message in priority order to the message queue. Each message has an assigned a priority. The highest priority message is be at the front of the queue.

The maximum number of messages that a message queue may accept is specified at creation by the mq maxmsg field of the attribute structure. If this amount is exceeded, the behavior of the process is determined according to what oflag was used when the message queue was opened. If the queue was opened with O NONBLOCK flag set, the process does not block, and an error is returned. If the O NONBLOCK flag was not set, the process does block and wait for space on the queue.

16.3.5 Receiving a Message from a Message Queue

The mq receive() function is used to receive the oldest of the highest priority message(s) from the message queue specified by mqdes. The messages are received in FIFO order within the priorities. The received message’s priority is stored in the location referenced by the msg prio. If the msg prio is a NULL, the priority is discarded. The message is removed and stored in an area pointed to by msg ptr whose length is of msg len. The msg len must be at least equal to the mq msgsize attribute of the message queue.

The blocking behavior of the message queue is set by O NONBLOCK at mq open or by setting O NONBLOCK in mq flags in a call to mq setattr. If this is a blocking queue, the process does block and wait on an empty queue. If this a non-blocking queue, the process does not block. Upon successful completion, mq receive returns the length of the selected message in bytes and the message is removed from the queue.

16.3.6 Notification of Receipt of a Message on an Empty Queue

The mq notify() function registers the calling process to be notified of message arrival at an empty message queue. Every message queue has the ability to notify one (and only one) process whenever the queue’s state changes from empty (0 messages) to nonempty. This means that the process does not have to block or constantly poll while it waits for a message. By calling mq notify, a notification request is attached to a message queue. When a message is received by an empty queue, if there are no processes blocked and waiting for the message, then the queue notifies the requesting process of a message arrival. There is only one signal sent by the message queue, after that the notification request is de-registered and another process can attach its notification request. After receipt of a notification, a process must re-register if it wishes to be notified again.

If there is a process blocked and waiting for the message, that process gets the message, and notification is not sent. Only one process can have a notification request attached to a message queue at any one time. If another process attempts to register a notification request, it fails. You can de-register for a message queue by passing a NULL to mq notify, this removes any notification request attached to the queue. Whenever the message queue is closed, all notification attachments are removed.

16.3.7 Setting the Attributes of a Message Queue

The mq setattr() function is used to set attributes associated with the open message queue description referenced by the message queue descriptor specified by mqdes. The *omqstat represents the old or previous attributes. If omqstat is non-NULL, the function mq setattr() stores, in the location referenced by omqstat, the previous message queue attributes and the current queue status. These values are the same as would be returned by a call to mq getattr() at that point.

There is only one mq attr.mq flag that can be altered by this call. This is the flag that deals with the blocking and non-blocking behavior of the message queue. If the flag is set then the message queue is non-blocking, and requests to send or receive do not block while waiting for resources. If the flag is not set, then message send and receive may involve waiting for an empty queue or waiting for a message to arrive.

16.3.8 Getting the Attributes of a Message Queue

The mq getattr() function is used to get status information and attributes of the message queue associated with the message queue descriptor. The results are returned in the mq attr structure referenced by the mqstat argument. All of these attributes are set at create time, except the blocking/non-blocking behavior of the message queue which can be dynamically set by using mq setattr. The attribute mq curmsg is set to reflect the number of messages on the queue at the time that mq getattr was called.

16.4 Directives

This section details the message passing manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

16.4.1 mq open - Open a Message Queue

CALLING SEQUENCE:

     #include <mqueue.h>
     mqd_t mq_open(
       const char     *name,
       int             oflag,
       mode_t          mode,
       struct mq_attr *attr
     );

STATUS CODES:

EACCES - Either the message queue exists and the permissions requested in oflags were denied, or the message does not exist and permission to create one is denied.

EEXIST - You tried to create a message queue that already exists.

EINVAL - An inappropriate name was given for the message queue, or the values of mqmaxmsg or mq msgsize were less than 0.

ENOENT - The message queue does not exist, and you did not specify to create it.

EINTR - The call to mq open was interrupted by a signal.

EMFILE - The process has too many files or message queues open. This is a process limit error.

ENFILE - The system has run out of resources to support more open message queues. This is a system error.

ENAMETOOLONG - mq name is too long.

DESCRIPTION:

The mq open () function establishes the connection between a process and a message queue with a message queue descriptor. If the message queue already exists, mq open opens it, if the message queue does not exist, mq open creates it. Message queues can have multiple senders and receivers. If mq open is successful, the function returns a message queue descriptor. Otherwise, the function returns a -1 and sets ’errno’ to indicate the error.

The name of the message queue is used as an argument. For the best of portability, the name of the message queue should begin with a "/" and no other "/" should be in the name. Different systems interpret the name in different ways.

The oflags contain information on how the message is opened if the queue already exists. This may be O RDONLY for read only, O WRONLY for write only, of O RDWR, for read and write.

In addition, the oflags contain information needed in the creation of a message queue. O_ NONBLOCK - If the non-block flag is set then the message queue is non-blocking, and requests to send and receive messages do not block waiting for resources. If the flag is not set then the message queue is blocking, and a request to send might have to wait for an empty message queue. Similarly, a request to receive might have to wait for a message to arrive on the queue. O_CREAT - This call specifies that the call the mq open is to create a new message queue. In this case the mode and attribute arguments of the function call are utilized. The message queue is created with a mode similar to the creation of a file, read and write permission creator, group, and others.

The geometry of the message queue is contained in the attribute structure. This includes mq msgsize that dictates the maximum size of a single message, and the mq maxmsg that dictates the maximum number of messages the queue can hold at one time. If a NULL is used in the mq attr argument, then the message queue is created with implementation defined defaults. O_EXCL - is always set if O CREAT flag is set. If the message queue already exists, O EXCL causes an error message to be returned, otherwise, the new message queue fails and appends to the existing one.

NOTES:

The mq open () function does not add or remove messages from the queue. When a new message queue is being created, the mq flag field of the attribute structure is not used.

16.4.2 mq close - Close a Message Queue

CALLING SEQUENCE:

     #include <mqueue.h>
     int mq_close(
       mqd_t mqdes
     );

STATUS CODES:

EINVAL - The descriptor does not represent a valid open message queue

DESCRIPTION:

The mq close function removes the association between the message queue descriptor, mqdes, and its message queue. If mq close() is successfully completed, the function returns a value of zero; otherwise, the function returns a value of -1 and sets errno to indicate the error.

NOTES:

If the process had successfully attached a notification request to the message queue via mq notify, this attachment is removed, and the message queue is available for another process to attach for notification. mq close has no effect on the contents of the message queue, all the messages that were in the queue remain in the queue.

16.4.3 mq unlink - Remove a Message Queue

CALLING SEQUENCE:

     #include <mqueue.h>
     int mq_unlink(
       const char *name
     );

STATUS CODES:

EINVAL - The descriptor does not represent a valid message queue

DESCRIPTION:

The mq unlink() function removes the named message queue. If the message queue is not open when mq unlink is called, then the queue is immediately eliminated. Any messages that were on the queue are lost, and the queue can not be opened again. If processes have the queue open when mq unlink is called, the removal of the queue is delayed until the last process using the queue has finished. However, the name of the message queue is removed so that no other process can open it. Upon successful completion, the function returns a value of zero. Otherwise, the named message queue is not changed by this function call, and the function returns a value of -1 and sets errno to indicate the error.

NOTES:

Calls to mq open() to re-create the message queue may fail until the message queue is actually removed. However, the mq unlink() call need not block until all references have been closed; it may return immediately.

16.4.4 mq send - Send a Message to a Message Queue

CALLING SEQUENCE:

     #include<mqueue.h>
     int mq_send(
       mqd_t        mqdes,
       const char  *msg_ptr,
       size_t       msg_len,
       unsigned int msg_prio
     );

STATUS CODES:

EBADF - The descriptor does not represent a valid message queue, or the queue was opened for read only O RDONLY EINVAL - The value of msg prio was greater than the MQ PRIO MAX. EMSGSIZE - The msg len is greater than the mq msgsize attribute of the message queue EAGAIN - The message queue is non-blocking, and there is no room on the queue for another message as specified by the mq maxmsg. EINTR - The message queue is blocking. While the process was waiting for free space on the queue, a signal arrived that interrupted the wait.

DESCRIPTION:

The mq send() function adds the message pointed to by the argument msg ptr to the message queue specified by mqdes. Each message is assigned a priority , from 0 to MQ PRIO MAX. MQ PRIO MAX is defined in <limits.h> and must be at least 32. Messages are added to the queue in order of their priority. The highest priority message is at the front of the queue.

The maximum number of messages that a message queue may accept is specified at creation by the mq maxmsg field of the attribute structure. If this amount is exceeded, the behavior of the process is determined according to what oflag was used when the message queue was opened. If the queue was opened with O NONBLOCK flag set, then the EAGAIN error is returned. If the O NONBLOCK flag was not set, the process blocks and waits for space on the queue, unless it is interrupted by a signal.

Upon successful completion, the mq send () function returns a value of zero. Otherwise, no message is enqueued, the function returns -1, and errno is set to indicate the error.

NOTES:

If the specified message queue is not full, mq send inserts the message at the position indicated by the msg prio argument.

16.4.5 mq receive - Receive a Message from a Message Queue

CALLING SEQUENCE:

     #include <mqueue.h>
     size_t mq_receive(
       mqd_t         mqdes,
       char         *msg_ptr,
       size_t        msg_len,
       unsigned int *msg_prio
     );

STATUS CODES:

EBADF - The descriptor does not represent a valid message queue, or the queue was opened for write only O WRONLY EMSGSIZE - The msg len is less than the mq msgsize attribute of the message queue EAGAIN - The message queue is non-blocking, and the queue is empty EINTR - The message queue is blocking. While the process was waiting for a message to arrive on the queue, a signal arrived that interrupted the wait.

DESCRIPTION:

The mq receive function is used to receive the oldest of the highest priority message(s) from the message queue specified by mqdes. The messages are received in FIFO order within the priorities. The received message’s priority is stored in the location referenced by the msg prio. If the msg prio is a NULL, the priority is discarded. The message is removed and stored in an area pointed to by msg ptr whose length is of msg len. The msg len must be at least equal to the mq msgsize attribute of the message queue.

The blocking behavior of the message queue is set by O NONBLOCK at mq open or by setting O NONBLOCK in mq flags in a call to mq setattr. If this is a blocking queue, the process blocks and waits on an empty queue. If this a non-blocking queue, the process does not block.

Upon successful completion, mq receive returns the length of the selected message in bytes and the message is removed from the queue. Otherwise, no message is removed from the queue, the function returns a value of -1, and sets errno to indicate the error.

NOTES:

If the size of the buffer in bytes, specified by the msg len argument, is less than the mq msgsize attribute of the message queue, the function fails and returns an error

16.4.6 mq notify - Notify Process that a Message is Available

CALLING SEQUENCE:

     #include <mqueue.h>
     int mq_notify(
       mqd_t                  mqdes,
       const struct sigevent *notification
     );

STATUS CODES:

EBADF - The descriptor does not refer to a valid message queue EBUSY - A notification request is already attached to the queue

DESCRIPTION:

If the argument notification is not NULL, this function registers the calling process to be notified of message arrival at an empty message queue associated with the specified message queue descriptor, mqdes.

Every message queue has the ability to notify one (and only one) process whenever the queue’s state changes from empty (0 messages) to nonempty. This means that the process does not have to block or constantly poll while it waits for a message. By calling mq notify, a notification request is attached to a message queue. When a message is received by an empty queue, if there are no processes blocked and waiting for the message, then the queue notifies the requesting process of a message arrival. There is only one signal sent by the message queue, after that the notification request is de-registered and another process can attach its notification request. After receipt of a notification, a process must re-register if it wishes to be notified again.

If there is a process blocked and waiting for the message, that process gets the message, and notification is not be sent. Only one process can have a notification request attached to a message queue at any one time. If another process attempts to register a notification request, it fails. You can de-register for a message queue by passing a NULL to mq notify; this removes any notification request attached to the queue. Whenever the message queue is closed, all notification attachments are removed.

Upon successful completion, mq notify returns a value of zero; otherwise, the function returns a value of -1 and sets errno to indicate the error.

NOTES:

It is possible for another process to receive the message after the notification is sent but before the notified process has sent its receive request.

16.4.7 mq setattr - Set Message Queue Attributes

CALLING SEQUENCE:

     #include <mqueue.h>
     int mq_setattr(
       mqd_t                 mqdes,
       const struct mq_attr *mqstat,
       struct mq_attr       *omqstat
     );

STATUS CODES:

EBADF - The message queue descriptor does not refer to a valid, open queue. EINVAL - The mq flag value is invalid.

DESCRIPTION:

The mq setattr function is used to set attributes associated with the open message queue description referenced by the message queue descriptor specified by mqdes. The *omqstat represents the old or previous attributes. If omqstat is non-NULL, the function mq setattr() stores, in the location referenced by omqstat, the previous message queue attributes and the current queue status. These values are the same as would be returned by a call to mq getattr() at that point.

There is only one mq attr.mq flag which can be altered by this call. This is the flag that deals with the blocking and non-blocking behavior of the message queue. If the flag is set then the message queue is non-blocking, and requests to send or receive do not block while waiting for resources. If the flag is not set, then message send and receive may involve waiting for an empty queue or waiting for a message to arrive.

Upon successful completion, the function returns a value of zero and the attributes of the message queue have been changed as specified. Otherwise, the message queue attributes is unchanged, and the function returns a value of -1 and sets errno to indicate the error.

NOTES:

All other fields in the mq attr are ignored by this call.

16.4.8 mq getattr - Get Message Queue Attributes

CALLING SEQUENCE:

     #include <mqueue.h>
     int mq_getattr(
       mqd_t mqdes,
       struct mq_attr *mqstat
     );

STATUS CODES:

EBADF - The message queue descriptor does not refer to a valid, open message queue.

DESCRIPTION:

The mqdes argument specifies a message queue descriptor. The mq getattr function is used to get status information and attributes of the message queue associated with the message queue descriptor. The results are returned in the mq attr structure referenced by the mqstat argument. All of these attributes are set at create time, except the blocking/non-blocking behavior of the message queue which can be dynamically set by using mq setattr. The attribute mq curmsg is set to reflect the number of messages on the queue at the time that mq getattr was called.

Upon successful completion, the mq getattr function returns zero. Otherwise, the function returns -1 and sets errno to indicate the error.

17 Thread Manager

17.1 Introduction

The thread manager implements the functionality required of the thread manager as defined by POSIX 1003.1b-1996. This standard requires that a compliant operating system provide the facilties to manage multiple threads of control and defines the API that must be provided. The services provided by the thread manager are:

• pthread_attr_init - Initialize a Thread Attribute Set

• pthread_attr_destroy - Destroy a Thread Attribute Set

• pthread_attr_setdetachstate - Set Detach State
• pthread_attr_getdetachstate - Get Detach State
• pthread_attr_setstacksize - Set Thread Stack Size
• pthread_attr_getstacksize - Get Thread Stack Size
• pthread_attr_setstackaddr - Set Thread Stack Address
• pthread_attr_getstackaddr - Get Thread Stack Address

• pthread_attr_setscope - Set Thread Scheduling Scope

• pthread_attr_getscope - Get Thread Scheduling Scope

• pthread_attr_setinheritsched - Set Inherit Scheduler Flag
• pthread_attr_getinheritsched - Get Inherit Scheduler Flag
• pthread_attr_setschedpolicy - Set Scheduling Policy
• pthread_attr_getschedpolicy - Get Scheduling Policy
• pthread_attr_setschedparam - Set Scheduling Parameters
• pthread_attr_getschedparam - Get Scheduling Parameters

• pthread_create - Create a Thread

• pthread_exit - Terminate the Current Thread

• pthread_detach - Detach a Thread

• pthread_join - Wait for Thread Termination

• pthread_self - Get Thread ID

• pthread_equal - Compare Thread IDs

• pthread_once - Dynamic Package Initialization

• pthread_setschedparam - Set Thread Scheduling Parameters • pthread_getschedparam - Get Thread Scheduling Parameters

17.2 Background

17.2.1 Thread Attributes

Thread attributes are utilized only at thread creation time. A thread attribute structure may be initialized and passed as an argument to the pthread_create routine.

stack address is the address of the optionally user specified stack area for this thread. If this value is NULL, then RTEMS allocates the memory for the thread stack from the RTEMS Workspace Area. Otherwise, this is the user specified address for the memory to be used for the thread’s stack. Each thread must have a distinct stack area. Each processor family has different alignment rules which should be followed.

stack size is the minimum desired size for this thread’s stack area. If the size of this area as specified by the stack size attribute is smaller than the minimum for this processor family and the stack is not user specified, then RTEMS will automatically allocate a stack of the minimum size for this processor family.

contention scope specifies the scheduling contention scope. RTEMS only supports the PTHREAD SCOPE PROCESS scheduling contention scope.

scheduling inheritance specifies whether a user specified or the scheduling policy and parameters of the currently executing thread are to be used. When this is PTHREAD INHERIT SCHED, then the scheduling policy and parameters of the currently executing thread are inherited by the newly created thread.

scheduling policy and parameters specify the manner in which the thread will contend for the processor. The scheduling parameters are interpreted based on the specified policy. All policies utilize the thread priority parameter.

17.3 Operations

There is currently no text in this section.

17.4 Services

This section details the thread manager’s services. A subsection is dedicated to each of this manager’s services and describes the calling sequence, related constants, usage, and status codes.

17.4.1 pthread attr init - Initialize a Thread Attribute Set

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_init(
       pthread_attr_t *attr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

DESCRIPTION:

The pthread_attr_init routine initializes the thread attributes object specified by attr with the default value for all of the individual attributes.

NOTES:

The settings in the default attributes are implementation defined. For RTEMS, the default attributes are as follows:

• stackadr is not set to indicate that RTEMS is to allocate the stack memory.

    • stacksize is set to PTHREAD_MINIMUM_STACK_SIZE.

• contentionscope is set to PTHREAD_SCOPE_PROCESS.

• inheritsched is set to PTHREAD_INHERIT_SCHED to indicate that the created thread inherits its scheduling attributes from its parent.

    • detachstate is set to PTHREAD_CREATE_JOINABLE.

17.4.2 pthread attr destroy - Destroy a Thread Attribute Set CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_destroy(
       pthread_attr_t *attr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid. EINVAL The attribute set is not initialized.

DESCRIPTION:

The pthread_attr_destroy routine is used to destroy a thread attributes object. The behavior of using an attributes object after it is destroyed is implementation dependent.

NOTES:

NONE

17.4.3 pthread attr setdetachstate - Set Detach State

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setdetachstate(
       pthread_attr_t *attr,
       int             detachstate
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The detachstate argument is invalid.

DESCRIPTION:

The pthread_attr_setdetachstate routine is used to value of the detachstate attribute.

This attribute controls whether the thread is created in a detached state.

The detachstate can be either PTHREAD_CREATE_DETACHED or PTHREAD_CREATE_JOINABLE.

The default value for all threads is PTHREAD_CREATE_JOINABLE.

NOTES:

If a thread is in a detached state, then the use of the ID with the pthread_detach or pthread_join routines is an error.

17.4.4 pthread attr getdetachstate - Get Detach State

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getdetachstate(
       const pthread_attr_t *attr,
       int                  *detachstate
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The detatchstate pointer argument is invalid.

DESCRIPTION:

The pthread_attr_getdetachstate routine is used to obtain the current value of the detachstate attribute as specified by the attr thread attribute object.

17.4.5 pthread attr setstacksize - Set Thread Stack Size

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setstacksize(
       pthread_attr_t *attr,
       size_t          stacksize
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

DESCRIPTION:

The pthread_attr_setstacksize routine is used to set the stacksize attribute in the attr thread attribute object.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_ATTR_ STACKSIZE to indicate that this routine is supported.

If the specified stacksize is below the minimum required for this CPU (PTHREAD_STACK_MIN, then the stacksize will be set to the minimum for this CPU.

17.4.6 pthread attr getstacksize - Get Thread Stack Size

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getstacksize(
       const pthread_attr_t *attr,
       size_t               *stacksize
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The stacksize pointer argument is invalid.

DESCRIPTION:

The pthread_attr_getstacksize routine is used to obtain the stacksize attribute in the attr thread attribute object.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_ATTR_ STACKSIZE to indicate that this routine is supported.

17.4.7 pthread attr setstackaddr - Set Thread Stack Address

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setstackaddr(
       pthread_attr_t *attr,
       void           *stackaddr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

DESCRIPTION:

The pthread_attr_setstackaddr routine is used to set the stackaddr attribute in the attr thread attribute object.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_ATTR_ STACKADDR to indicate that this routine is supported.

It is imperative to the proper operation of the system that each thread have sufficient stack space.

17.4.8 pthread attr getstackaddr - Get Thread Stack Address

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getstackaddr(
       const pthread_attr_t  *attr,
       void                 **stackaddr
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The stackaddr pointer argument is invalid.

DESCRIPTION:

The pthread_attr_getstackaddr routine is used to obtain the stackaddr attribute in the attr thread attribute object.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_ATTR_ STACKADDR to indicate that this routine is supported.

17.4.9 pthread attr setscope - Set Thread Scheduling Scope

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setscope(
       pthread_attr_t *attr,
       int             contentionscope
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The contention scope specified is not valid.

ENOTSUP The contention scope specified (PTHREAD SCOPE SYSTEM) is not supported.

DESCRIPTION:

The pthread_attr_setscope routine is used to set the contention scope field in the thread attribute object attr to the value specified by contentionscope.

The contentionscope must be either PTHREAD_SCOPE_SYSTEM to indicate that the thread is to be within system scheduling contention or PTHREAD_SCOPE_PROCESS indicating that the thread is to be within the process scheduling contention scope.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.10 pthread attr getscope - Get Thread Scheduling Scope

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getscope(
       const pthread_attr_t *attr,
       int                  *contentionscope
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The contentionscope pointer argument is invalid.

DESCRIPTION:

The pthread_attr_getscope routine is used to obtain the value of the contention scope field in the thread attributes object attr. The current value is returned in

contentionscope.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.11 pthread attr setinheritsched - Set Inherit Scheduler Flag

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setinheritsched(
       pthread_attr_t *attr,
       int             inheritsched
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The specified scheduler inheritance argument is invalid.

DESCRIPTION:

The pthread_attr_setinheritsched routine is used to set the inherit scheduler field in the thread attribute object attr to the value specified by inheritsched.

The contentionscope must be either PTHREAD_INHERIT_SCHED to indicate that the thread is to inherit the scheduling policy and parameters fromthe creating thread, or PTHREAD_ EXPLICIT_SCHED to indicate that the scheduling policy and parameters for this thread are to be set from the corresponding values in the attributes object. If contentionscope is PTHREAD_INHERIT_SCHED, then the scheduling attributes in the attr structure will be ignored at thread creation time.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.12 pthread attr getinheritsched - Get Inherit Scheduler Flag

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getinheritsched(
       const pthread_attr_t *attr,
       int                  *inheritsched
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The inheritsched pointer argument is invalid.

DESCRIPTION:

The pthread_attr_getinheritsched routine is used to object the current value of the inherit scheduler field in the thread attribute object attr.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.13 pthread attr setschedpolicy - Set Scheduling Policy

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setschedpolicy(
       pthread_attr_t *attr,
       int             policy
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

ENOTSUP The specified scheduler policy argument is invalid.

DESCRIPTION:

The pthread_attr_setschedpolicy routine is used to set the scheduler policy field in the thread attribute object attr to the value specified by policy.

Scheduling policies may be one of the following:

    • SCHED_DEFAULT
    • SCHED_FIFO
    • SCHED_RR
    • SCHED_SPORADIC
    • SCHED_OTHER

The precise meaning of each of these is discussed elsewhere in this manual.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.14 pthread attr getschedpolicy - Get Scheduling Policy

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getschedpolicy(
       const pthread_attr_t *attr,
       int                  *policy
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The specified scheduler policy argument pointer is invalid.

DESCRIPTION:

The pthread_attr_getschedpolicy routine is used to obtain the scheduler policy field from the thread attribute object attr. The value of this field is returned in policy.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.15 pthread attr setschedparam - Set Scheduling Parameters

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_setschedparam(
       pthread_attr_t           *attr,
       const struct sched_param  param
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The specified scheduler parameter argument is invalid.

DESCRIPTION:

The pthread_attr_setschedparam routine is used to set the scheduler parameters field in the thread attribute object attr to the value specified by param.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.16 pthread attr getschedparam - Get Scheduling Parameters

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_attr_getschedparam(
       const pthread_attr_t *attr,
       struct sched_param   *param
     );

STATUS CODES:

EINVAL The attribute pointer argument is invalid.

EINVAL The attribute set is not initialized.

EINVAL The specified scheduler parameter argument pointer is invalid.

DESCRIPTION:

The pthread_attr_getschedparam routine is used to obtain the scheduler parameters field from the thread attribute object attr. The value of this field is returned in param.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.17 pthread create - Create a Thread

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_create(
       pthread_t             *thread,
       const pthread_attr_t  *attr,
       void                 (*start_routine)( void *),
       void                  *arg
     );

STATUS CODES:

EINVAL The attribute set is not initialized.

EINVAL The user specified a stack address and the size of the area was not large enough to meet this processor’s minimum stack requirements.

EINVAL The specified scheduler inheritance policy was invalid.

ENOTSUP The specified contention scope was PTHREAD SCOPE PROCESS.

EINVAL The specified thread priority was invalid.

EINVAL The specified scheduling policy was invalid.

EINVAL The scheduling policy was SCHED SPORADIC and the specified replenishment period is less than the initial budget.

EINVAL The scheduling policy was SCHED SPORADIC and the specified low priority is invalid.

EAGAIN The system lacked the necessary resources to create another thread, or the self imposed limit on the total number of threads in a process PTHREAD THREAD MAX would be exceeded.

EINVAL Invalid argument passed.

DESCRIPTION:

The pthread_create routine is used to create a new thread with the attributes specified by attr. If the attr argument is NULL, then the default attribute set will be used. Modification of the contents of attr after this thread is created does not have an impact on this thread.

The thread begins execution at the address specified by start_routine with arg as its only argument. If start_routine returns, then it is functionally equivalent to the thread executing the pthread_exit service.

Upon successful completion, the ID of the created thread is returned in the thread argument.

NOTES:

There is no concept of a single main thread in RTEMS as there is in a tradition UNIX system. POSIX requires that the implicit return of the main thread results in the same effects as if there were a call to exit. This does not occur in RTEMS.

The signal mask of the newly created thread is inherited from its creator and the set of pending signals for this thread is empty.

17.4.18 pthread exit - Terminate the Current Thread

CALLING SEQUENCE:

     #include <pthread.h>
     void pthread_exit(
       void *status
     );

STATUS CODES:

NONE

DESCRIPTION:

The pthread_exit routine is used to terminate the calling thread. The status is made available to any successful join with the terminating thread.

When a thread returns from its start routine, it results in an implicit call to the pthread_ exit routine with the return value of the function serving as the argument to pthread_exit.

NOTES:

Any cancellation cleanup handlers that hace been pushed and not yet popped shall be popped in reverse of the order that they were pushed. After all cancellation cleanup handlers have been executed, if the thread has any thread-specific data, destructors for that data will be invoked.

Thread termination does not release or free any application visible resources including byt not limited to mutexes, file descriptors, allocated memory, etc.. Similarly, exitting a thread does not result in any process-oriented cleanup activity.

There is no concept of a single main thread in RTEMS as there is in a tradition UNIX system. POSIX requires that the implicit return of the main thread results in the same effects as if there were a call to exit. This does not occur in RTEMS.

All access to any automatic variables allocated by the threads is lost when the thread exits. Thus references (i.e. pointers) to local variables of a thread should not be used in a global manner without care. As a specific example, a pointer to a local variable should NOT be used as the return value.

17.4.19 pthread detach - Detach a Thread

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_detach(
       pthread_t thread
     );

STATUS CODES:

ESRCH The thread specified is invalid.

EINVAL The thread specified is not a joinable thread.

DESCRIPTION:

The pthread_detach routine is used to to indicate that storage for thread can be reclaimed when the thread terminates without another thread joinging with it.

NOTES:

If any threads have previously joined with the specified thread, then they will remain joined with that thread. Any subsequent calls to pthread_join on the specified thread will fail.

17.4.20 pthread join - Wait for Thread Termination

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_join(
       pthread_t    thread,
       void       **value_ptr
     );

STATUS CODES:

ESRCH The thread specified is invalid.

EINVAL The thread specified is not a joinable thread.

EDEADLK A deadlock was detected or thread is the calling thread.

DESCRIPTION:

The pthread_join routine suspends execution of the calling thread until thread terminates. If thread has already terminated, then this routine returns immediately. The value returned by thread (i.e. passed to pthread_exit is returned in value_ptr.

When this routine returns, then thread has been terminated.

NOTES:

The results of multiple simultaneous joins on the same thread is undefined.

If any threads have previously joined with the specified thread, then they will remain joined with that thread. Any subsequent calls to pthread_join on the specified thread will fail.

If value ptr is NULL, then no value is returned.

17.4.21 pthread self - Get Thread ID CALLING SEQUENCE:

     #include <pthread.h>
     pthread_t pthread_self( void );

STATUS CODES:

The value returned is the ID of the calling thread.

DESCRIPTION:

This routine returns the ID of the calling thread.

NOTES:

NONE

17.4.22 pthread equal - Compare Thread IDs

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_equal(
        pthread_t t1,
        pthread_t t2
     );

STATUS CODES:

zero The thread ids are not equal.

non-zero The thread ids are equal.

DESCRIPTION:

The pthread_equal routine is used to compare two thread IDs and determine if they are equal.

NOTES:

The behavior is undefined if the thread IDs are not valid.

17.4.23 pthread once - Dynamic Package Initialization

CALLING SEQUENCE:

     #include <pthread.h>
     pthread_once_t once_control = PTHREAD_ONCE_INIT;
     int pthread_once(
        pthread_once_t   *once_control,
        void            (*init_routine)(void)
     );

STATUS CODES:

NONE

DESCRIPTION:

The pthread_once routine is used to provide controlled initialization of variables. The first call to pthread_once by any thread with the same once_control will result in the init_routine being invoked with no arguments. Subsequent calls to pthread_once with the same once_control will have no effect.

The init_routine is guaranteed to have run to completion when this routine returns to the caller.

NOTES:

The behavior of pthread_once is undefined if once_control is automatic storage (i.e. on a task stack) or is not initialized using PTHREAD_ONCE_INIT.

17.4.24 pthread setschedparam - Set Thread Scheduling Parameters

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_setschedparam(
       pthread_t           thread,
       int                 policy,
       struct sched_param *param
     );

STATUS CODES:

EINVAL The scheduling parameters indicated by the parameter param is invalid.

EINVAL The value specified by policy is invalid.

EINVAL The scheduling policy was SCHED SPORADIC and the specified replenishment period is less than the initial budget.

EINVAL The scheduling policy was SCHED SPORADIC and the specified low priority is invalid.

ESRCH The thread indicated was invalid.

DESCRIPTION:

The pthread_setschedparam routine is used to set the scheduler parameters currently associated with the thread specified by thread to the policy specified by policy. The contents of param are interpreted based upon the policy argument.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

17.4.25 pthread getschedparam - Get Thread Scheduling Parameters

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_getschedparam(
       pthread_t           thread,
       int                *policy,
       struct sched_param *param
     );

STATUS CODES:

EINVAL The policy pointer argument is invalid.

EINVAL The scheduling parameters pointer argument is invalid.

ESRCH The thread indicated by the parameter thread is invalid.

DESCRIPTION:

The pthread_getschedparam routine is used to obtain the scheduler policy and parameters associated with thread. The current policy and associated parameters values returned in policy and param, respectively.

NOTES:

As required by POSIX, RTEMS defines the feature symbol _POSIX_THREAD_PRIORITY_ SCHEDULING to indicate that the family of routines to which this routine belongs is supported.

18 Key Manager

18.1 Introduction

The key manager ... The directives provided by the key manager are:

• pthread_key_create - Create Thread Specific Data Key • pthread_key_delete - Delete Thread Specific Data Key • pthread_setspecific - Set Thread Specific Key Value • pthread_getspecific - Get Thread Specific Key Value

18.2 Background

There is currently no text in this section.

18.3 Operations

There is currently no text in this section.

18.4 Directives

This section details the key manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

18.4.1 pthread key create - Create Thread Specific Data Key

CALLING SEQUENCE:

     #include <pthread.h>
     int pthread_key_create(
       pthread_key_t *key,
       void (*destructor)( void )
     );

STATUS CODES:

EAGAIN There were not enough resources available to create another key.

ENOMEM Insufficient memory exists to create the key.

18.4.2 pthread key delete - Delete Thread Specific Data Key CALLING SEQUENCE:

#include <pthread.h>
int pthread_key_delete(
pthread_key_t key,
);

STATUS CODES:

EINVAL The key was invalid

DESCRIPTION: NOTES:

18.4.3 pthread setspecific - Set Thread Specific Key Value

CALLING SEQUENCE:

#include <pthread.h>
int pthread_setspecific(
pthread_key_t key,
const void *value
);

STATUS CODES:

EINVAL The specified key is invalid.

DESCRIPTION:

18.4.4 pthread getspecific - Get Thread Specific Key Value CALLING SEQUENCE:

#include <pthread.h>
void *pthread_getspecific(
pthread_key_t key
);

STATUS CODES:

NULL There is no thread-specific data associated with the specified key. non-NULL The data associated with the specified key.

DESCRIPTION: NOTES:

19 Thread Cancellation Manager

19.1 Introduction

The thread cancellation manager is ... The directives provided by the thread cancellation manager are:

• pthread_cancel - Cancel Execution of a Thread

• pthread_setcancelstate - Set Cancelability State

• pthread_setcanceltype - Set Cancelability Type

• pthread_testcancel - Create Cancellation Point

• pthread_cleanup_push - Establish Cancellation Handler • pthread_cleanup_pop - Remove Cancellation Handler

19.2 Background

There is currently no text in this section.

19.3 Operations

There is currently no text in this section.

19.4 Directives

This section details the thread cancellation manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

19.4.1 pthread cancel - Cancel Execution of a Thread CALLING SEQUENCE:

     int pthread_cancel(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

19.4.2 pthread setcancelstate - Set Cancelability State CALLING SEQUENCE:

     int pthread_setcancelstate(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

19.4.3 pthread setcanceltype - Set Cancelability Type CALLING SEQUENCE:

     int pthread_setcanceltype(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

19.4.4 pthread testcancel - Create Cancellation Point CALLING SEQUENCE:

     int pthread_testcancel(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

19.4.5 pthread cleanup push - Establish Cancellation Handler CALLING SEQUENCE:

     int pthread_cleanup_push(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

19.4.6 pthread cleanup pop - Remove Cancellation Handler CALLING SEQUENCE:

     int pthread_cleanup_push(
     );

STATUS CODES:

E The

DESCRIPTION: NOTES:

20 Services Provided by C Library (libc)

20.1 Introduction

This section lists the routines that provided by the Newlib C Library.

20.2 Standard Utility Functions (stdlib.h)

• abort - Abnormal termination of a program

• abs - Integer absolute value (magnitude)

• assert - Macro for Debugging Diagnostics

• atexit - Request execution of functions at program exit

• atof - String to double or float

• atoi - String to integer

• bsearch - Binary search

• calloc - Allocate space for arrays

• div - Divide two integers

• ecvtbuf - Double or float to string of digits

• ecvt - Double or float to string of digits (malloc result)

• __env_lock - Lock environment list for getenv and setenv • gvcvt - Format double or float as string

• exit - End program execution

• getenv - Look up environment variable

• labs - Long integer absolute value (magnitude)

• ldiv - Divide two long integers

• malloc - Allocate memory

• realloc - Reallocate memory

• free - Free previously allocated memory

• mallinfo - Get information about allocated memory

• __malloc_lock - Lock memory pool for malloc and free

• mbstowcs - Minimal multibyte string to wide string converter • mblen - Minimal multibyte length

• mbtowc - Minimal multibyte to wide character converter

• qsort - Sort an array

• rand - Pseudo-random numbers

• strtod - String to double or float

• strtol - String to long

• strtoul - String to unsigned long

• system - Execute command string

• wcstombs - Minimal wide string to multibyte string converter • wctomb - Minimal wide character to multibyte converter

20.3 Character Type Macros and Functions (ctype.h)

• isalnum - Alphanumeric character predicate

• isalpha - Alphabetic character predicate

• isascii - ASCII character predicate

• iscntrl - Control character predicate

• isdigit - Decimal digit predicate

• islower - Lower-case character predicate

• isprint - Printable character predicates (isprint, isgraph)

• ispunct - Punctuation character predicate

• isspace - Whitespace character predicate

• isupper - Uppercase character predicate

• isxdigit - Hexadecimal digit predicate

• toascii - Force integers to ASCII range

• tolower - Translate characters to lower case

• toupper - Translate characters to upper case

20.4 Input and Output (stdio.h)

• clearerr - Clear file or stream error indicator

• fclose - Close a file

• feof - Test for end of file

• ferror - Test whether read/write error has occurred

• fflush - Flush buffered file output

• fgetc - Get a character from a file or stream

• fgetpos - Record position in a stream or file

• fgets - Get character string from a file or stream

• fiprintf - Write formatted output to file (integer only)

• fopen - Open a file

• fdopen - Turn an open file into a stream

• fputc - Write a character on a stream or file

• fputs - Write a character string in a file or stream

• fread - Read array elements from a file

• freopen - Open a file using an existing file descriptor

• fseek - Set file position

• fsetpos - Restore position of a stream or file

• ftell - Return position in a stream or file

• fwrite - Write array elements from memory to a file or stream • getc - Get a character from a file or stream (macro)

• getchar - Get a character from standard input (macro)

• gets - Get character string from standard input (obsolete)

• iprintf - Write formatted output (integer only)

• mktemp - Generate unused file name

• perror - Print an error message on standard error

• putc - Write a character on a stream or file (macro)

• putchar - Write a character on standard output (macro)

• puts - Write a character string on standard output

• remove - Delete a file’s name

• rename - Rename a file

• rewind - Reinitialize a file or stream

• setbuf - Specify full buffering for a file or stream

• setvbuf - Specify buffering for a file or stream

• siprintf - Write formatted output (integer only)

• printf - Write formatted output

• scanf - Scan and format input

• tmpfile - Create a temporary file

• tmpnam - Generate name for a temporary file

• vprintf - Format variable argument list

20.5 Strings and Memory (string.h)

• bcmp - Compare two memory areas

• bcopy - Copy memory regions

• bzero - Initialize memory to zero

• index - Search for character in string

• memchr - Find character in memory

• memcmp - Compare two memory areas

• memcpy - Copy memory regions

• memmove - Move possibly overlapping memory

• memset - Set an area of memory

• rindex - Reverse search for character in string

• strcasecmp - Compare strings ignoring case

• strcat - Concatenate strings

• strchr - Search for character in string

• strcmp - Character string compare

• strcoll - Locale specific character string compare

• strcpy - Copy string

• strcspn - Count chars not in string

• strerror - Convert error number to string

• strlen - Character string length • strlwr - Convert string to lower case

• strncasecmp - Compare strings ignoring case

• strncat - Concatenate strings

• strncmp - Character string compare

• strncpy - Counted copy string

• strpbrk - Find chars in string

• strrchr - Reverse search for character in string

• strspn - Find initial match

• strstr - Find string segment

• strtok - Get next token from a string

• strupr - Convert string to upper case

• strxfrm - Transform string

20.6 Signal Handling (signal.h)

• raise - Send a signal

• signal - Specify handler subroutine for a signal

20.7 Time Functions (time.h)

• asctime - Format time as string

• clock - Cumulative processor time

• ctime - Convert time to local and format as string

• difftime - Subtract two times

• gmtime - Convert time to UTC (GMT) traditional representation • localtime - Convert time to local representation

• mktime - Convert time to arithmetic representation

• strftime - Flexible calendar time formatter

• time - Get current calendar time (as single number)

20.8 Locale (locale.h)

• setlocale - Select or query locale

20.9 Reentrant Versions of Functions

• Equivalent for errno variable:

      • errno_r - XXX

• Locale functions:

      • localeconv_r - XXX
      • setlocale_r - XXX

• Equivalents for stdio variables:

      • stdin_r - XXX

• stdout_r - XXX
• stderr_r - XXX

• Stdio functions:

• fdopen_r - XXX
• perror_r - XXX
• tempnam_r - XXX
• fopen_r - XXX
• putchar_r - XXX
• tmpnam_r - XXX
• getchar_r - XXX
• puts_r - XXX
• tmpfile_r - XXX
• gets_r - XXX
• remove_r - XXX
• vfprintf_r - XXX
• iprintf_r - XXX
• rename_r - XXX
• vsnprintf_r - XXX
• mkstemp_r - XXX
• snprintf_r - XXX
• vsprintf_r - XXX
• mktemp_t - XXX
• sprintf_r - XXX

• Signal functions:

• init_signal_r - XXX
• signal_r - XXX
• kill_r - XXX
• _sigtramp_r - XXX
• raise_r - XXX

• Stdlib functions:

• calloc_r - XXX
• mblen_r - XXX
• srand_r - XXX
• dtoa_r - XXX
• mbstowcs_r - XXX
• strtod_r - XXX
• free_r - XXX
• mbtowc_r - XXX
• strtol_r - XXX

• getenv_r - XXX
• memalign_r - XXX
• strtoul_r - XXX
• mallinfo_r - XXX
• mstats_r - XXX
• system_r - XXX
• malloc_r - XXX
• rand_r - XXX
• wcstombs_r - XXX
• malloc_r - XXX
• realloc_r - XXX
• wctomb_r - XXX
• malloc_stats_r - XXX
• setenv_r - XXX

• String functions:

• strtok_r - XXX

• System functions:

• close_r - XXX
• link_r - XXX
• unlink_r - XXX
• execve_r - XXX
• lseek_r - XXX
• wait_r - XXX
• fcntl_r - XXX
• open_r - XXX
• write_r - XXX
• fork_r - XXX
• read_r - XXX
• fstat_r - XXX
• sbrk_r - XXX
• gettimeofday_r - XXX
• stat_r - XXX
• getpid_r - XXX
• times_r - XXX

• Time function:

• asctime_r - XXX

20.10 Miscellaneous Macros and Functions

• unctrl - Return printable representation of a character

20.11 Variable Argument Lists

• Stdarg (stdarg.h):

      • va_start - XXX
      • va_arg - XXX
      • va_end - XXX

• Vararg (varargs.h):

      • va_alist - XXX
      • va_start-trad - XXX
      • va_arg-trad - XXX
      • va_end-trad - XXX

20.12 Reentrant System Calls

• open_r - XXX
• close_r - XXX
• lseek_r - XXX
• read_r - XXX
• write_r - XXX
• fork_r - XXX
• wait_r - XXX
• stat_r - XXX
• fstat_r - XXX
• link_r - XXX
• unlink_r - XXX
• sbrk_r - XXX

21 Services Provided by the Math Library (libm)

21.1 Introduction

This section lists the routines that provided by the Newlib Math Library (libm).

21.2 Standard Math Functions (math.h)

• acos - Arccosine

• acosh - Inverse hyperbolic cosine

• asin - Arcsine

• asinh - Inverse hyperbolic sine

• atan - Arctangent

• atan2 - Arctangent of y/x

• atanh - Inverse hyperbolic tangent

• jN - Bessel functions (jN and yN)

• cbrt - Cube root

• copysign - Sign of Y and magnitude of X • cosh - Hyperbolic cosine

• erf - Error function (erf and erfc)

• exp - Exponential

• expm1 - Exponential of x and - 1

• fabs - Absolute value (magnitude)

• floor - Floor and ceiling (floor and ceil)

• fmod - Floating-point remainder (modulo)

• frexp - Split floating-point number

• gamma - Logarithmic gamma function

• hypot - Distance from origin

• ilogb - Get exponent

• infinity - Floating infinity

• isnan - Check type of number

• ldexp - Load exponent

• log - Natural logarithms

• log10 - Base 10 logarithms

• log1p - Log of 1 + X

• matherr - Modifiable math error handler • modf - Split fractional and integer parts • nan - Floating Not a Number

• nextafter - Get next representable number • pow - X to the power Y • remainder - remainder of X divided by Y • scalbn - scalbn

• sin - Sine or cosine (sin and cos)

• sinh - Hyperbolic sine

• sqrt - Positive square root

• tan - Tangent

• tanh - Hyperbolic tangent

22 Status of Implementation

This chapter provides an overview of the status of the implementation of the POSIX API for RTEMS. The POSIX 1003.1b Compliance Guide provides more detailed information regarding the implementation of each of the numerous functions, constants, and macros specified by the POSIX 1003.1b standard. RTEMS supports many of the process and user/group oriented services in a "single user/single process" manner. This means that although these services may be of limited usefulness or functionality, they are provided and do work in a coherent manner. This is significant when porting existing code from UNIX to RTEMS.

• Implementation • The current implementation of dup() is insufficient.

• FIFOs mkfifo() are not currently implemented.

• Asynchronous IO is not implemented.

• The flockfile() family is not implemented

• getc/putc unlocked family is not implemented

• Shared Memory is not implemented

• Mapped Memory is not implemented

• NOTES:

• For Shared Memory and Mapped Memory services, it is unclear what

level of support is appropriate and possible for RTEMS.

• Functional Testing • Tests for unimplemented services

• Performance Testing • There are no POSIX Performance Tests.

• Documentation

• Many of the service description pages are not complete in this manual. These need to be completed and information added to the background and operations sections.

• Example programs (not just tests) would be very nice.

Command and Variable Index 319

Command and Variable Index

_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

A

access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 aio_cancel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 aio_error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 aio_fsync. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 aio_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 aio_return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 aio_suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 aio_write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 alarm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34, 35 asctime_r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160

C

cfgetispeed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 cfgetospeed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 cfsetispeed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 cfsetospeed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 clock_getres . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 clock_gettime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 clock_settime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 closedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 creat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 ctermid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 ctime_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

D

dup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 dup2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

E

execl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 execle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 execlp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 execv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 execve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 execvp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

F

fchdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 fchmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 fcntl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 fdatasync. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 fdopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 fileno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 flockfile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 fork. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 fpathconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 fstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 fsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 ftruncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 ftrylockfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 funlockfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

G

getc_unlocked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 getchar_unlocked. . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 getcwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 getdents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 getegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 geteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 getgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 getgrgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 getgrgid_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 getgrnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 getgrnam_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 getgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 getlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 getlogin_r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 getpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 getpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 getppid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 getpwnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 getpwnam_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 getpwuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 getpwuid_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gettimeofday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 getuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 gmtime_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

I

isatty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

K

kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

L

link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 lio_listio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 localtime_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 longjmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 lseek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

320 RTEMS POSIX API User’s Guide

lstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

M

microseonds alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 mkfifo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 mlock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 mlockall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 mmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 mount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 mprotect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 mq_attr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 mq_close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 mq_getattr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 mq_notify. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 mq_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 mq_receive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 mq_send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 mq_setattr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 mq_unlink. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 mqd_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 msync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 munlock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 munlockall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 munmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

N

nanosleep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

O

open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 opendir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

P

pathconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 pthread_atfork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 pthread_attr_destroy . . . . . . . . . . . . . . . . . . . . . . . 268

pthread_attr_getdetachstate . . . . . . . . . . . . . . . 270
pthread_attr_getinheritsched . . . . . . . . . . . . . . 278
pthread_attr_getschedparam . . . . . . . . . . . . . . . . 282
pthread_attr_getschedpolicy . . . . . . . . . . . . . . . 280

pthread_attr_getscope . . . . . . . . . . . . . . . . . . . . . . 276

pthread_attr_getstackaddr . . . . . . . . . . . . . . . . . 274
pthread_attr_getstacksize . . . . . . . . . . . . . . . . . 272

pthread_attr_init . . . . . . . . . . . . . . . . . . . . . . . . . . 267

pthread_attr_setdetachstate . . . . . . . . . . . . . . . 269
pthread_attr_setinheritsched . . . . . . . . . . . . . . 277
pthread_attr_setschedparam . . . . . . . . . . . . . . . . 281
pthread_attr_setschedpolicy . . . . . . . . . . . . . . . 279

pthread_attr_setscope . . . . . . . . . . . . . . . . . . . . . . 275

pthread_attr_setstackaddr . . . . . . . . . . . . . . . . . 273

pthread_attr_setstacksize . . . . . . . . . . . . . . . . . 271

pthread_cancel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 pthread_cleanup_pop . . . . . . . . . . . . . . . . . . . . . . . . 305 pthread_cleanup_push . . . . . . . . . . . . . . . . . . . . . . . 304 pthread_cond_broadcast. . . . . . . . . . . . . . . . . . . . . 213 pthread_cond_destroy . . . . . . . . . . . . . . . . . . . . . . . 211 pthread_cond_init . . . . . . . . . . . . . . . . . . . . . . . . . . 210 pthread_cond_signal . . . . . . . . . . . . . . . . . . . . . . . . 212 pthread_cond_timedwait. . . . . . . . . . . . . . . . . . . . . 215 pthread_cond_wait . . . . . . . . . . . . . . . . . . . . . . . . . . 214

pthread_condattr_destroy . . . . . . . . . . . . . . . . . . 207
pthread_condattr_getpshared . . . . . . . . . . . . . . . 209

pthread_condattr_init . . . . . . . . . . . . . . . . . . . . . . 206

pthread_condattr_setpshared . . . . . . . . . . . . . . . 208

pthread_create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 pthread_detach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 pthread_equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 pthread_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 pthread_getschedparam . . . . . . . . . . . . . . . . . . . . . . 292 pthread_join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 pthread_kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 pthread_mutex_destroy . . . . . . . . . . . . . . . . . . . . . . 198

pthread_mutex_getprioceiling . . . . . . . . . . . . . . 204

pthread_mutex_init . . . . . . . . . . . . . . . . . . . . . . . . . 197 pthread_mutex_lock . . . . . . . . . . . . . . . . . . . . . . . . . 199

pthread_mutex_setprioceiling . . . . . . . . . . . . . . 203
pthread_mutex_timedlock . . . . . . . . . . . . . . . . . . . 201

pthread_mutex_trylock . . . . . . . . . . . . . . . . . . . . . . 200 pthread_mutex_unlock . . . . . . . . . . . . . . . . . . . . . . . 202

pthread_mutexattr_destroy . . . . . . . . . . . . . . . . . 190
pthread_mutexattr_getprioceiling . . . . . . . . . 194
pthread_mutexattr_getprotocol. . . . . . . . . . . . . 192
pthread_mutexattr_getpshared . . . . . . . . . . . . . . 196

pthread_mutexattr_init. . . . . . . . . . . . . . . . . . . . . 189

pthread_mutexattr_setprioceiling . . . . . . . . . 193
pthread_mutexattr_setprotocol. . . . . . . . . . . . . 191
pthread_mutexattr_setpshared . . . . . . . . . . . . . . 195

pthread_once . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 pthread_self . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 pthread_setcancelstate. . . . . . . . . . . . . . . . . . . . . 301 pthread_setcanceltype . . . . . . . . . . . . . . . . . . . . . . 302 pthread_setschedparam . . . . . . . . . . . . . . . . . . . . . . 291 pthread_sigmask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 pthread_testcancel . . . . . . . . . . . . . . . . . . . . . . . . . 303 putc_unlocked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 putchar_unlocked. . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

R

rand_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 readdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 readlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 readv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 rewinddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

Command and Variable Index 321

S

scandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 sched_get_priority_max. . . . . . . . . . . . . . . . . . . . . 232 sched_get_priority_min. . . . . . . . . . . . . . . . . . . . . 231 sched_rr_get_interval . . . . . . . . . . . . . . . . . . . . . . 233 sched_yield . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 sem_close. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 sem_destroy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 sem_getvalue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 sem_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 sem_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 sem_post . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 sem_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 sem_timedwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 sem_trywait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 sem_unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 sem_wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 setgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 setjmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 setlocale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 setpgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 setsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 setuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 shm_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 shm_unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 sigaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 sigaddset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 sigdelset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 sigemptyset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 sigfillset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 sigismember . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 siglongjmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 sigpending. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 sigprocmask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 sigqueue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 sigsetjmp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 sigsuspend. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 sigtimedwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 sigwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 sigwaitinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 strtok_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 symlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 sysconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

T

tcdrain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 tcflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 tcflush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 tcgetattr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 tcgetpgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 tcsendbreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 tcsetattr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 tcsetpgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 telldir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ttyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 ttyname_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 tzset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

U

umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 uname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 unmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 usecs alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 usleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 utime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

W

wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 waitpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 writev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

Concept Index 323

Concept Index

A

acquire ownership of file stream. . . . . . . . . . . . . . . 147 add a signal to a signal set. . . . . . . . . . . . . . . . . . . . . 17 associate stream with file descriptor. . . . . . . . . . . 146 asynchronous file synchronization . . . . . . . . . . . . . 128 asynchronous read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 asynchronous write . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

B

broadcast a condition variable . . . . . . . . . . . . . . . . 213

C

cancel asynchronous i/o request. . . . . . . . . . . . . . . 126 cancel execution of a thread . . . . . . . . . . . . . . . . . . 300 change access and/or modification times of an

inode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 change memory protection . . . . . . . . . . . . . . . . . . . . 224 changes file mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 changes permissions of a file . . . . . . . . . . . . . . . . . . . 91 changes the current working directory . . . . . . 71, 72 changes the owner and/or group of a file. . . . . . . 93 check permissions for a file. . . . . . . . . . . . . . . . . . . . . 89 close a message queue . . . . . . . . . . . . . . . . . . . . . . . . 258 close a named semaphore . . . . . . . . . . . . . . . . . . . . . 180 closes a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 compare thread ids . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 create a directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 create a new file or rewrite an existing one . . . . . 76 create a process. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 create a thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 create an inter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 create cancellation point . . . . . . . . . . . . . . . . . . . . . . 303 create session and set process group id . . . . . . . . . 51 creates a link to a file . . . . . . . . . . . . . . . . . . . . . . . . . . 78 creates a symbolic link to a file . . . . . . . . . . . . . . . . 79

D

delay process execution . . . . . . . . . . . . . . . . . . 239, 240 delay with high resolution . . . . . . . . . . . . . . . . . . . . 241 delete a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 delete a signal from a signal set . . . . . . . . . . . . . . . . 18 destroy a condition variable. . . . . . . . . . . . . . . . . . . 211 destroy a condition variable attribute set. . . . . . 207 destroy a mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 destroy a mutex attribute set . . . . . . . . . . . . . . . . . 190 destroy a thread attribute set . . . . . . . . . . . . . . . . . 268 destroy an unnamed semaphore . . . . . . . . . . . . . . . 178 detach a thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 determine if file descriptor is terminal . . . . . . . . . . 60 determine terminal device name . . . . . . . . . . . . . . . 58 discards terminal data . . . . . . . . . . . . . . . . . . . . . . . . 138 duplicates an open file descriptor. . . . . . . . . 105, 106 dynamic package initialization . . . . . . . . . . . . . . . . 290 dynamically set the priority ceiling. . . . . . . . . . . . 203

E

empty a signal set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ends directory read operation . . . . . . . . . . . . . . . . . . 70 establish cancellation handler . . . . . . . . . . . . . . . . . 304 examine and change process blocked signals …. 24 examine and change signal action . . . . . . . . . . . . . . 22 examine and change thread blocked signals. . . . . 25 examine pending signals . . . . . . . . . . . . . . . . . . . . . . . 27 execute a file . . . . . . . . . . . . . . . . . . . . . . 5, 6, 7, 8, 9, 10

F

fill a signal set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

G

generate terminal pathname . . . . . . . . . . . . . . . . . . . 57 get character from stdin without locking . . . . . . 151 get character without locking . . . . . . . . . . . . . . . . . 150 get clock resolution . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 get configurable system variables. . . . . . . . . . . . . . . 61 get detach state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 get directory entries . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 get effective group id . . . . . . . . . . . . . . . . . . . . . . . . . . 44 get effective user id . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 get environment variables . . . . . . . . . . . . . . . . . . . . . . 55 get group file entry for id . . . . . . . . . . . . . . . . . . . . . 166 get group file entry for name. . . . . . . . . . . . . . . . . . 168 get inherit scheduler flag. . . . . . . . . . . . . . . . . . . . . . 278 get maximum priority value . . . . . . . . . . . . . . . . . . 232 get message queue attributes . . . . . . . . . . . . . . . . . 264 get minimum priority value . . . . . . . . . . . . . . . . . . . 231 get parent process id. . . . . . . . . . . . . . . . . . . . . . . . . . . 40 get password file entry for uid . . . . . . . . . . . . . . . . 170 get process group id . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 get process id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 get process shared attribute . . . . . . . . . . . . . . . . . . 209 get process times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 get real group id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 get scheduling parameters . . . . . . . . . . . . . . . . . . . . 282 get scheduling policy . . . . . . . . . . . . . . . . . . . . . . . . . 280 get supplementary group ids . . . . . . . . . . . . . . . . . . . 47 get system name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 get the blocking protocol . . . . . . . . . . . . . . . . . . . . . 192 get the current priority ceiling . . . . . . . . . . . . . . . . 204 get the priority ceiling . . . . . . . . . . . . . . . . . . . . . . . . 194 get the time of day . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 get the value of a semaphore. . . . . . . . . . . . . . . . . . 186 get the visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

324 RTEMS POSIX API User’s Guide

get thread id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 get thread scheduling parameters . . . . . . . . . . . . . 292 get thread scheduling scope . . . . . . . . . . . . . . . . . . . 276 get thread stack address . . . . . . . . . . . . . . . . . . . . . . 274 get thread stack size . . . . . . . . . . . . . . . . . . . . . . . . . . 272 get time in seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 get timeslicing quantum . . . . . . . . . . . . . . . . . . . . . . 233 get user id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 get user name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 get user name, reentrant . . . . . . . . . . . . . . . . . . . . . . . 49 gets configuration values for files . . . . . . . . . . . 97, 99 gets current working directory . . . . . . . . . . . . . . . . . 73 gets file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 88 gets foreground process group id . . . . . . . . . . . . . . 140 gets information about a file . . . . . . . . . . . . . . . . . . . 86 gets terminal attributes. . . . . . . . . . . . . . . . . . . . . . . 134

I

initialize a condition variable . . . . . . . . . . . . . . . . . 210 initialize a condition variable attribute set …. 206 initialize a mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 initialize a mutex attribute set . . . . . . . . . . . . . . . . 189 initialize a thread attribute set. . . . . . . . . . . . . . . . 267 initialize an unnamed semaphore. . . . . . . . . . . . . . 177 initialize time conversion information . . . . . . . . . 158 is signal a member of a signal set . . . . . . . . . . . . . . 20

L

list directed i/o. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 lock a mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 lock a mutex with timeout . . . . . . . . . . . . . . . . . . . . 201 lock a range of the process address space. . . . . . 220 lock the address space of a process . . . . . . . . . . . . 218

M

makes a directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 makes a fifo special file . . . . . . . . . . . . . . . . . . . . . . . . 82 manipulates an open file descriptor . . . . . . . . . . . 111 map process addresses to a memory object …. 222 memory object synchronization . . . . . . . . . . . . . . . 225 microsecond delay process execution . . . . . . . . . . 240 mount a file system. . . . . . . . . . . . . . . . . . . . . . . . . . . 117

N

non . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155, 157, 183 notify process that a message is available . . . . . 262

O

obtain file descriptor number for this file . . . . . . 145 obtain the name of a symbolic link destination . . 80 obtain time of day . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 open a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 open a message queue . . . . . . . . . . . . . . . . . . . . . . . . 256 open a named semaphore . . . . . . . . . . . . . . . . . . . . . 179 open a shared memory object . . . . . . . . . . . . . . . . . 226 opens a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

P

password file entry for name . . . . . . . . . . . . . . . . . . 172 poll to acquire ownership of file stream. . . . . . . . 148 poll to lock a mutex . . . . . . . . . . . . . . . . . . . . . . . . . . 200 put character to stdin without locking . . . . . . . . 153 put character without locking. . . . . . . . . . . . . . . . . 152

Q

queue a signal to a process. . . . . . . . . . . . . . . . . . . . . 33

R

reads a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 reads from a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 reads terminal input baud rate. . . . . . . . . . . . . . . . 130 reads terminal output baud rate . . . . . . . . . . . . . . 131 receive a message from a message queue . . . . . . 261 reentrant determine terminal device name . . . . . . 59 reentrant extract token from string . . . . . . . . . . . 159 reentrant get group file entry . . . . . . . . . . . . . . . . . 167 reentrant get group file entry for name . . . . . . . . 169 reentrant get password file entry for name …. 173 reentrant get password file entry for uid. . . . . . . 171 reentrant get user name. . . . . . . . . . . . . . . . . . . . . . . . 49 reentrant local time conversion. . . . . . . . . . . . . . . . 163 reentrant random number generation . . . . . . . . . 164 reentrant struct tm to ascii time conversion … 160 reentrant time t to ascii time conversion . . . . . . 161 reentrant utc time conversion . . . . . . . . . . . . . . . . . 162 register fork handlers . . . . . . . . . . . . . . . . . . . . . . . . . . 11 release ownership of file stream . . . . . . . . . . . . . . . 149 remove a message queue . . . . . . . . . . . . . . . . . . . . . . 259 remove a shared memory object . . . . . . . . . . . . . . 227 remove cancellation handler . . . . . . . . . . . . . . . . . . 305 removes a directory entry . . . . . . . . . . . . . . . . . . . . . . 83 renames a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 reposition read/write file offset. . . . . . . . . . . . . . . . 113 resets the readdir() pointer . . . . . . . . . . . . . . . . . . . . 67 retrieve error status of asynchronous i/o operation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 retrieve return status asynchronous i/o operation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 return current location in directory stream . . . . . 69

S

save context for non . . . . . . . . . . . . . . . . . . . . . . . . . . 154 save context with signal status for non . . . . . . . . 156 scan a directory for matching entries . . . . . . . . . . . 68 schedule alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 schedule alarm in microseonds . . . . . . . . . . . . . . . . . 35

Concept Index 325

send a message to a message queue . . . . . . . . . . . 260 send a signal to a process . . . . . . . . . . . . . . . . . . . . . . 26 send a signal to a thread. . . . . . . . . . . . . . . . . . . . . . . 23 sends a break to a terminal . . . . . . . . . . . . . . . . . . . 136 set cancelability state. . . . . . . . . . . . . . . . . . . . . . . . . 301 set cancelability type . . . . . . . . . . . . . . . . . . . . . . . . . 302 set detach state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 set environment variables . . . . . . . . . . . . . . . . . . . . . . 56 set group id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 set inherit scheduler flag . . . . . . . . . . . . . . . . . . . . . . 277 set message queue attributes. . . . . . . . . . . . . . . . . . 263 set process group id for job control. . . . . . . . . . . . . 52 set process shared attribute. . . . . . . . . . . . . . . . . . . 208 set scheduling parameters. . . . . . . . . . . . . . . . . . . . . 281 set scheduling policy. . . . . . . . . . . . . . . . . . . . . . . . . . 279 set terminal attributes . . . . . . . . . . . . . . . . . . . . . . . . 135 set the blocking protocol. . . . . . . . . . . . . . . . . . . . . . 191 set the current locale . . . . . . . . . . . . . . . . . . . . . . . . . 144 set the priority ceiling . . . . . . . . . . . . . . . . . . . . . . . . 193 set the visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 set thread scheduling parameters . . . . . . . . . . . . . 291 set thread scheduling scope . . . . . . . . . . . . . . . . . . . 275 set thread stack address . . . . . . . . . . . . . . . . . . . . . . 273 set thread stack size . . . . . . . . . . . . . . . . . . . . . . . . . . 271 set time of day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 set user id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 sets a file creation mask. . . . . . . . . . . . . . . . . . . . . . . . 77 sets foreground process group id . . . . . . . . . . . . . . 141 sets terminal input baud rate . . . . . . . . . . . . . . . . . 132 sets terminal output baud rate. . . . . . . . . . . . . . . . 133 signal a condition variable . . . . . . . . . . . . . . . . . . . . 212 suspend process execution . . . . . . . . . . . . . . . . . . . . . 29 suspends/restarts terminal output.. . . . . . . . . . . . 139 synchronize file complete in . . . . . . . . . . . . . . . . . . . 114 synchronize file in . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 synchronize file systems. . . . . . . . . . . . . . . . . . . . . . . 116 synchronously accept a signal . . . . . . . . . . . . . . 30, 31 synchronously accept a signal with timeout …. 32

T

terminate a process . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 terminate the current thread. . . . . . . . . . . . . . . . . . 285 truncate a file to a specified length . . . . . . . . . 95, 96

U

unlink a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 unlock a mutex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 unlock a range of the process address space … 221 unlock a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . 185 unlock the address space of a process . . . . . . . . . 219 unmap previously mapped addresses . . . . . . . . . . 223 unmount file systems . . . . . . . . . . . . . . . . . . . . . . . . . 118 usecs delay process execution . . . . . . . . . . . . . . . . . 240

V

vectored read from a file . . . . . . . . . . . . . . . . . . . . . . 119 vectored write to a file. . . . . . . . . . . . . . . . . . . . . . . . 120

W

wait for a signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 wait for asynchronous i/o request . . . . . . . . . . . . . 127 wait for process termination. . . . . . . . . . . . . . . . 12, 13 wait for thread termination . . . . . . . . . . . . . . . . . . . 287 wait on a condition variable . . . . . . . . . . . . . . . . . . 214 wait on a semaphore. . . . . . . . . . . . . . . . . . . . . . . . . . 182 wait on a semaphore for a specified time . . . . . . 184 wait with timeout a condition variable . . . . . . . . 215 waits for all output to be transmitted to the

terminal.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 writes to a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

Y

yield the processor. . . . . . . . . . . . . . . . . . . . . . . . . . . . 234